### v2

# On the existence of weighted-cscK metrics

Abstract.
In this paper, we prove that on a smooth Kähler manifold, the ${\mathbb{G}}$-coercivity of the weighted Mabuchi functional implies the existence of the $(\mathrm{v},\mathrm{w})$-weighted-cscK (extremal) metric with $\mathrm{v}$ log-concave (firstly studied in [Lah19]), e.g, extremal metrics, Kähler–Ricci solitons, $\mu$-cscK metrics.
[…]
1. Introduction
In this paper, we apply the variational approach and continuity method to study the weighted-cscK (extremal) metrics, the study of which is initiated in papers [Lah19, Lah23, AJL23]. The weighted-cscK (extremal) metric is a natural generalization of the cscK (extremal) metric. The (generalized) Yau–Tian–Donaldson conjecture predicts an equivalence between the existence of cscK metric(s) and the ${\mathbb{G}}$-uniform K-stability condition. By the work of many authors (see an incomplete list: [Tia97, DR17, BDL20, CC21a, CC21b, Li22, BJ23]), much progress has been made, and it is known that the existence of cscK metric(s) is equivalent to the ${\mathbb{G}}$-coercivity of the Mabuchi functional, and can be implied by the ${\mathbb{G}}$-uniformly K-stability for models.
The category of weighted-cscK metrics includes a large class of important metrics, such as Kähler–Einstein metrics, cscK metrics, Kähler–Ricci solitons, Mabuchi solitons, $\mu$-cscK metrics, etc. Indeed, the log-concavity assumption of our main theorem 1.2 is satisfied in all these examples. It provides a universal framework to treat these different canonical metrics. This viewpoint has some interesting applications, e.g, in [Li21], the author finds the Ricci-flat Kähler cone metric on an irregular cone by transforming it (within the weighted-cscK category) to a $g$-soliton on a quasi-regular cone. Because of its importance, it is attempting to generalize the Yau–Tian–Donaldson type theory to the weighted-cscK case which would give an algebraic criterion for the resolution of the equation.
Let $(X,\omega)$ be a compact Kähler manifold. Let $T$ be an $\ell$-dimensional real torus generated by Hamiltonian vector fields of $(X,\omega)$. The complexified torus $T_{\mathbb{C}}$ acts holomorphically on $(X,\omega)$. The Lie algebra $\mathfrak{t}:=\mathrm{Lie}(T)$ is identified with $N_{\mathbb{R}}\simeq\mathbb{R}^{\ell}$. Similarly, we identify the dual $\mathfrak{t}^{*}$ with $M_{\mathbb{R}}\simeq\mathbb{R}^{\ell}$ with the corresponding dual real coordinates, which will be denote by $y:=(y_{1},\cdots,y_{\ell})$.
Let $\{\xi^{\alpha}\}$ be a standard basis of $\mathfrak{t}$, and let $m^{\xi^{\alpha}}_{\phi}\in C^{\infty}(X)$ be the associated Hamiltonian function for a Kähler form $\omega_{\phi}\in[\omega]$. We have
$\iota_{\xi^{\alpha}}\omega_{\phi}=\dfrac{\sqrt{-1}}{2\pi}\overline{\partial}m^{\xi^{\alpha}}_{\phi}.$
By [Ati82], the image of $m_{\phi}:X\rightarrow\mathfrak{t}^{*}=\mathbb{R}^{\ell}$ is a convex polytope $P$, which is called the associated moment polytope. It is explicitly given by
$m_{\phi}(x):=(m_{\phi}^{\xi^{1}}(x),\cdots,m_{\phi}^{\xi^{\ell}}(x)),\quad P:=m_{\phi}(X).$
We use $\alpha$, $\beta$ to represent the index of the Lie algebra $\mathfrak{t}$, and $i,j$ to represent the index of base manifold $X$. Let $y=(y_{\alpha})\in P$ denote for the coordinate. For any smooth function $\mathrm{v}$ on $P$, we will always use $\mathrm{v}_{,\alpha}$ to denote the derivative of $\mathrm{v}$ over $y_{\alpha}$.
Definition 1.1.
We define the Lie algebra
${{\mathrm{aut}}}_{T}(X)=\{\xi\in{{\mathrm{aut}}(X)}:[\xi,c]=0,\forall c\in\mathfrak{t}_{\mathbb{C}}\}.$
Let ${\rm Aut}_{T}(X)$ be the connected subgroup of $\rm Aut(X)$ that is generated by ${\mathrm{aut}}_{T}(X)$. Let ${\mathbb{G}}=K_{\mathbb{C}}\subset{\rm Aut}_{T}(X)$ be a reductive Lie group, and $K$ contains a maximal torus of ${\rm Aut}_{T}(X)$. Denote ${\mathbb{T}}$ as the center of ${\mathbb{G}}$. Note that, since $T$ is a subgroup of the center of ${\rm Aut}_{T}(X)$ and $K$ contains a maximal torus, we have $T\subset K$.
In [Lah19], the author considers the following weighted-cscK equation
${S_{\mathrm{v}}(\phi)}={\mathrm{w}(m_{\phi})}\cdot\ell_{\rm ext}(m_{\phi}),$ (1.1)
where $\mathrm{v},\mathrm{w}\in C^{\infty}(P,{\mathbb{R}}_{>0})$, $\ell_{\mathrm{ext}}$ is an affine function on $P$, and the weighted-scalar curvature is given by
$S_{\mathrm{v}}(\phi):=\mathrm{v}(m_{\phi})S(\phi)-2\mathrm{v}_{,\alpha}(m_{\phi})\Delta_{\phi}(m_{\phi}^{\xi^{\alpha}})-\mathrm{v}_{,\alpha\beta}(m_{\phi})\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}.$
Note that when $\mathrm{v}=\mathrm{w}=const$, the solution of (1.1) is the extremal metric.
Equation (1.1) can be equivalently written as a system of second-order PDEs:
$\mathrm{v}(m_{\phi})(\omega+{\rm dd^{c}}\phi)^{n}=e^{F}\omega^{n},$ (1.2a)
$\left(\Delta_{\phi}-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}\right)F=-\ell_{\mathrm{ext}}(m_{\phi})\frac{\mathrm{w}(m_{\phi})}{\mathrm{v}(m_{\phi})}+\frac{\mathrm{tr}_{\mathrm{v},\phi}\mathrm{Ric}(\omega)}{\mathrm{v}(m_{\phi})},$ (1.2b)
where for a $(1,1)$-form $\eta$, $\mathrm{tr}_{\mathrm{v},\phi}(\eta)$ is defined by
$\mathrm{tr}_{\mathrm{v},\phi}(\eta):=\mathrm{v}(m_{\phi})\mathrm{tr}_{\phi}\eta+\langle(\mathrm{d}\mathrm{v})(m_{\phi}),m_{\eta}\rangle,$
where $\langle(\mathrm{d}\mathrm{v})(m_{\phi}),m_{\eta}\rangle$ is uniformly bounded independent of $\phi$. For the detail definition of $m_{\eta}$ see Subsection 2.A.
From the variational point of view, the equation (1.1) can also be considered as the Euler-Lagrangian of the weighted Mabuchi functional ${\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}}}$ (see Definition 2.5). In [Lah23], the author shows that the weighted-cscK metric is a minimizer of ${\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}}}$ and such metric is unique up to automorphism. It is proved in [AJL23] that the existence of the solution of (1.1) implies the ${\mathbb{G}}$-coercivity of ${\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}}}$. In this article, we aim to prove the inverse direction and conclude the following theorem:
Theorem 1.2.
Assume $\mathrm{v}$ is log-concave. The following statements are equivalent:
(1). Up to the automorphism group ${\rm Aut}_{T}(X)$, there exists a unique solution of the weighted-cscK (extremal) equation (1.1).
(2). The weighted Mabuchi functional for weighted-cscK (extremal) equation is $\mathbb{G}$-coercive, i.e,
${\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}}}(\phi)\geq\delta\cdot{\mathbb{J}}_{\mathrm{v},{\mathbb{T}}}(\phi)-C,$
for some constants $\delta>0$, $C>0$, and any $\phi\in{\mathcal{E}}^{1}_{K}(X,\omega)$.
To show (2) implies (1), we consider the following weighted version of the continuity path of Chen [Che18] for weighted extremal metrics,
$t\left(\frac{S_{\mathrm{v}}(\phi)}{\mathrm{w}(m_{\phi})}-\ell_{\rm ext}(m_{\phi})\right)-(1-t)\left(\frac{\mathrm{tr}_{\mathrm{v},\phi}(\theta)}{\mathrm{w}(m_{\phi})}-\underline{\theta}\right)=0,$ (1.3)
where $\theta$ is a $K$-invariant smooth positive $(1,1)$-form in the same class of $\omega$.
Remark 1.3.
The idea of the continuity path (1.3) can be seen from the variational picture. Since $-\mathrm{Ric}(\omega)$ may not be positive, replacing a small part of the term $-{\mathbb{E}}_{\mathrm{v}}^{\mathrm{Ric}}$ in the weighted Mabuchi functional by ${\mathbb{E}}_{\mathrm{v}}^{\theta}$ for some positive (1,1)-form $\theta$ will make the weighted Mabuchi functional more coercive. Intuitively, this makes (1.3) always solvable for $t\in[0,1)$.
Such continuity method is used in [CC21a, CC21b] for the resolution of the cscK problem. The idea of using continuity path to solve Monge-Ampère type equations dates back to the fundamental work of [Yau78].
Similar with [CC21a], the equation (1.3) can be re-written as a coupled PDE system
$\mathrm{v}(m_{\phi})\omega_{\phi}^{n}=e^{F}\omega^{n},$ (1.4a)
$\left(\Delta_{\phi}-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}\right)F=-\ell_{\mathrm{ext}}(m_{\phi})\frac{\mathrm{w}(m_{\phi})}{\mathrm{v}(m_{\phi})}+(1-\frac{1}{t})\frac{\mathrm{tr}_{\mathrm{v},\phi}(\theta)}{\mathrm{v}(m_{\phi})}$
$\qquad\qquad\qquad\qquad\qquad\qquad-(1-\frac{1}{t})\underline{\theta}\frac{\mathrm{w}(m_{\phi})}{\mathrm{v}(m_{\phi})}+\frac{\mathrm{tr}_{\mathrm{v},\phi}\mathrm{Ric}(\omega)}{\mathrm{v}(m_{\phi})}.$ (1.4b)
Compared with the equations in [CC21a], the first equation (1.4a) is a weighted Monge-Ampère equation, and the second linear equation (1.4b) also contains first-order derivative term of $F$ (which is a third-order derivative of $\phi$). This makes the $C^{0}$-estimate and $C^{2}$-estimate of $\phi$, the gradient estimate of $F$ more challenging.
We prove a uniform $C^{0}$-estimate by modifying Guo-Phong’s argument in [GP22]. In addition, in the calculation of the $C^{2}$-estimate of $\phi$ and the gradient estimate of $F$, we need to modify Chen-Cheng’s method in [CC21b] (in particular, we need to choose new test functions and integral weights in order to deal with the third-order derivative terms (of $\phi$)).
Organization
In Section 2, we recall some backgrounds on weighted-cscK metrics and re-write the weighted-cscK equation to a coupled PDE system.
In Section 3, we consider a modified coupled PDE system and prove the uniform $C^{0}$ estimate and $W^{2,p}$-estimate for $\phi$, and the gradient estimate for $F+f_{*}-\log(\mathrm{v}(m_{\phi}))$. Essentially, we only need the log-concavity assumption in the proof of $W^{2,p}$-estimate.
In Section 4, we prove Theorem 1.2 by using the continuity method. We give a sketched proof of the openness part in Subsection 4.A, where we follow the argument in [Che18], [Has19]. We prove the closedness part in Subsection 4.B, where we follow the strategy of [CC21b].
In Section 5 (Appendix), we provide $C^{2}$-estimate and higher-order estimates for the weighted-cscK equation. In addition, we give a detailed computation of the linearized operators used in the openness part.
Relation with Di-Nezza–Jubert–Lahdili’s work
As our project was nearing completion, we were informed that Eleonora Di-Nezza, Simon Jubert and Abdellah Lahdili also proved the a priori estimates in [DNJL24] by using a method quite different from ours. Our two groups finished our projects independently.
Acknowledgement
The authors would like to thank Abdellah Lahdili for sharing his idea on showing the openness near the starting point of the continuity path around 2020. The authors would like to thank Eleonora Di Nezza, Simon Jubert, Abdellah Lahdili for pointing out an error (which causes a problem in $W^{2,p}$ estimate if we don’t assume the log-concavity of $\mathrm{v}$) in the first version of our draft. The authors would like to thank Jianchun Chu, Lifan Guan, Chang Li for their helpful discussions. The authors would like to thank Prof. Gang Tian, Akito Futaki, Chi Li, Eiji Inoue, Kewei Zhang for their interest and helpful comments. J. Han is supported by National Key R&D Program of China 2023YFA1009900, NSFC-20701003A012401.
2. Preliminaries
In this section, we will collect the foundations of the variational approach and continuity method that are needed in further analysis. In particular, we will review the weighted Monge-Ampère measure $\mathrm{v}(m_{\phi})\omega_{\phi}^{n}$, ${\mathcal{E}}^{1}$-space and $\mathrm{d}_{1}$-distance, the functionals ${\mathbb{M}}_{\mathrm{v},\mathrm{w}},{\mathbb{J}}_{\mathrm{v}},{\mathbb{E}}_{\mathrm{v}}$, coercivity, weighted-cscK (extremal) equations and continuity path.
2.A. Weighted Monge-Ampère measures, space of potentials, and functionals
The variational approach has successful applications in the resolution of Kähler-Einstein problem, see [Tia97, BBGZ13, BBE+19, BBJ21]. In this subsection, we will define the notions needed in the variational approach. We will follow the lines in [BN14, HL23, AJL23].
We will consider $K$-invariant $\omega$-psh functions:
[…]
In [BN14], the authors define the weighted Monge-Ampère measure by using Kiselman’s partial Legendre transform and approximation by series of step functions. Then in [HL23], the authors present an alternative definition by using a more direct fibration construction, which is more adapted in defining the non-Archimedean functionals.
Proposition 2.1 ([BN14, HL23]).
Let $\mathrm{v}$ be a smooth positive function on the moment polytope $P$. For any $\omega$-psh function $\phi\in\mathrm{Psh}_{K}(X,\omega)$, the weighted Monge-Ampère measure ${\rm MA}_{\mathrm{v}}(\phi):=\mathrm{v}(m_{\phi})\omega_{\phi}^{n}$ is well-defined as a Radon measure.
The following weighted functionals and potential spaces are defined in [BN14, HL23, Lah19, AJL23].
Definition 2.2.
For $\phi\in\mathrm{Psh}_{K}(X,\omega)$, $\phi=\phi_{0}+u$, $\phi_{t}=\phi_{0}+tu$ ($0\leq t\leq 1$), we define the weighted energy functionals:
${\mathbb{E}}_{\mathrm{v}}(\phi,\phi_{0})=\frac{1}{\mathbb{V}_{\mathrm{v}}(X)}\int_{0}^{1}\int_{X}(\phi-\phi_{0})\mathrm{v}(m_{\phi_{t}})\frac{\omega_{\phi_{t}}^{n}}{n!}\mathrm{d}t,$ (2.2)
${\mathbb{E}}_{\mathrm{v}}^{\eta}(\phi,\phi_{0})=\frac{1}{\mathbb{V}_{\mathrm{v}}(X)}\int_{0}^{1}\int_{X}(\phi-\phi_{0})\Big{(}\mathrm{v}(m_{\phi_{t}})\eta\wedge\frac{\omega_{\phi_{t}}^{n-1}}{(n-1)!}+\langle\mathrm{d}\mathrm{v}(m_{\phi_{t}}),m_{\eta}\rangle_{\phi_{t}}\frac{\omega_{\phi_{t}}^{n}}{n!}\Big{)}\mathrm{d}t,$ (2.3)
where $\eta$ is a smooth $(1,1)$-form, which can be locally expressed as $\eta={\rm dd^{c}}\psi$. And $m_{\eta}(\xi):=-J\xi(\psi)$, which is globally well-defined. And $\langle d\mathrm{v}(m_{\phi},m_{\eta}\rangle_{\phi}=-\sum_{\alpha}\mathrm{v}_{,\alpha}J\xi^{\alpha}(\psi)$.
We denote ${\mathbb{E}}_{\mathrm{v}}(\phi)={\mathbb{E}}_{\mathrm{v}}(\phi,0)$. (The same convention applies for other energy functionals.) We define the weighted I-functional:
${\mathbb{I}}_{\mathrm{v}}(\phi,\phi_{0})=\frac{1}{\mathbb{V}_{\mathrm{v}}(X)}\int_{X}(\phi-\phi_{0})\left(\mathrm{v}(m_{\phi_{0}})\frac{\omega_{\phi_{0}}^{n}}{n!}-\mathrm{v}(m_{\phi})\frac{\omega_{\phi}^{n}}{n!}\right),$ (2.4)
and the weighted J-functional:
${\mathbb{J}}_{\mathrm{v}}(\phi,\phi_{0})$ (2.5)
$=$ $\frac{1}{\mathbb{V}_{\mathrm{v}}(X)}\int_{X}(\phi-\phi_{0})\mathrm{v}(m_{\phi_{0}})\frac{\omega_{\phi_{0}}^{n}}{n!}-{\mathbb{E}}_{\mathrm{v}}(\phi,\phi_{0})=:\Lambda_{\mathrm{v}}(\phi,\phi_{0})-{\mathbb{E}}_{\mathrm{v}}(\phi,\phi_{0})$
$=$ $\int_{X}\sum_{k=0}^{n-1}\int_{0}^{1}\int_{0}^{t}\mathrm{v}(m_{\phi_{s}})\frac{1}{k!(n-k-1)!}(1-s)^{k}s^{n-k-1}\mathrm{d}s\mathrm{d}t\cdot\mathrm{d}(\phi-\phi_{0})\wedge\mathrm{d}^{\rm c}(\phi-\phi_{0})\wedge\omega^{k}\wedge\omega_{\phi}^{n-k-1}.$
We also define the reduced weighted J-functional:
${\mathbb{J}}_{\mathrm{v},{\mathbb{T}}}(\phi)=\inf_{\sigma\in{\mathbb{T}}}{\mathbb{J}}_{\mathrm{v}}(\sigma^{*}\phi).$ (2.6)
Remark 2.3.
There exists a constant $C$ which depends on $\mathrm{v}$, such that
$\frac{1}{C}{\mathbb{J}}(\phi)\leq{\mathbb{J}}_{\mathrm{v}}(\phi)\leq C{\mathbb{J}}(\phi),$
and
$\frac{1}{C}\big{(}{\mathbb{I}}(\phi)-{\mathbb{J}}(\phi)\big{)}\leq{\mathbb{I}}_{\mathrm{v}}(\phi)-{\mathbb{J}}_{\mathrm{v}}(\phi)\leq C\big{(}{\mathbb{I}}(\phi)-{\mathbb{J}}(\phi)\big{)}.$
Definition 2.4.
We define the following spaces of potentials:
${\mathcal{E}}_{K}(X,\omega)=\{\phi\in\mathrm{Psh}_{K}(X,\omega):\int_{X}\mathrm{v}(m_{\phi})\frac{\omega_{\phi}^{n}}{n!}=\int_{X}\mathrm{v}(m_{\omega})\frac{\omega^{n}}{n!}=:\mathbb{V}_{\mathrm{v}}(X)\},$ (2.7)
${\mathcal{E}}^{1}_{K}(X,\omega)=\{\phi\in{\mathcal{E}}_{K}(X,\omega):{\mathbb{E}}_{\mathrm{v}}(\phi)>-\infty\}.$ (2.8)
Definition 2.5.
For any $\phi\in{\mathcal{E}}^{1}_{K}(X,\omega)$, we define the (weighted) Mabuchi functional (for weighted-cscK problem) as:
${\mathbb{M}}_{\mathrm{v},\mathrm{w}}(\phi)=$ $\int_{X}\log(\frac{\mathrm{v}(m_{\phi})\omega_{\phi}^{n}}{\omega^{n}})\mathrm{v}(m_{\phi})\frac{\omega_{\phi}^{n}}{n!}-\int_{X}\log(\mathrm{v}(m_{\omega}))\mathrm{v}(m_{\omega})\frac{\omega^{n}}{n!}$
$-{\mathbb{E}}_{\mathrm{v}}^{\mathrm{Ric}(\omega)}(\phi)+{\mathbb{E}}_{\mathrm{w}}(\phi).$
We also define the (weighted) Mabuchi for (weighted-extremal problem) as:
${\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}}}(\phi)=$ $\int_{X}\log(\frac{\mathrm{v}(m_{\phi})\omega_{\phi}^{n}}{\omega^{n}})\mathrm{v}(m_{\phi})\frac{\omega_{\phi}^{n}}{n!}-\int_{X}\log(\mathrm{v}(m_{\omega}))\mathrm{v}(m_{\omega})\frac{\omega^{n}}{n!}$
$-{\mathbb{E}}_{\mathrm{v}}^{\mathrm{Ric}(\omega)}(\phi)+{\mathbb{E}}_{\mathrm{w}\cdot\ell_{\mathrm{ext}}}(\phi).$
We will denote
${\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}}}(\phi)={\mathbb{H}}_{\mathrm{v}}(\phi)-{\mathbb{E}}_{\mathrm{v}}^{\mathrm{Ric}(\omega)}(\phi)+{\mathbb{E}}_{\mathrm{w}\cdot\ell_{\mathrm{ext}}}(\phi)-C_{0},$
where
${\mathbb{H}}_{\mathrm{v}}(\phi)=\int_{X}\log(\frac{\mathrm{v}(m_{\phi})\omega_{\phi}^{n}}{\omega^{n}})\mathrm{v}(m_{\phi})\frac{\omega_{\phi}^{n}}{n!},\;C_{0}=\int_{X}\log(\mathrm{v}(m_{\omega}))\mathrm{v}(m_{\omega})\frac{\omega^{n}}{n!}.$
In [Che00, Dar15, BN14, BB17, BDL17, Lah19, HL23, Lah23], the authors show that the space ${\mathcal{E}}^{1}_{K}(X,\omega)$ is granted with a Finsler structure. The functionals have controlled behaviours with respect to such Finsler structure.
Proposition 2.6.
Let $\phi_{0},\phi_{1}$ be two $\omega$-psh functions in ${\mathcal{E}}^{1}_{K}(X,\omega)$. Then there exists a unique geodesic segment $\phi_{t}$ $(0\leq t\leq 1)$ in ${\mathcal{E}}^{1}_{K}(X,\omega)$ that connects $\phi_{0}$ and $\phi_{1}$. The space ${\mathcal{E}}^{1}_{K}(X,\omega)$ is geodesically complete and associated with a distance $\mathrm{d}_{1}$. In addition, along any geodesic, ${\mathbb{E}}_{\mathrm{v}}$ is affine, ${\mathbb{J}}_{\mathrm{v}}$, ${\mathbb{M}}_{\mathrm{v},\mathrm{w}}$ are convex.
Remark 2.7.
Since $\mathrm{v}$ has positive lower and upper bounds, the strong topology defined by ${\mathbb{E}}_{\mathrm{v}}$ is equivalent to the strong topology defined by ${\mathbb{E}}$. And the entropies ${\mathbb{H}}$, ${\mathbb{H}}_{\mathrm{v}}$ are comparable, i.e., there exists $C>0$, such that $\frac{1}{C}({\mathbb{H}}+1)\leq{\mathbb{H}}_{\mathrm{v}}\leq C({\mathbb{H}}+1)$.
Definition 2.8 (Coercivity).
Let ${\mathbb{F}}$ be a functional defined over the potential space ${\mathcal{E}}^{1}_{K}(X,\omega)$. We say that ${\mathbb{F}}$ is coercive if there exist constants $\delta>0,C$, such that for any $\phi\in{\mathcal{E}}^{1}_{K}(X,\omega)$,
${\mathbb{F}}(\phi)\geq\delta\cdot{\mathbb{J}}_{\mathrm{v}}(\phi)-C.$
We say that ${\mathbb{F}}$ is $\mathbb{G}$-coercive if there exist constants $\delta>0,C$, such that for any $\phi\in{\mathcal{E}}^{1}_{K}(X,\omega)$,
${\mathbb{F}}(\phi)\geq\delta\cdot{\mathbb{J}}_{\mathrm{v},{\mathbb{T}}}(\phi)-C.$
Remark 2.9.
All the definitions and propositions of this subsection can be generalized to the setting of projective klt varieties. See for example [HL23, Section 2].
2.B. Weighted-cscK equations
In this subsection, following the idea in [CC21a], we re-write weighted-cscK equation (1.1) to a coupled PDE system, in which the first one is a $g$-Monge-Ampère equation (as studied in [BN14], [HL23]) and the second one is a linear elliptic equation. Compared with the cscK equation studied in [CC21a], one main difference is that, there are first order derivative terms appearing in the second equation. In the following of this subsection, we will collect some computations that are frequently used in latter sections.
Let
$F=\log\left(\frac{\mathrm{v}(m_{\phi})\omega_{\phi}^{n}}{\omega^{n}}\right)\quad i.e.,\mathrm{v}(m_{\phi})\omega_{\phi}^{n}=e^{F}\omega^{n}.$
Then the linearization of L.H.S. is given by
$\left.\frac{\mathrm{d}}{\mathrm{d}t}\right|_{t=0}(\mathrm{v}(m_{\phi_{t}})\omega_{\phi_{t}}^{n})=$ $\mathrm{v}_{,\alpha}(m_{\phi})\left.\frac{\mathrm{d}}{\mathrm{d}t}\right|_{t=0}\langle m_{\phi+t\psi},\xi^{\alpha}\rangle\omega_{\phi}^{n}+\mathrm{v}(m_{\phi})\Delta_{\phi}\psi\omega_{\phi}^{n}$
$=$ $\left(\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}\mathrm{d}^{\rm c}\psi(\xi^{\alpha})+\Delta_{\phi}(\psi)\right)\mathrm{v}(m_{\phi})\omega_{\phi}^{n}$
$=$ $\left(\Delta_{\phi}-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}\right)(\psi)\mathrm{v}(m_{\phi})\omega_{\phi}^{n}$
where $\phi_{t}=\phi+t\psi$ and $\Delta_{\phi}:=\mathrm{tr}_{\phi}\sqrt{-1}\partial\overline{\partial}$, we have used [Lah19, Lemma 1]
$m_{\phi}^{\xi}=m_{\omega}^{\xi}+t\mathrm{d}^{\rm c}\phi(\xi).$
Note that
$|\nabla^{\phi}\log\mathrm{v}(m_{\phi})|_{\phi}^{2}=\frac{\mathrm{v}_{,\alpha}(m_{\phi})\mathrm{v}_{,\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})^{2}}\langle\xi^{\beta},\xi^{\alpha}\rangle_{\phi}.$ (2.9)
We denote $m_{\phi}^{\xi}=\langle m_{\phi},\xi\rangle$.
One has
$\Delta_{\phi}\log\mathrm{v}(m_{\phi})=\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}\Delta_{\phi}(m^{\xi^{\alpha}}_{\phi})+\frac{\mathrm{v}_{,\alpha\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}-\frac{\mathrm{v}_{,\alpha}(m_{\phi})\mathrm{v}_{,\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})^{2}}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}$ (2.10)
and
$-J\xi^{\alpha}\log\left(\frac{\omega_{\phi}^{n}}{\omega^{n}}\right)=\mathrm{d}^{\rm c}\left(\log\left(\frac{\omega_{\phi}^{n}}{\omega^{n}}\right)\right)(\xi^{\alpha})=\Delta_{\phi}(m^{\xi^{\alpha}}_{\phi})-\Delta_{\omega}(m^{\xi^{\alpha}}_{\omega}),$ (2.11)
(see [Lah19, Lemma 5]). Thus one obtains
$-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}(F)=$ $\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}\mathrm{d}^{c}F(\xi^{\alpha})$
$=$ $\frac{\mathrm{v}_{,\alpha}(m_{\phi})\mathrm{v}_{,\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})^{2}}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}+\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}\Delta_{\phi}(m^{\xi^{\alpha}}_{\phi})-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}\Delta_{\omega}(m^{\xi^{\alpha}}_{\omega})$ (2.12)
Then we have
$\left(\Delta_{\phi}-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}\right)F=$ $\Delta_{\phi}\log\mathrm{v}(m_{\phi})+\Delta_{\phi}\log\left(\frac{\omega_{\phi}^{n}}{\omega^{n}}\right)-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}(F)$
$=$ $2\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}\Delta_{\phi}(m^{\xi^{\alpha}}_{\phi})+\frac{\mathrm{v}_{,\alpha\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}\Delta_{\omega}(m^{\xi^{\alpha}}_{\omega})$
$-S(\phi)+\mathrm{tr}_{\phi}\mathrm{Ric}(\omega)$
$=$ $-\frac{S_{\mathrm{v}}(\phi)}{\mathrm{v}(m_{\phi})}-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}\Delta_{\omega}(m^{\xi^{\alpha}}_{\omega})+\mathrm{tr}_{\phi}\mathrm{Ric}(\omega)$
$=$ $-\frac{S_{\mathrm{v}}(\phi)}{\mathrm{v}(m_{\phi})}+\mathrm{v}(m_{\phi})^{-1}\Big{(}\langle(\mathrm{d}\mathrm{v})(m_{\phi}),m_{\mathrm{Ric}(\omega)}\rangle+\mathrm{v}(m_{\phi})\mathrm{tr}_{\phi}\mathrm{Ric}(\omega)\Big{)}$
$=$ $-\frac{S_{\mathrm{v}}(\phi)}{\mathrm{v}(m_{\phi})}+\frac{\mathrm{tr}_{\mathrm{v},\phi}\mathrm{Ric}(\omega)}{\mathrm{v}(m_{\phi})}$
where
$S_{\mathrm{v}}(\phi)=\mathrm{v}(m_{\phi})S(\phi)-2\mathrm{v}_{,\alpha}\Delta_{\phi}(m_{\phi}^{\xi^{\alpha}})-\mathrm{v}_{,\alpha\beta}(m_{\phi})\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}.$
and $m_{\mathrm{Ric}(\omega)}:=\Delta_{\omega}(m_{\omega})$.
The continuity path (1.3) can be reformulated into the following coupled PDE system
$\mathrm{v}(m_{\phi})\omega_{\phi}^{n}=e^{F}\omega^{n},$ (2.13a)
$\left(\Delta_{\phi}-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}\right)F=-\ell_{\mathrm{ext}}(m_{\phi})\frac{\mathrm{w}(m_{\phi})}{\mathrm{v}(m_{\phi})}+\frac{\mathrm{tr}_{\mathrm{v},\phi}(\eta_{t})}{\mathrm{v}(m_{\phi})}-(1-\frac{1}{t})\underline{\theta}\frac{\mathrm{w}(m_{\phi})}{\mathrm{v}(m_{\phi})},$ (2.13b)
where we denote $\eta_{t}:=(1-\frac{1}{t})\theta+\mathrm{Ric}(\omega)$. By (2.12) and (2.13b),
$\Delta_{\phi}F=$ $-\ell_{\mathrm{ext}}(m_{\phi})\frac{\mathrm{w}(m_{\phi})}{\mathrm{v}(m_{\phi})}+\frac{\mathrm{tr}_{\mathrm{v},\phi}(\eta_{t})}{\mathrm{v}(m_{\phi})}-(1-\frac{1}{t})\underline{\theta}\frac{\mathrm{w}(m_{\phi})}{\mathrm{v}(m_{\phi})}$
$-\frac{\mathrm{v}_{,\alpha}(m_{\phi})\mathrm{v}_{,\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})^{2}}\langle\xi^{\beta},\xi^{\alpha}\rangle_{\phi}-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}\Delta_{\phi}(m_{\phi}^{\xi^{\alpha}}-m_{\omega}^{\xi^{\alpha}})$
and
$\Delta_{\phi}(F-\log\mathrm{v}(m_{\phi}))=$ $-\ell_{\mathrm{ext}}(m_{\phi})\frac{\mathrm{w}(m_{\phi})}{\mathrm{v}(m_{\phi})}+\frac{\mathrm{tr}_{\mathrm{v},\phi}(\eta_{t})}{\mathrm{v}(m_{\phi})}-(1-\frac{1}{t})\underline{\theta}\frac{\mathrm{w}(m_{\phi})}{\mathrm{v}(m_{\phi})}$
$-\frac{\mathrm{v}_{,\alpha}(m_{\phi})\mathrm{v}_{,\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})^{2}}\langle\xi^{\beta},\xi^{\alpha}\rangle_{\phi}-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}\Delta_{\phi}m_{\phi}^{\xi^{\alpha}}+\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}\Delta_{\omega}m_{\omega}^{\xi^{\alpha}}$
$-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}\Delta_{\phi}(m^{\xi^{\alpha}}_{\phi})-\frac{\mathrm{v}_{,\alpha\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}+\frac{\mathrm{v}_{,\alpha}(m_{\phi})\mathrm{v}_{,\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})^{2}}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}$
$=$ $-\left((\ell_{\mathrm{ext}}+(1-\frac{1}{t})\underline{\theta})\frac{\mathrm{w}}{\mathrm{v}}-(1-\frac{1}{t})\frac{\langle(\mathrm{d}\mathrm{v})(m_{\phi}),m_{\theta}\rangle}{\mathrm{v}}\right)+\mathrm{tr}_{\phi}\eta_{t}$
$-2\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}\Delta_{\phi}m_{\phi}^{\xi^{\alpha}}-\frac{\mathrm{v}_{,\alpha\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}.$
Note that, by (2.12),
$\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}(F-\log\mathrm{v}(m_{\phi}))=$ $\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}(F)-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}(\log\mathrm{v}(m_{\phi}))$
$=$ $\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}(F)-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}\frac{J\xi^{\alpha}(\mathrm{v}(m_{\phi}))}{\mathrm{v}(m_{\phi})}$
$=$ $\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}(F)-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})^{2}}\mathrm{v}_{,\beta}(m_{\phi})J\xi^{\alpha}(\langle m_{\phi},\xi^{\beta}\rangle)$
$=$ $\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}(F)+\frac{\mathrm{v}_{,\alpha}(m_{\phi})\mathrm{v}_{,\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})^{2}}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}$
$=$ $-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}\Delta_{\phi}(m_{\phi}^{\xi^{\alpha}})+\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}\Delta_{\omega}(m_{\omega}^{\xi^{\alpha}}).$ (2.14)
It follows that
$\Delta_{\phi}(F-\log\mathrm{v}(m_{\phi}))=$ $-f_{t}^{\prime}+\mathrm{tr}_{\phi}\eta_{t}+2\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}(F-\log\mathrm{v}(m_{\phi}))-\frac{\mathrm{v}_{,\alpha\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi},$ (2.15)
where
$f_{t}^{\prime}:=(\ell_{\mathrm{ext}}+(1-\frac{1}{t})\underline{\theta})\frac{\mathrm{w}}{\mathrm{v}}-(1-\frac{1}{t})\frac{\langle(\mathrm{d}\mathrm{v})(m_{\phi}),m_{\theta}\rangle}{\mathrm{v}}+2\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}\Delta_{\omega}(m_{\omega}^{\xi^{\alpha}}).$
In addition, by (2.10) and (2.14)
$\Delta_{\phi}\log\mathrm{v}(m_{\phi})=$ $-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}(F-\log\mathrm{v}(m_{\phi}))+\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}\Delta_{\omega}(m_{\omega}^{\xi^{\alpha}}).$
$+\frac{\mathrm{v}_{,\alpha\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}-\frac{\mathrm{v}_{,\alpha}(m_{\phi})\mathrm{v}_{,\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})^{2}}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}$ (2.16)
One also has
$\Delta_{\phi}(F+\log\mathrm{v}(m_{\phi}))=$ $-f_{t}+\mathrm{tr}_{\phi}\eta_{t}-2\frac{\mathrm{v}_{,\alpha}(m_{\phi})\mathrm{v}_{,\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})^{2}}\langle\xi^{\beta},\xi^{\alpha}\rangle_{\phi}+\frac{\mathrm{v}_{,\alpha\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi},$ (2.17)
where
$f_{t}=(\ell_{\mathrm{ext}}+(1-\frac{1}{t})\underline{\theta})\frac{\mathrm{w}}{\mathrm{v}}-(1-\frac{1}{t})\frac{\langle(\mathrm{d}\mathrm{v})(m_{\phi}),m_{\theta}\rangle}{\mathrm{v}}.$
We have
$\left|2\frac{\mathrm{v}_{,\alpha}(m_{\phi})\mathrm{v}_{,\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})^{2}}-\frac{\mathrm{v}_{,\alpha\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})}\right|\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}\leq A_{\mathrm{v},\xi}^{\prime}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}\leq A_{\mathrm{v},\xi}\mathrm{tr}\omega_{\phi},$ (2.18)
for some constant $A_{\mathrm{v},\xi}>0$ only depends on $\mathrm{v}$ and $\{\xi^{\alpha}\}$.
In addition, if $\mathrm{v}$ is log-concave, then
$\left(\frac{\mathrm{v}_{,\alpha\beta}}{\mathrm{v}}-\frac{\mathrm{v}_{,\alpha}\mathrm{v}_{,\beta}}{\mathrm{v}^{2}}\right)\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}\leq 0.$ (2.19)
Indeed, for any $p\in X$, let $\{X^{i}\}$ be the orthnormal basis of $T_{p}X$ with respect to $g_{\phi}$ and write $\xi^{\alpha}=\xi^{\alpha}_{i}X^{i}$. Thus, one has
$\left(\frac{\mathrm{v}_{,\alpha\beta}}{\mathrm{v}}-\frac{\mathrm{v}_{,\alpha}\mathrm{v}_{,\beta}}{\mathrm{v}^{2}}\right)\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}=\sum_{i=1}^{n}\left(\frac{\mathrm{v}_{,\alpha\beta}}{\mathrm{v}}-\frac{\mathrm{v}_{,\alpha}\mathrm{v}_{,\beta}}{\mathrm{v}^{2}}\right)\xi^{\alpha}_{i}\xi^{\beta}_{\overline{i}}.$
Since $\mathrm{v}$ is log-concave, then $\left(\frac{\mathrm{v}_{,\alpha\beta}}{\mathrm{v}}-\frac{\mathrm{v}_{,\alpha}\mathrm{v}_{,\beta}}{\mathrm{v}^{2}}\right)_{\alpha,\beta}$ is a negative definite matrix. It follows that
$\left(\frac{\mathrm{v}_{,\alpha\beta}}{\mathrm{v}}-\frac{\mathrm{v}_{,\alpha}\mathrm{v}_{,\beta}}{\mathrm{v}^{2}}\right)\xi^{\alpha}_{i}\xi^{\beta}_{\overline{i}}=(\xi^{1}_{i},\ldots,\xi^{r}_{i})\cdot\left(\frac{\mathrm{v}_{,\alpha\beta}}{\mathrm{v}}-\frac{\mathrm{v}_{,\alpha}\mathrm{v}_{,\beta}}{\mathrm{v}^{2}}\right)\cdot(\xi^{1}_{\overline{i}},\ldots,\xi^{r}_{\overline{i}})^{T}\leq 0$
for each fixed $i$. Therefore, (2.19) follows.
In the study of the closedness part of the continuity method, it is also necessary to study the following modified (by automorphism) coupled equations (see Lemma 4.7):
$\mathrm{v}(m_{\phi})\omega_{\phi}^{n}=e^{F}\omega^{n},$ (2.20a)
$\left(\Delta_{\phi}-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}\right)F=-\ell_{\mathrm{ext}}(m_{\phi})\frac{\mathrm{w}(m_{\phi})}{\mathrm{v}(m_{\phi})}+(1-\frac{1}{t})\frac{\mathrm{tr}_{\mathrm{v},\phi}(\theta)}{\mathrm{v}(m_{\phi})}$
$\qquad\qquad\qquad\qquad\qquad\quad-\frac{\mathrm{tr}_{\mathrm{v},\phi}({\rm dd^{c}}f_{*})}{\mathrm{v}(m_{\phi})}-(1-\frac{1}{t})\underline{\theta}\frac{\mathrm{w}(m_{\phi})}{\mathrm{v}(m_{\phi})}+\frac{\mathrm{tr}_{\mathrm{v},\phi}\mathrm{Ric}(\omega)}{\mathrm{v}(m_{\phi})},$ (2.20b)
where $f_{*}\in\mathrm{Psh}(X,(\frac{1-t}{t})\theta)$, $\sup_{X}f_{*}=0$.
2.C. Examples
The examples in this subsection all satisfy the log-concavity assumption of $\mathrm{v}$. By choosing the functions $\mathrm{v},\mathrm{w}$ accordingly, the expression of the weighted-cscK metric can be formulated into Kähler–Einstein metric, cscK metric, Kähler–Ricci soliton, extremal metrics. We refer the interested readers to the examples listed in[Lah19, Section 3].
Among others, one interesting example is studied in [Ino22], where the author defined the constant $\mu_{\xi}^{\lambda}$-scalar curvature Kähler metric ($\mu$-cscK for short) as a generalization of cscK metric. The $\mu_{\xi}^{\lambda}$-scalar curvature is defined by
$S_{\xi}^{\lambda}(\omega_{\phi})=S(\omega_{\phi})-2\Delta_{\phi}m_{\phi}^{\xi}-J\xi(m_{\phi}^{\xi})-\lambda m_{\phi}^{\xi},$
where $J\xi+\sqrt{-1}\xi$ is a holomorphic vector field, $m_{\phi}^{\xi}$ is the corresponding Hamiltonian function of $\xi$, $\lambda\in{\mathbb{R}}$. The $\mu$-cscK metric can be expressed as a weighted-cscK metric by letting $\mathrm{v}=\mathrm{w}=e^{m_{\phi}^{\xi}}$. When $X$ is a Fano manifold, one interesting feature of the $\mu$-cscK metric $\omega_{\phi}$ is that, it interpolates the extremal metric and the Kähler–Ricci soliton metric under the assumption of the existence. In particular, as $\lambda\to-\infty$, $\omega_{\phi}$ converges to the extremal metric; as $\lambda$ converges to the average scalar curvature $\underline{S}>0$, $\omega_{\phi}$ converges to the Kähler–Ricci soliton.
Another interesting feature is, when the $\mu$-cscK metric exists, the vector field $\xi$ is a critical point of the volume functional $\mathrm{Vol}^{\lambda}(\xi)$, where
$\log(\mathrm{Vol}^{\lambda})(\xi)=\frac{1}{\int_{X}e^{m_{\phi}^{\xi}}\omega_{\phi}^{n}}\int_{X}\Big{(}S(\omega_{\phi})-\Delta_{\phi}m_{\phi}^{\xi}-\lambda m_{\phi}^{\xi}\Big{)}e^{m_{\phi}^{\xi}}\omega_{\phi}^{n}+\lambda\cdot\log(\int_{X}e^{m_{\phi}^{\xi}}\omega_{\phi}^{n}).$
After a normalization, $\log(\mathrm{Vol}^{\lambda})$ has the same expression as Perelman’s $W$-functional [Per02]. When restricted on Fano manifolds, $\log(\mathrm{Vol}^{\lambda})$ is equal to Tian–Zhu’s volume functional [TZ02] up to the difference of a constant.
Some other related works can be found in [He19, Ino21, Liu22, SD22, Jub23, Zha23, Fut24, GJSS24, His24, NN24].
3. Apriori estimates
In this section, we provide uniform $C^{0}$-estimate, $W^{2,p}$-estimate of $\phi$ and gradient estimate of $F+f_{*}-\log(\mathrm{v}(m_{\phi}))$ for equation (2.20a), (2.20b). The proof of the uniform $C^{0}$-estimate is based on a modification of Guo–Phong’s method [GP22]. The proof of $W^{2,p}$-estimate and gradient estimate follows from Chen–Cheng’s method [CC21b] with a new choice of test functions and integral weights.
3.A. Uniform $C^{0}$-estimate
Let $\omega_{X}$ be a fixed smooth Kähler form, and let $\omega=\chi+s\omega_{X}$ be a smooth Kähler form, where $\chi$ is a smooth $(1,1)$-form in a nef class, $s>0$. We will first consider the uniform $C^{0}$-estimate of the following PDE system
$\begin{cases}\mathrm{v}(m_{\phi})\;\omega_{\phi}^{n}=c_{\omega,\mathrm{v}}\;e^{\widetilde{F}}\omega_{X}^{n},\\ (\Delta_{\phi}-\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha})\big{(}\widetilde{F}+f_{*}\big{)}=-f+\mathrm{tr}_{\phi}\eta,\end{cases}$ (3.1)
where $c_{\omega,\mathrm{v}}=\int_{X}\mathrm{v}\;\omega^{n}/{\int_{X}\omega_{X}^{n}}$. We assume that $f$ is a smooth function on $X$. Furthermore, we also assume that there exists a constant $p>1$ such that $\int_{X}e^{-pf_{*}}\omega^{n}<C(p)$.
Remark 3.1.
In the modified continuity path, for the $(\frac{1-t}{t})\theta$-psh function $f_{*}$, for arbitrarily large $p$, we have $\int_{X}e^{-pf_{*}}\omega^{n}$ uniformly bounded from above when $t$ is sufficiently close to $1$ (see Proposition 4.8). Furthermore, the function $\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha}(f_{*})$ is determined by the class $[\theta]$, which is bounded. Therefore the modified continuity path (2.20a), (2.20b) can be reformulated into (3.1) when $t$ is sufficiently close to $1$.
Theorem 3.2.
Assume $\omega\leq\kappa\omega_{X}$ for some $\kappa>0$ and $\int_{X}\widetilde{F}e^{\widetilde{F}}\omega_{X}^{n}\leq K_{1}$, $\eta\geq-K_{2}\omega$, $\int_{X}e^{-pf_{*}}\omega^{n}<C(p)$ for some $p>1$, $\sup_{X}f_{*}=0$, $\sup_{X}\phi=0$. Then $\sup_{X}|\phi|\leq C$, $\sup_{X}(\widetilde{F}+f_{*})\leq C$, where the constant $C$ depends on $\omega_{X},n,\kappa,f,\mathrm{v},C(p),K_{1},K_{2}$.
If we further assume $\eta\leq K_{3}\omega$, then $\widetilde{F}+f_{*}$ is bounded from below by another constant depending further on $K_{3}$.
Remark 3.3.
The format of the statement of Theorem 3.2 is to make it suitable for our future application in singular cases.
Proof.
By applying [GP22, Theorem 2] to equation $\omega_{\phi}^{n}=c_{\omega}e^{\widetilde{F}}\omega_{X}^{n}$, for $l>n$, we have $\sup_{X}|\phi|\leq C$, where $C$ depends on $\omega_{X},n,l,\kappa,{\rm Ent}_{l}(e^{\widetilde{F}})$.
Therefore, we just need to modify [GP22, Theorem 3] to use $\int_{X}\widetilde{F}e^{\widetilde{F}}\omega_{X}^{n}$ to bound ${\rm Ent}_{l}(\widetilde{F})$.
We need to use the auxiliary function $\psi_{k}$ which satisfies the following equation
$(\omega+{\rm dd}^{c}\psi_{k})^{n}=\frac{\tau_{k}\Big{(}-\phi+\delta(\widetilde{F}+f_{*})\Big{)}}{A_{k}}\cdot c_{\omega,\mathrm{v}}\cdot e^{\widetilde{F}}\omega_{X}^{n},\quad\sup_{X}\psi_{k}=0,$ (3.2)
where $\delta=\frac{1}{10}K_{2}^{-1}$, $\tau_{k}:{\mathbb{R}}\rightarrow{\mathbb{R}}_{+}$ is a sequence of positive smooth functions that converges monotonically decreasingly to the function $x\cdot{\rm Id}_{{\mathbb{R}}_{+}}(x)$. And
$A_{k}=\frac{c_{\omega,\mathrm{v}}}{V_{\omega}}\int_{X}\tau_{k}\Big{(}-\phi+\delta(\widetilde{F}+f_{*})\Big{)}\cdot e^{\widetilde{F}}\omega_{X}^{n},$
which converges to
$A_{\infty}$ $=\frac{c_{\omega,\mathrm{v}}}{V_{\omega}}\int_{\Omega}\Big{(}-\phi+\delta(\widetilde{F}+f_{*})\Big{)}e^{\widetilde{F}}\omega_{X}^{n}$
$\leq\frac{c_{\omega,\mathrm{v}}}{V_{\omega}}\int_{\Omega}\Big{(}-\phi+\delta\cdot\widetilde{F}\Big{)}e^{\widetilde{F}}\omega_{X}^{n},$
where $\Omega=\{x\in X:\widetilde{F}+f_{*}>0\}$ and $0\leq A_{\infty}\leq C(K_{1})$, which is due to the entropy bound and the weak compactness implied by it.
Consider
$\Psi=-\epsilon(-\psi_{k}+\Lambda)^{\frac{n}{n+1}}-\phi+\delta(\widetilde{F}+f_{*}),$ (3.3)
where $\epsilon=(\frac{(n+1)(n+C_{f,\delta})}{n^{2}})^{\frac{n}{n+1}}A_{k}^{\frac{1}{n+1}}$, $\Lambda^{\frac{1}{n+1}}=\frac{2n}{n+1}\epsilon$, where the constant $C_{f,\delta}$ is chosen such that $\left|-\delta f+\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha}(\phi)\right|\leq C_{f,\delta}$.
The next is to show that $\Psi\leq 0$. It suffices to assume $\sup_{X}\Psi$ is achieved at a point $x_{0}\in\Omega$. We have
$\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha}(\Psi)(x_{0})=0,$
which implies
$\delta\cdot\Delta_{\phi}(\widetilde{F}+f_{*})=-\delta\cdot f+\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha}(\phi)+\delta\cdot\mathrm{tr}_{\phi}\eta-\frac{n\epsilon}{n+1}\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha}(\psi_{k})\cdot(-\psi_{k}+\Lambda)^{-\frac{1}{n+1}}.$
Then at $x_{0}$, we have
00 0 $\geq\Delta_{\phi}\Psi$
$\geq\frac{\epsilon n}{n+1}(-\psi_{k}+\Lambda)^{\frac{-1}{n+1}}({\rm tr}_{\phi}\omega_{\psi_{k}}-{\rm tr}_{\phi}\omega)-(n-{\rm tr}_{\phi}\omega)+\delta\Delta_{\phi}(\widetilde{F}+f_{*})$
$\geq\frac{\epsilon n^{2}}{n+1}(-\psi_{k}+\Lambda)^{\frac{-1}{n+1}}\Big{(}\frac{-\phi+\delta(\widetilde{F}+f_{*})}{A_{k}}\Big{)}^{\frac{1}{n}}-n-C_{f,\delta}$
$\quad+\left(1-\frac{\epsilon n}{n+1}(-\psi_{k}+\Lambda)^{\frac{-1}{n+1}}-\frac{1}{10}\right){\rm tr}_{\phi}\omega-\frac{n\epsilon}{n+1}\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha}(\psi_{k})(-\psi_{k}+\Lambda)^{\frac{-1}{n+1}}$
$\geq\frac{\epsilon n^{2}}{n+1}(-\psi_{k}+\Lambda)^{\frac{-1}{n+1}}\left(\Big{(}\frac{-\phi+\delta(\widetilde{F}+f_{*})}{A_{k}}\Big{)}^{\frac{1}{n}}-\frac{\mathrm{v}_{\alpha}}{n\mathrm{v}}J\xi^{\alpha}(\psi_{k})\right)-(n+C_{f,\delta}).$
Note that $|\frac{\mathrm{v}_{\alpha}}{n\mathrm{v}}J\xi^{\alpha}(\psi_{k})|$ is bounded by some constant $C^{\prime}$ uniformly. This implies
$-\phi+\delta(\widetilde{F}+f_{*})\leq\left(\frac{(n+1)(n+C_{f,\delta})}{\epsilon n^{2}}\right)^{n}A_{k}(-\psi_{k}+\Lambda)^{\frac{n}{n+1}}+C^{\prime}=\epsilon(-\psi_{k}+\Lambda)^{\frac{n}{n+1}}+C^{\prime}.$
This concludes that $\sup_{X}\Psi=\Psi(x_{0})\leq C^{\prime}$. Therefore
$\delta(\widetilde{F}+f_{*})\leq-\phi+\delta(\widetilde{F}+f_{*})\leq C\big{(}(-\psi_{k}+1)^{\frac{n}{n+1}}+1)\leq-\widetilde{\epsilon}\psi_{k}+C_{\widetilde{\epsilon}},$
where $\widetilde{\epsilon}$ can be chosen arbitrarily small. Denote $q^{*}>1$ such that $\frac{1}{q}+\frac{1}{q^{*}}=1$. Choose $r\,\widetilde{\epsilon}$ such that $r\delta>1$ and $rq\widetilde{\epsilon}\leq\alpha(X,\kappa\omega_{X})$ for some $q^{*}>1$ that is close to $1$, and $p=q^{*}r\delta>1$. Then we have
$\int_{X}e^{r\delta\widetilde{F}}\omega_{X}^{n}$ $=\int_{X}e^{r\delta(\widetilde{F}+f_{*}-f_{*})}\omega_{X}^{n}$
$\leq\Big{(}\int_{X}e^{qr\delta(\widetilde{F}+f_{*})}\omega_{X}^{n}\Big{)}^{\frac{1}{q}}\cdot\Big{(}\int_{X}e^{-q^{*}r\delta f_{*}}\omega_{X}^{n}\Big{)}^{\frac{1}{q^{*}}}$
$\leq C\Big{(}\int_{X}e^{-qr\widetilde{\epsilon}\psi_{k}}\omega_{X}^{n}\Big{)}^{\frac{1}{q}}\leq C^{\prime}.$
In particular, this implies $\mathrm{Ent}_{l}(\widetilde{F})$ is uniformly bounded. Then by [GPT23, Theorem 1] we have $\sup_{X}|\phi|\leq C$, where the constant $C$ depends on parameters listed in the statement of the theorem. (Alternatively, we can also use the $C^{0}$-estimate in [EGZ09]).
In order to prove the upper and lower bounds of $\widetilde{F}+f_{*}$, we need to use the mean-value type inequality proved in [GP22].
Let $u:=\widetilde{F}+f_{*}-K_{2}\phi$, which satisfies $(\Delta_{\phi}-\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha})u\geq-a$ for some constant $a\geq 0$. Apply [GP22, Lemma 3] to $u$, we can obtain the upper bound of $\widetilde{F}+f_{*}$.
Similarly, let $u:=-(\widetilde{F}+f_{*})-K_{3}\phi$, which satisfies $(\Delta_{\phi}-\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha})u\geq-a$ for some $a\geq 0$, we can obtain a lower bound of $\widetilde{F}+f_{*}$. ∎
3.B. $W^{2,p}$-estimate
For the convenience of the computation, since
$\frac{\mathrm{tr}_{\mathrm{v},\phi}({\rm dd^{c}}f_{*})}{\mathrm{v}(m_{\phi})}=\Delta_{\phi}f_{*}+\frac{\langle(\mathrm{d}\mathrm{v})(m_{\phi}),\mathrm{d}^{c}f_{*}\rangle}{\mathrm{v}}=\Delta_{\phi}f_{*}-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}f_{*},$
we rewrite the coupled system (2.20a), (2.20b) into
$\mathrm{v}(m_{\phi})\omega_{\phi}^{n}$ $=e^{F}\omega^{n},$ (3.4)
$\left(\Delta_{\phi}-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}\right)F$ $=\widetilde{f}_{t}+\mathrm{tr}_{\phi}(\eta_{t})-\Delta_{\phi}f_{*}+\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha}f_{*},$
where $f_{*}\in\mathrm{Psh}(X,(\frac{1}{t}-1)\theta)$, $\sup_{X}f_{*}=0$,
$\widetilde{f}_{t}=-(\ell_{\mathrm{ext}}+(1-\frac{1}{t})\underline{\theta})\frac{\mathrm{w}}{\mathrm{v}}+(1-\frac{1}{t})\frac{\langle(\mathrm{d}\mathrm{v})(m_{\phi}),m_{\theta}\rangle}{\mathrm{v}}-\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}\Delta_{\omega}m_{\omega}^{\xi^{\alpha}}$
which is a bounded function, $\eta_{t}=(1-\frac{1}{t})\theta+\mathrm{Ric}(\omega)$, and $\int_{X}e^{-pf_{*}}\omega^{n}<C$ for a sufficiently large positive constant $p$. (which can be achieved when $t\to 1$ in the modified continuity path.)
Lemma 3.4.
Let $\phi$ be a solution of (3.4), then
$\Delta_{\phi}\log\mathrm{tr}\omega_{\phi}\geq\frac{\Delta(F-\log\mathrm{v}(m_{\phi}))-S(\omega)}{\mathrm{tr}\omega_{\phi}}-B\mathrm{tr}_{\phi}\omega,$
where $B>0$ is a constant depending on the lower bound of the holomorphic bisectional curvature of $\omega$.
Proof.
By [Szfrm[o]–4, Lemma 3.7], then
$\Delta_{\phi}\log\mathrm{tr}\omega_{\phi}\geq-B\mathrm{tr}_{\phi}\omega-\frac{g^{j\overline{k}}\mathrm{Ric}(\phi)_{j\overline{k}}}{\mathrm{tr}\omega_{\phi}},$
where $B>0$ is a constant depending on the lower bound of the holomorphic bisectional curvature of $\omega$. Since
$\mathrm{Ric}(\omega_{\phi})=\mathrm{Ric}(\omega)-\sqrt{-1}\partial\overline{\partial}\log\frac{\omega_{\phi}^{n}}{\omega^{n}}=\mathrm{Ric}(\omega)-\sqrt{-1}\partial\overline{\partial}(F-\log\mathrm{v}(m_{\phi})),$
then
$g^{j\overline{k}}\mathrm{Ric}(\phi)_{j\overline{k}}=S(\omega)-\Delta(F-\log\mathrm{v}(m_{\phi})).$
Therefore the lemma follows. ∎
Theorem 3.5.
Assume $\mathrm{v}$ is log-concave. Let $\phi$ be a solution of (3.4), for any $p\geq 1$, then
$\int_{X}e^{(p-1)f_{*}}(n+\Delta\phi)^{p}\omega^{n}\leq C,$ (3.5)
where $C$ is a constant depending on $p$, $\mathrm{v}$, $\mathrm{w}$, $\|\phi\|_{C^{0}}$, $\|F+f_{*}\|_{C^{0}}$, an upper bound of $\theta$ and $\mathrm{Ric}(\omega)$, and a lower bound of the holomorphic bisectional curvature of $\omega$.
Remark 3.6.
By Hölder inequality, we also have
$\int_{X}(n+\Delta\phi)^{p}\omega^{n}\leq C,$ (3.6)
for a constant $C$ that depends on the same parameters as stated above.
Proof.
We consider
$u:=e^{-a(F+\log\mathrm{v}(m_{\phi})+\delta f_{*}+b\phi)}\mathrm{tr}\omega_{\phi},$
where $a,b>1$ are uniform constants determined later. Since
$\sqrt{-1}\partial\overline{\partial}e^{f}=e^{f}\sqrt{-1}\partial\overline{\partial}f+e^{f}\sqrt{-1}\partial f\wedge\overline{\partial}f,$
then one has
$\Delta_{\phi}u=\Delta_{\phi}e^{\log u}=$ $e^{\log u}\Delta_{\phi}\log u+u\nabla^{\phi}(\log u)\cdot_{\phi}\nabla^{\phi}(\log u)$
$=$ $-au\Delta_{\phi}(F+\log\mathrm{v}(m_{\phi})+\delta f_{*}+b\phi)+u\Delta_{\phi}\log\mathrm{tr}\omega_{\phi}+u^{-1}|\nabla^{\phi}u|_{\phi}^{2}.$ (3.7)
By (2.17), (3.4) and Remark 3.1, we have
$\Delta_{\phi}(F+\log\mathrm{v}(m_{\phi})+\delta f_{*}+b\phi)$
$=-f_{t}+\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha}f_{*}+\mathrm{tr}_{\phi}\eta_{t}-2|\nabla^{\phi}\log\mathrm{v}(m_{\phi})|_{\phi}^{2}+\frac{\mathrm{v}_{,\alpha\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}+bn-b\mathrm{tr}_{\phi}\omega-(1-\delta)\Delta_{\phi}f_{*}$
$\leq(C+bn-f_{t})+(A_{1}-b)\mathrm{tr}_{\phi}\omega-2|\nabla^{\phi}\log\mathrm{v}(m_{\phi})|_{\phi}^{2}+\frac{\mathrm{v}_{,\alpha\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}-(1-\delta)\Delta_{\phi}f_{*},$ (3.8)
where $A_{1}>0$ is a constant satisfying $\eta_{t}\leq A_{1}\omega$. By Lemma 3.4, (3.7) and (3.8),
$\Delta_{\phi}u$ $\geq e^{-a(F+\log\mathrm{v}+\delta f_{*}+b\phi)}\Big{(}a(f_{t}-bn-C+(1-\delta)\Delta_{\phi}f_{*})\mathrm{tr}\omega_{\phi}+(ab-aA_{1}-B)\mathrm{tr}_{\phi}\omega\mathrm{tr}\omega_{\phi}$
$\qquad\quad+\Delta F^{\prime}-S(\omega)\Big{)}+2au|\nabla^{\phi}\log\mathrm{v}(m_{\phi})|_{\phi}^{2}-au\frac{\mathrm{v}_{,\alpha\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}+u^{-1}|\nabla^{\phi}u|_{\phi}^{2}\ ,$ (3.9)
where we denote $F^{\prime}:=F-\log\mathrm{v}(m_{\phi})$.
Note that
$\frac{1}{2p+1}\Delta_{\phi}u^{2p+1}=$ $u^{2p}\Delta_{\phi}u+2pu^{2p-1}|\nabla^{\phi}u|_{\phi}^{2}.$
Take the integral on the above equality with respect to the weighted measure $e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$, together with (3.9), we obtain
$\int_{X}2pu^{2p-1}\cdot|\nabla_{\phi}u|_{\phi}^{2}\cdot e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
$=$ $\int_{X}u^{2p}\Big{(}-\Delta_{\phi}u-\langle\nabla_{\phi}u,2a\nabla_{\phi}\log\mathrm{v}\rangle_{\phi}\Big{)}\cdot e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
$\leq$ $-2a\int_{X}u^{2p}\langle\nabla_{\phi}u,\nabla_{\phi}\log\mathrm{v}\rangle_{\phi}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
$-\int_{X}u^{2p}e^{-a(F+\log\mathrm{v}+\delta f_{*}+b\phi)}\Bigl{\{}a\Big{(}f_{t}-bn-C+(1-\delta)\Delta_{\phi}f_{*}\Big{)}\mathrm{tr}\omega_{\phi}+(ab-aA_{1}-B)\mathrm{tr}_{\phi}\omega\mathrm{tr}\omega_{\phi}$
$+\Delta(F-\log\mathrm{v})-S(\omega)\Bigr{\}}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
$-2a\int_{X}u^{2p+1}|\nabla_{\phi}\log\mathrm{v}|_{\phi}^{2}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}-\int_{X}u^{2p-1}|\nabla_{\phi}u|_{\phi}^{2}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
$+a\int_{X}u^{2p+1}\frac{\mathrm{v}_{,\alpha\beta}}{\mathrm{v}}\langle\xi_{\alpha},\xi_{\beta}\rangle_{\phi}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}.$
Using the fact that $|S(\omega)|<C$, and by choosing $b>1$ sufficiently large so that $ab-aA_{1}-B\geq\frac{ab}{2}$ and $|a(f_{t}-bn-C)|\leq(n+1)ab$, the inequality above can be simplified into
$(2p+1-a)\int_{X}u^{2p-1}\cdot|\nabla_{\phi}u|_{\phi}^{2}\cdot e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
$\leq$ $(n+1)ab\int_{X}u^{2p+1}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}-a(1-\delta)\int_{X}u^{2p+1}\Delta_{\phi}f_{*}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
$-\frac{ab}{2}\int_{X}u^{2p+1}\mathrm{tr}_{\phi}\omega e^{2a\log\mathrm{v}}\omega_{\phi}^{n}+C\int_{X}u^{2p}e^{-a(F+\log\mathrm{v}+\delta f_{*}+b\phi)}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
$-\int_{X}u^{2p}e^{-a(F+\log\mathrm{v}+\delta f_{*}+b\phi)}\Delta(F-\log\mathrm{v})e^{2a\log\mathrm{v}}\omega_{\phi}^{n},$ (3.10)
where we use the fact (2.19) which is ensured by the log-concavity of $\mathrm{v}$, and
$-2a\int_{X}u^{2p}\langle\nabla_{\phi}u,\nabla_{\phi}\log\mathrm{v}\rangle_{\phi}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
$\leq$ ${a}\int_{X}u^{2p+1}|\nabla_{\phi}\log\mathrm{v}|_{\phi}^{2}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}+a\int_{X}u^{2p-1}|\nabla_{\phi}u|_{\phi}^{2}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}.$
Next, we handle the term involving $\Delta(F-\log\mathrm{v})$ in the above inequality by repeating the integration by parts. Recall that $\omega_{\phi}^{n}=e^{F-\log\mathrm{v}}\omega^{n}$.
$-\int_{X}u^{2p}e^{-a(F+\log\mathrm{v}+\delta f_{*}+b\phi)}\Delta(F-\log\mathrm{v})e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
$=$ $-\int_{X}u^{2p}e^{(1-a)(F-\log\mathrm{v})-a\delta f_{*}-ab\phi}\frac{1}{1-a}\Delta\Big{(}(1-a)(F-\log\mathrm{v})\Big{)}\omega^{n}$
$=$ $-\int_{X}u^{2p}e^{(1-a)(F-\log\mathrm{v})-a\delta f_{*}-ab\phi}\cdot\frac{1}{1-a}\Delta\Big{(}(1-a)(F-\log\mathrm{v})-a\delta f_{*}-ab\phi\Big{)}\omega^{n}$
$-\int_{X}u^{2p}e^{(1-a)(F-\log\mathrm{v})-a\delta f_{*}-ab\phi}\cdot\frac{a\delta\Delta f_{*}+ab\Delta\phi}{1-a}\omega^{n}$
=:absent:=:= : $I+II.$
One has
$I=$ $-\int_{X}u^{2p}e^{(1-a)(F-\log\mathrm{v})-a\delta f_{*}-ab\phi}\cdot\frac{1}{1-a}\Delta\Big{(}(1-a)(F-\log\mathrm{v})-a\delta f_{*}-ab\phi\Big{)}\omega^{n}$
$=$ $-\int_{X}\frac{u^{2p}e^{(1-a)(F-\log\mathrm{v})-a\delta f_{*}-ab\phi}}{a-1}\cdot|\nabla\big{(}(1-a)(F-\log\mathrm{v})-a\delta f_{*}-ab\phi\big{)}|^{2}\omega^{n}$
$-\int_{X}\frac{2p}{a-1}u^{2p-1}e^{(1-a)(F-\log\mathrm{v})-a\delta f_{*}-ab\phi}\cdot\langle\nabla u,\nabla\big{(}(1-a)(F-\log\mathrm{v})-a\delta f_{*}-ab\phi\big{)}\rangle\omega^{n}$
$\leq$ $-\int_{X}\frac{u^{2p}e^{(1-a)(F-\log\mathrm{v})-a\delta f_{*}-ab\phi}}{a-1}\cdot|\nabla\big{(}(1-a)(F-\log\mathrm{v})-a\delta f_{*}-ab\phi\big{)}|^{2}\omega^{n}$
$+\int_{X}\frac{u^{2p}e^{(1-a)(F-\log\mathrm{v})-a\delta f_{*}-ab\phi}}{(a-1)}\cdot|\nabla\big{(}(1-a)(F-\log\mathrm{v})-a\delta f_{*}-ab\phi\big{)}|^{2}\omega^{n}$
$+\int_{X}\frac{p^{2}}{a-1}u^{2p-2}e^{(1-a)(F-\log\mathrm{v})-a\delta f_{*}-ab\phi}\cdot|\nabla u|^{2}\omega^{n}$
$\leq$ $\frac{p^{2}}{a-1}\int_{X}u^{2p-2}e^{-a(F-\log\mathrm{v})-a\delta f_{*}-ab\phi}\cdot|\nabla u|^{2}\omega_{\phi}^{n}$
$=$ $\frac{p^{2}}{a-1}\int_{X}u^{2p-2}e^{-a(F+\log\mathrm{v}+\delta f_{*}+b\phi)}\cdot|\nabla u|^{2}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
$\leq$ $\frac{p^{2}}{a-1}\int_{X}u^{2p-2}e^{-a(F+\log\mathrm{v}+\delta f_{*}+b\phi)}\cdot|\nabla_{\phi}u|_{\phi}^{2}(n+\Delta\phi)e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
$=$ $\frac{p^{2}}{a-1}\int_{X}u^{2p-1}|\nabla_{\phi}u|_{\phi}^{2}e^{2a\log\mathrm{v}}\omega_{\phi}^{n},$
where we have used
$|\nabla^{\phi}u|_{\phi}^{2}\mathrm{tr}\omega_{\phi}\geq|\nabla u|^{2}$
for the sixth inequality.
Then the inequality (3.10) can be simplified into
$\int_{X}(2p+1-a)u^{2p-1}\cdot|\nabla_{\phi}u|_{\phi}^{2}\cdot e^{2a\log\mathrm{v}}\omega_{\phi}^{n}+\frac{ab}{2}\int_{X}u^{2p+1}\mathrm{tr}_{\phi}\omega e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
$\leq$ $(n+1)ab\int_{X}u^{2p+1}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}+C\int_{X}u^{2p}e^{-a(F+\log\mathrm{v}+\delta f_{*}+b\phi)}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
$-a(1-\delta)\int_{X}u^{2p+1}\Delta_{\phi}f_{*}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}+\frac{p^{2}}{a-1}\int_{X}u^{2p-1}|\nabla_{\phi}u|_{\phi}^{2}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}+II.$ (3.11)
Note that
$II-a(1-\delta)\int_{X}u^{2p+1}\Delta_{\phi}f_{*}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}+C\int_{X}u^{2p}e^{-a(F+\log\mathrm{v}+\delta f_{*}+b\phi)}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
$=$ $-\int_{X}u^{2p}e^{-a(F+\log\mathrm{v}+\delta f_{*}+b\phi)}\Big{(}\frac{a\delta\Delta f_{*}}{1-a}+\frac{ab\Delta\phi}{1-a}\Big{)}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
$-a(1-\delta)\int_{X}u^{2p+1}\Delta_{\phi}f_{*}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}+C\int_{X}u^{2p}e^{-a(F+\log\mathrm{v}+\delta f_{*}+b\phi)}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
$\leq$ $\int_{X}u^{2p}e^{-a(F+\log\mathrm{v}+\delta f_{*}+b\phi)}\Big{(}C-\frac{ab}{1-a}\cdot\Delta\phi\Big{)}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
$+\int_{X}u^{2p}e^{-a(F+\log\mathrm{v}+\delta f_{*}+b\phi)}\Big{(}-a(1-\delta)\Delta_{\phi}f_{*}\cdot(n+\Delta\phi)-\frac{a\delta}{1-a}\cdot\Delta f_{*}\Big{)}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
$\leq$ $Cb\int_{X}u^{2p+1}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}+C\int_{X}u^{2p+1}\mathrm{tr}_{\phi}\omega e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$ (3.12)
where we have used the following facts
$C-\frac{ab}{1-a}\Delta\phi$ $\leq C+\frac{ab}{a-1}(n+\Delta\phi)-\frac{abn}{a-1}\leq Cb(n+\Delta\phi),$
since $b$ is sufficiently large, and
$-a(1-\delta)\Delta_{\phi}f_{*}(n+\Delta\phi)+\frac{a}{a-1}\cdot\delta\cdot\Delta f_{*}=-\Delta_{\phi}f_{*}\cdot(n+\Delta\phi)+\Delta f_{*}$
$=$ $-\sum_{i\neq j}\frac{(f_{*})_{i\overline{i}}(1+\phi_{j\overline{j}})}{1+\phi_{i\overline{i}}}\leq(\frac{1-t}{t})\sum_{i\neq j}\frac{\theta_{i\overline{i}}(1+\phi_{j\overline{j}})}{1+\phi_{i\overline{i}}}\leq C\cdot\mathrm{tr}_{\phi}\omega\cdot(n+\Delta\phi),$
where we let $\delta=\frac{a-1}{a}$.
Choose $a=p+1$, then $(2p+1-a)-\frac{p^{2}}{a-1}=0$. By (3.11), (3.12) and choosing $b$ sufficiently large, we have
$\frac{ab}{8}\int_{X}u^{2p+1}\mathrm{tr}_{\phi}\omega e^{2a\log\mathrm{v}}\omega_{\phi}^{n}\leq C^{\prime}b\int_{X}u^{2p+1}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}.$
By
$\mathrm{tr}\omega_{\phi}\leq\frac{1}{(n-1)!}(\mathrm{tr}_{\phi}\omega)^{n-1}\frac{\omega_{\phi}^{n}}{\omega^{n}},$
we obtain
$\int_{X}u^{2p+1}(n+\Delta\phi)^{\frac{1}{n-1}}e^{\frac{-1}{n-1}(F-\log\mathrm{v})}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}\leq C\int_{X}u^{2p+1}\omega_{\phi}^{n}$
It follows that
$\int_{X}e^{(-a(2p+1)+\frac{n-2}{n-1})F-(2p+1)(a-1)f_{*}-ab(2p+1)\phi+(2a+\frac{1}{n-1}-a(2p+1))\log\mathrm{v}}(n+\Delta\phi)^{2p+1+\frac{1}{n-1}}\omega^{n}$
$\leq C\int_{X}e^{(1-a(2p+1))F-a(2p+1)\log\mathrm{v}-(2p+1)(a-1)f_{*}-ab(2p+1)\phi}(n+\Delta\phi)^{2p+1}\omega^{n}.$
Use $\|\phi\|_{C^{0}},\|F+f_{*}\|_{C^{0}},|\log\mathrm{v}|\leq C$, then we derive
$\int_{X}e^{(2p+1-\frac{n-2}{n-1})f_{*}}(n+\Delta\phi)^{2p+1+\frac{1}{n-1}}\omega^{n}\leq C\int_{X}e^{2pf_{*}}(n+\Delta\phi)^{2p+1}\omega^{n}.$
Take $p=k\cdot\frac{1}{n-1}$, then the inductive argument concludes the proof.
3.C. Gradient estimate of $F-\log\mathrm{v}+f_{*}$
Theorem 3.7.
Assume $\mathrm{v}$ is log-concave. Let $\phi$ be a solution of (3.4), then
$\sup_{X}(|\nabla^{\phi}(F-\log\mathrm{v}(m_{\phi})+f_{*})|_{\phi})\leq C,$
where $C$ is a constant depending on $\omega$, $[\theta]$, $\mathrm{v}$, $\mathrm{w}$, $\|\phi\|_{C^{0}}$, $\|F+f_{*}\|_{C^{0}}$, $\|\mathrm{tr}\omega_{\phi}\|_{L^{p}}$ for some $p$.
Furthermore, for any $p>0$, with a constant $C_{p}$ that depends on the same parameters above,
$\int_{X}|\nabla^{\phi}(F+f_{*})|_{\phi}^{P}\omega^{n}<C_{p}.$ (3.13)
Remark 3.8.
We want to emphasize that, the proof of Theorem 3.7 only depends on the $C^{0}$-estimate Theorem 3.2 and $W^{2,p}$-estimate Theorem 3.5, but not on the log-concavity assumption. We put the log-concavity assumption in the statement of Theorem 3.7 only for its compatibility with our main result.
Proof.
Let $W=F^{\prime}+f_{*}=F-\log\mathrm{v}+f_{*}$.
Let’s begin by reviewing some preliminaries that are needed along the proof. We have the following formulas and inequalities:
$\mathrm{tr}\omega_{\phi}\geq ne^{\frac{F^{\prime}}{n}}\geq C>0,$
which is due to the arithmetic-geometric inequality and the lower bound of $F^{\prime}$,
$\mathrm{Ric}(\phi)_{i\overline{j}}$ $=\mathrm{Ric}_{i\overline{j}}-F^{\prime}_{i\overline{j}}=\mathrm{Ric}_{i\overline{j}}-W_{i\overline{j}}+(f_{*})_{i\overline{j}},$
$\mathrm{tr}_{\phi}\omega$ $\leq n^{-1}e^{-F^{\prime}}(\mathrm{tr}\omega_{\phi})^{n-1}\leq C(\mathrm{tr}\omega_{\phi})^{n-1},$
$|\mathrm{Ric}(\omega)(\nabla^{\phi}W,\nabla^{\phi}W)|$ $\leq C\cdot e^{-F^{\prime}}\cdot(\mathrm{tr}\omega_{\phi})^{n-1}\cdot|\nabla^{\phi}W|_{\phi}^{2},$
$\mathrm{Ric}(\phi)_{i\overline{j}}W_{\overline{i}}W_{j}$ $=\mathrm{Ric}(\omega)(\nabla^{\phi}W,\nabla^{\phi}W)-\frac{W_{i\overline{j}}W_{\overline{i}}W_{j}}{(1+\phi_{i\overline{i}})(1+\phi_{j\overline{j}})}+\frac{(f_{*})_{i\overline{j}}W_{\overline{i}}W_{j}}{(1+\phi_{i\overline{i}})(1+\phi_{j\overline{j}})}$
$\geq-Ce^{-F^{\prime}}(\mathrm{tr}\omega_{\phi})^{n-1}|\nabla^{\phi}W|_{\phi}^{2}-\frac{W_{i\overline{j}}W_{\overline{i}}W_{j}}{(1+\phi_{i\overline{i}})(1+\phi_{j\overline{j}})}.$
Since $f_{*}$ is a $(\frac{1-t}{t})\theta$-psh function , $\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha}(f_{*})$ is determined by class $[\theta]$, which is bounded. And by Theorem 3.2, $|W|_{C^{0}}$ is bounded. With the preliminaries in mind and under an orthonormal frame of $g_{\phi}$, using Bochner formula (see the proof of [CC21a, Proposition 4.1]) implies
$e^{\frac{-W}{2}}\Delta_{\phi}(e^{\frac{W}{2}}|\nabla^{\phi}W|_{\phi}^{2})\geq$ $(\Delta_{\phi}W)_{i}W_{\overline{i}}+(\Delta_{\phi}W)_{\overline{i}}W_{i}+\mathrm{Ric}(\phi)_{i\overline{j}}W_{\overline{i}}W_{j}+|W_{i\overline{j}}|_{\phi}^{2}+|W_{ij}|_{\phi}^{2}$
$+\frac{1}{4}|\nabla^{\phi}W|_{\phi}^{4}+\frac{1}{2}(W_{i}W_{\overline{j}}W_{\overline{i}j}+W_{\overline{i}}W_{j}W_{i\overline{j}}+W_{i}W_{j}W_{\overline{i}\overline{j}}+W_{\overline{i}}W_{\overline{j}}W_{ij})$
$+\frac{1}{2}\Delta_{\phi}W\cdot|\nabla^{\phi}W|_{\phi}^{2}$
$\geq$ $(\Delta_{\phi}W)_{i}W_{\overline{i}}+(\Delta_{\phi}W)_{\overline{i}}W_{i}+|W_{i\overline{j}}|_{\phi}^{2}$
$+\frac{1}{2}\Delta_{\phi}W\cdot|\nabla^{\phi}W|_{\phi}^{2}-Ce^{-F^{\prime}}(\mathrm{tr}\omega_{\phi})^{n-1}\cdot|\nabla^{\phi}W|_{\phi}^{2}$
$\geq$ $2\nabla^{\phi}W\cdot_{\phi}\nabla^{\phi}\Delta_{\phi}W+\frac{|W_{i\overline{j}}|^{2}}{(1+\phi_{i\overline{i}})(1+\phi_{j\overline{j}})}-Ce^{-F^{\prime}}(\mathrm{tr}\omega_{\phi})^{n-1}\cdot|\nabla^{\phi}W|_{\phi}^{2}$
$+\frac{1}{2}\big{(}-C-Ce^{-F^{\prime}}(\mathrm{tr}\omega_{\phi})^{n-1}+2\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha}(W)\Big{)}|\nabla^{\phi}W|_{\phi}^{2}$
$\geq$ $2\nabla^{\phi}W\cdot_{\phi}\nabla^{\phi}\Delta_{\phi}W+\frac{|W_{i\overline{j}}|^{2}}{(1+\phi_{i\overline{i}})(1+\phi_{j\overline{j}})}$
$-C(1+e^{-F^{\prime}}(\mathrm{tr}\omega_{\phi})^{n-1})\cdot|\nabla^{\phi}W|_{\phi}^{2}+\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha}(W)|\nabla^{\phi}W|_{\phi}^{2}.$ (3.14)
Note that
$\Delta_{\phi}W=$ $\Delta_{\phi}(F+f_{*})-\Delta_{\phi}\log\mathrm{v}$
$=$ $\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha}(F+f_{*}-\log\mathrm{v})+\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha}\log\mathrm{v}-\Delta_{\phi}\log\mathrm{v}+\widetilde{f}_{t}+\mathrm{tr}_{\phi}\eta_{t}$
$=$ $\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha}(W)+\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha}(F^{\prime}+f_{*})-\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha}(f_{*})+\left(\frac{\mathrm{v}_{,\alpha}\mathrm{v}_{,\beta}}{\mathrm{v}^{2}}-\frac{\mathrm{v}_{,\alpha\beta}}{\mathrm{v}}\right)\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}$ (3.15)
$-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}\Delta_{\omega}(m_{\omega}^{\xi^{\alpha}})-\frac{\mathrm{v}_{,\alpha}\mathrm{v}_{,\beta}}{\mathrm{v}^{2}}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}+\widetilde{f}_{t}+\mathrm{tr}_{\phi}\eta_{t},$
$2\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha}(W)-C-C\mathrm{tr}_{\phi}\omega-C\mathrm{tr}\omega_{\phi}\leq\Delta_{\phi}W\leq 2\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha}(W)+C+C\mathrm{tr}_{\phi}\omega+C\mathrm{tr}\omega_{\phi},$
where we use (3.4) and (2.16) for the second equality, (2.18) for the last inequality.
To control the derivative term $J\xi^{\alpha}(W)|\nabla^{\phi}W|_{\phi}^{2}$, we need to add another barrier function. One computes
$\Delta_{\phi}(e^{-\log\mathrm{v}(m_{\phi})}\cdot e^{\frac{W}{2}}|\nabla^{\phi}W|_{\phi}^{2})$
$=$ $g_{\phi}^{i\overline{j}}\partial_{i}\left(e^{-\log\mathrm{v}(m_{\phi})}(-(\log\mathrm{v}(m_{\phi}))_{\overline{j}})(e^{\frac{W}{2}}|\nabla^{\phi}W|_{\phi}^{2})+e^{-\log\mathrm{v}(m_{\phi})}\partial_{\overline{j}}(e^{\frac{W}{2}}|\nabla^{\phi}W|_{\phi}^{2})\right)$
$=$ $e^{\frac{W}{2}-\log\mathrm{v}(m_{\phi})}|\nabla^{\phi}W|_{\phi}^{2}|\nabla^{\phi}\log\mathrm{v}(m_{\phi})|_{\phi}^{2}+e^{-\log\mathrm{v}(m_{\phi})}(-\Delta_{\phi}\log\mathrm{v}(m_{\phi}))e^{\frac{W}{2}}|\nabla^{\phi}W|_{\phi}^{2}$
$-g_{\phi}^{i\overline{j}}e^{-\log\mathrm{v}(m_{\phi})}(\log\mathrm{v}(m_{\phi}))_{\overline{j}}(e^{\frac{W}{2}}|\nabla^{\phi}W|_{\phi}^{2})_{i}-g_{\phi}^{i\overline{j}}e^{-\log\mathrm{v}(m_{\phi})}(\log\mathrm{v}(m_{\phi}))_{i}(e^{\frac{W}{2}}|\nabla^{\phi}W|_{\phi}^{2})_{\overline{j}}$
$+e^{-\log\mathrm{v}(m_{\phi})}\Delta_{\phi}(e^{\frac{W}{2}}|\nabla^{\phi}W|_{\phi}^{2})$
$=$ $e^{\frac{W}{2}-\log\mathrm{v}(m_{\phi})}|\nabla^{\phi}W|_{\phi}^{2}|\nabla^{\phi}\log\mathrm{v}(m_{\phi})|_{\phi}^{2}-2e^{-\log\mathrm{v}(m_{\phi})}\nabla^{\phi}\log\mathrm{v}(m_{\phi})\cdot_{\phi}\nabla^{\phi}(e^{\frac{W}{2}}|\nabla^{\phi}W|_{\phi}^{2})$
$+e^{-\log\mathrm{v}(m_{\phi})}\left(\left(\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}(W)-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(\phi)}J\xi^{\alpha}(f_{*})-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}\Delta_{\omega}(m_{\omega}^{\xi^{\alpha}})\right.\right.$
$\left.\left.-\frac{\mathrm{v}_{,\alpha\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}+\frac{\mathrm{v}_{,\alpha}(m_{\phi})\mathrm{v}_{,\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})^{2}}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}\right)e^{\frac{W}{2}}|\nabla^{\phi}W|_{\phi}^{2}+\Delta_{\phi}(e^{\frac{W}{2}}|\nabla^{\phi}W|_{\phi}^{2})\right)$
$\geq$ $e^{\frac{W}{2}-\log\mathrm{v}(m_{\phi})}|\nabla^{\phi}W|_{\phi}^{2}|\nabla^{\phi}\log\mathrm{v}(m_{\phi})|_{\phi}^{2}-2e^{-\log\mathrm{v}(m_{\phi})}\nabla^{\phi}\log\mathrm{v}(m_{\phi})\cdot_{\phi}\nabla^{\phi}(e^{\frac{W}{2}}|\nabla^{\phi}W|_{\phi}^{2})$
$+e^{\frac{W}{2}-\log\mathrm{v}(m_{\phi})}2\nabla^{\phi}W\cdot_{\phi}\nabla^{\phi}\Delta_{\phi}W-C_{\mathrm{v},\xi}(1+(\mathrm{tr}\omega_{\phi})^{n-1})e^{\frac{W}{2}-\log\mathrm{v}(m_{\phi})}|\nabla^{\phi}W|_{\phi}^{2}$
$+\frac{|W_{i\overline{j}}|^{2}}{(1+\phi_{i\overline{i}})(1+\phi_{j\overline{j}})}e^{\frac{F^{\prime}}{2}-\log\mathrm{v}(m_{\phi})}+2\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}(W)e^{\frac{W}{2}-\log\mathrm{v}(m_{\phi})}|\nabla^{\phi}W|_{\phi}^{2},$ (3.16)
where we use (2.16) for the third equality, (2.18) and (3.14) for the last inequality.
Set
$u=e^{\frac{W}{2}-\log\mathrm{v}(m_{\phi})}|\nabla^{\phi}W|_{\phi}^{2}+1,$
then we have
$\Delta_{\phi}u\geq$ $-C_{\mathrm{v},\xi}(\mathrm{tr}\omega_{\phi})^{n-1}u+2e^{\frac{W}{2}-\log\mathrm{v}(m_{\phi})}\nabla^{\phi}W\cdot_{\phi}\nabla^{\phi}\Delta_{\phi}W+2\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}(W)u$
$-2\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha}(W)+e^{\frac{W}{2}-\log\mathrm{v}(m_{\phi})}|\nabla^{\phi}W|_{\phi}^{2}|\nabla^{\phi}\log\mathrm{v}(m_{\phi})|_{\phi}^{2}$
$-2e^{-\log\mathrm{v}(m_{\phi})}\nabla^{\phi}\log\mathrm{v}(m_{\phi})\cdot_{\phi}\nabla^{\phi}(e^{\frac{W}{2}}|\nabla^{\phi}W|_{\phi}^{2}).$ (3.17)
For any $p>0$,
$\frac{1}{2p+1}\Delta_{\phi}(u^{2p+1})=u^{2p}\Delta_{\phi}u+2pu^{2p-1}|\nabla^{\phi}u|_{\phi}^{2}.$
Taking the integration with respect to the weighted measure $e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}$, one has
$\int_{X}2pu^{2p-1}|\nabla^{\phi}u|_{\phi}^{2}e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}$
$=$ $\int_{X}u^{2p}(-\Delta_{\phi}u)e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}+\frac{1}{2p+1}\int_{X}\Delta_{\phi}(u^{2p+1})e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}.$ (3.18)
On the one hand,
$\frac{1}{2p+1}\int_{X}\Delta_{\phi}(u^{2p+1})e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}$
$=$ $-\frac{1}{2p+1}\int_{X}\nabla^{\phi}(u^{2p+1})\cdot_{\phi}\nabla^{\phi}e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}$
$=$ $-\int_{X}2u^{2p}\nabla^{\phi}(e^{-\log\mathrm{v}(m_{\phi})}e^{\frac{W}{2}}|\nabla^{\phi}W|_{\phi}^{2})\cdot_{\phi}\nabla^{\phi}\log\mathrm{v}(m_{\phi})e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}$
$=$ $-2\int_{X}u^{2p}\nabla^{\phi}(e^{\frac{W}{2}}|\nabla^{\phi}W|_{\phi}^{2})\cdot_{\phi}\nabla^{\phi}\log\mathrm{v}(m_{\phi})e^{\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}$
$+2\int_{X}u^{2p}e^{\frac{W}{2}}|\nabla^{\phi}W|_{\phi}^{2}|\nabla^{\phi}\log\mathrm{v}(m_{\phi})|_{\phi}^{2}e^{\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}.$ (3.19)
On the other hand, by (3.17),
$\int_{X}u^{2p}(-\Delta_{\phi}u)e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}$
$\leq$ $C\int_{X}u^{2p+1}(\mathrm{tr}\omega_{\phi})^{n-1}e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}-2\int_{X}u^{2p+1}\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}(W)e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}$
$-\int_{X}u^{2p}e^{\frac{W}{2}}|\nabla^{\phi}W|_{\phi}^{2}|\nabla^{\phi}\log\mathrm{v}(m_{\phi})|_{\phi}^{2}e^{\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}$
$+2\int_{X}u^{2p}\nabla^{\phi}\log\mathrm{v}(m_{\phi})\cdot_{\phi}\nabla^{\phi}(e^{\frac{W}{2}}|\nabla^{\phi}W|_{\phi}^{2})e^{\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}$
$-2\int_{X}u^{2p}e^{\frac{W}{2}}\nabla^{\phi}W\cdot_{\phi}\nabla^{\phi}\Delta_{\phi}We^{\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}.$ (3.20)
where we have used
$|J\xi^{\alpha}(W)|\leq C|\nabla W|\leq C(\mathrm{tr}\omega_{\phi})^{\frac{1}{2}}|\nabla^{\phi}W|_{\phi}\leq C(\mathrm{tr}\omega_{\phi})^{\frac{1}{2}}C\sqrt{u}\leq C(\mathrm{tr}\omega_{\phi})^{\frac{1}{2}}u\leq C(\mathrm{tr}\omega_{\phi})^{n-1}u.$ (3.21)
By (3.C) and (3.20), we have
$\int_{X}2pu^{2p-1}|\nabla^{\phi}u|_{\phi}^{2}e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}$
$\leq$ $C\int_{X}u^{2p+1}(\mathrm{tr}\omega_{\phi})^{n-1}e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}-2\int_{X}u^{2p+1}\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}(W)e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}$
$-2\int_{X}u^{2p}e^{\frac{W}{2}}\nabla^{\phi}W\cdot_{\phi}\nabla^{\phi}\Delta_{\phi}We^{\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}.$ (3.22)
Using the integration by parts,
$-2\int_{X}u^{2p}e^{\frac{W}{2}}\nabla^{\phi}W\cdot_{\phi}\nabla^{\phi}\Delta_{\phi}We^{\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}$
$=$ $\int_{X}4pu^{2p-1}e^{\frac{W}{2}}\Delta_{\phi}W(\nabla^{\phi}W\cdot_{\phi}\nabla^{\phi}u)e^{\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}+\int_{X}2u^{2p}e^{\frac{W}{2}}(\Delta_{\phi}W)^{2}e^{\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}$
$+\int u^{2p}e^{\frac{W}{2}}|\nabla^{\phi}W|_{\phi}^{2}(\Delta_{\phi}W)e^{\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}$
$+2\int_{X}u^{2p}e^{\frac{W}{2}}(\Delta_{\phi}W)(\nabla^{\phi}W\cdot_{\phi}\nabla^{\phi}\log\mathrm{v}(m_{\phi}))e^{\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}.$ (3.23)
We estimate the four terms on the right side of (3.23). First,
$\int_{X}4pu^{2p-1}e^{\frac{W}{2}}\Delta_{\phi}W(\nabla^{\phi}W\cdot_{\phi}\nabla^{\phi}u)e^{\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}$
$\leq$ $\int_{X}pu^{2p-1}|\nabla^{\phi}u|_{\phi}^{2}e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}+\int_{X}4pu^{2p-1}e^{W}(\Delta_{\phi}W)^{2}|\nabla^{\phi}W|_{\phi}^{2}\omega_{\phi}^{n}$
$\leq$ $\int_{X}pu^{2p-1}|\nabla^{\phi}u|_{\phi}^{2}e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}+\int_{X}4pu^{2p}e^{\frac{W}{2}}(\Delta_{\phi}W)^{2}e^{\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}.$ (3.24)
By formula (3.15),
$\int u^{2p}e^{\frac{W}{2}}|\nabla^{\phi}W|_{\phi}^{2}(\Delta_{\phi}W)e^{\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}$
$\leq$ $\int u^{2p}e^{\frac{W}{2}}|\nabla^{\phi}W|_{\phi}^{2}\left(C+C\mathrm{tr}_{\phi}\omega+C\mathrm{tr}\omega_{\phi}+2\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}(W)\right)e^{\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}$
$=$ $2\int_{X}u^{2p+1}\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}(W)e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}-2\int_{X}u^{2p}\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}(W)e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}$
$+\int u^{2p}e^{\frac{W}{2}}|\nabla^{\phi}W|_{\phi}^{2}(C+C\mathrm{tr}_{\phi}(\omega)+C\mathrm{tr}\omega_{\phi})e^{\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}$
$\leq$ $2\int_{X}u^{2p+1}\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}(W)e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}+C_{\mathrm{v},\xi}\int_{X}u^{2p+1}(\mathrm{tr}\omega_{\phi})^{n-1}e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n},$ (3.25)
where we have used (3.21) for the last inequality above. Again by formula (3.15),
$|\Delta_{\phi}W|\leq C+B_{\mathrm{v},\xi}(\mathrm{tr}\omega_{\phi})^{n-1}+2\left|\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}\right||J\xi^{\alpha}(W)|\leq$ $B_{\mathrm{v},\xi}(\mathrm{tr}\omega_{\phi})^{n-1}+C_{\mathrm{v}}|\nabla W|$
$\leq$ $B_{\mathrm{v},\xi}(\mathrm{tr}\omega_{\phi})^{n-1}+C_{\mathrm{v}}(\mathrm{tr}\omega_{\phi})^{\frac{1}{2}}|\nabla^{\phi}W|_{\phi}.$
where $B_{\mathrm{v},\xi}$ is a positive constant such that
$|\mathrm{tr}_{\phi}\eta_{t}|+\left|\left(\frac{\mathrm{v}_{,\alpha}(m_{\phi})\mathrm{v}_{\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})^{2}}-\frac{\mathrm{v}_{,\alpha\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})}\right)\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}\right|\leq B_{\mathrm{v},\xi}(\mathrm{tr}\omega_{\phi})^{n-1}.$
It follows that
$\int_{X}u^{2p}e^{\frac{W}{2}}(\Delta_{\phi}W)^{2}e^{\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}$
$\leq$ $\int_{X}u^{2p}e^{\frac{W}{2}}C\Big{(}(\mathrm{tr}\omega_{\phi})^{2n-2}+(\mathrm{tr}\omega_{\phi})^{n-1}\sqrt{\mathrm{tr}\omega_{\phi}}|\nabla^{\phi}W|_{\phi}+\mathrm{tr}\omega_{\phi}|\nabla^{\phi}W|_{\phi}^{2}\Big{)}e^{\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}$
$\leq$ $\int_{X}C(\mathrm{tr}\omega_{\phi})^{2(n-1)}u^{2p+1}e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}.$ (3.26)
By Hölder inequality,
$2\int_{X}u^{2p}e^{\frac{W}{2}}(\Delta_{\phi}W)(\nabla^{\phi}W\cdot_{\phi}\nabla^{\phi}\log\mathrm{v}(m_{\phi}))e^{\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}$
$\leq$ $\int_{X}u^{2p}e^{\frac{W}{2}}(\Delta_{\phi}W)^{2}|\nabla^{\phi}\log\mathrm{v}(m_{\phi})|_{\phi}^{2}e^{\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}+\int_{X}u^{2p}e^{\frac{W}{2}}|\nabla^{\phi}W|_{\phi}^{2}e^{\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}$
$\leq$ $\int_{X}C(\mathrm{tr}\omega_{\phi})^{3(n-1)}u^{2p+1}e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}+\int_{X}u^{2p+1}e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n},$ (3.27)
where we have used (3.C) and
$|\nabla^{\phi}\log\mathrm{v}(m_{\phi})|_{\phi}^{2}=\frac{\mathrm{v}_{,\alpha}(m_{\phi})\mathrm{v}_{,\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})^{2}}\langle\xi^{\beta},\xi^{\alpha}\rangle_{\phi}\leq C_{\mathrm{v}}(1+\phi_{i\overline{i}})\xi_{\alpha,i}\xi_{\beta,\overline{i}}\leq C_{\mathrm{v},\xi}\mathrm{tr}\omega_{\phi}.$ (3.28)
By (3.23), (3.24), (3.25), (3.C) and (3.27), we obtain
$\begin{split}&-2\int_{X}u^{2p}e^{\frac{W}{2}}\nabla^{\phi}W\cdot_{\phi}\nabla^{\phi}\Delta_{\phi}We^{\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}\\ \leq&\int_{X}pu^{2p-1}|\nabla^{\phi}u|_{\phi}^{2}e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}+C(p+1)\int_{X}(\mathrm{tr}\omega_{\phi})^{3(n-1)}u^{2p+1}e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}\\ &+2\int_{X}u^{2p+1}\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}(W)e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}+\int_{X}u^{2p+1}e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}\end{split}$ (3.29)
Therefore, by (3.C), (3.29), we deduce
$\int_{X}pu^{2p-1}|\nabla^{\phi}u|_{\phi}^{2}e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}\leq C(p+1)\int_{X}(\mathrm{tr}\omega_{\phi})^{3(n-1)}u^{2p+1}e^{\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}.$
It follows that
$\int_{X}|\nabla^{\phi}(u^{p+\frac{1}{2}})|_{\phi}^{2}\omega^{n}\leq\frac{C_{\mathrm{v},\xi}(p+\frac{1}{2})^{2}(p+1)}{p}\int_{X}C_{\xi}(\mathrm{tr}\omega_{\phi})^{3(n-1)}u^{2p+1}\omega_{\phi}^{n}.$
In the inequality above, we use $\omega^{n}=e^{f_{*}}\cdot e^{-(F^{\prime}+f_{*})}\omega_{\phi}^{n}\leq C\omega_{\phi}^{n}$, since $F^{\prime}+f_{*}$ is uniformly bounded, and $f_{*}\leq 0$.
For any $20<\varepsilon<$ 2, by Hölder inequality, we have
$\int_{X}(\mathrm{tr}\omega_{\phi})^{3(n-1)}u^{2p+1}\omega_{\phi}^{n}$ $=\int_{X}(\mathrm{tr}\omega_{\phi})^{3(n-1)}u^{2p+1}e^{-f_{*}}e^{F^{\prime}+f_{*}}\omega^{n}$
$\leq C\left(\int_{X}u^{(p+\frac{1}{2})(2+\varepsilon)}\omega^{n}\right)^{\frac{2}{2+\varepsilon}}\left(\int_{X}(\mathrm{tr}\omega_{\phi})^{3(n-1)\frac{2+\varepsilon}{\varepsilon}}e^{-\frac{2+\varepsilon}{\varepsilon}f_{*}}\omega^{n}\right)^{\frac{\varepsilon}{2+\varepsilon}}.$
Denote $v:=u^{p+\frac{1}{2}}$, then the above inequalities imply
$\int_{X}|\nabla^{\phi}v|_{\phi}^{2}\omega^{n}\leq\frac{C_{\mathrm{v},\xi}(p+\frac{1}{2})^{3}}{p}\left(\int_{X}v^{(2+\varepsilon)}\omega^{n}\right)^{\frac{2}{2+\varepsilon}}\left(\int_{X}(\mathrm{tr}\omega_{\phi})^{3(n-1)\frac{2+\varepsilon}{\varepsilon}}e^{-\frac{2+\varepsilon}{\varepsilon}f_{*}}\omega^{n}\right)^{\frac{\varepsilon}{2+\varepsilon}}$
Similar with (4.26) in the proof of [CC21a, Proposition 4.1], we also have
$\left(\int_{X}|\nabla v|^{2-\varepsilon}\omega^{n}\right)^{\frac{2}{2-\varepsilon}}\leq$ $n^{\frac{\varepsilon}{2-\varepsilon}}\left(\int_{X}(\mathrm{tr}\omega_{\phi})^{\frac{2-\varepsilon}{\varepsilon}}\omega^{n}\right)^{\frac{\varepsilon}{2-\varepsilon}}\int_{X}|\nabla^{\phi}v|_{\phi}^{2}\omega^{n}$
$\leq$ $C_{\mathrm{v},\xi}p^{2}K_{\varepsilon}\left(\int_{X}v^{(2+\varepsilon)}\omega^{n}\right)^{\frac{2}{2+\varepsilon}},$
where
$K_{\varepsilon}:=n^{\frac{\varepsilon}{2-\varepsilon}}\left(\int_{X}(\mathrm{tr}\omega_{\phi})^{\frac{2-\varepsilon}{\varepsilon}}\omega^{n}\right)^{\frac{\varepsilon}{2-\varepsilon}}\left(\int_{X}(\mathrm{tr}\omega_{\phi})^{3(n-1)\frac{2+\varepsilon}{\varepsilon}}e^{-\frac{2+\varepsilon}{\varepsilon}f_{*}}\omega^{n}\right)^{\frac{\varepsilon}{2+\varepsilon}}$
is bounded, which can be seen by using Theorem 3.5, the assumption of $f_{*}$ and Hölder inequality.
Applying the Sobolev inequality ([Aub13, Theorem 2.21]) with exponent $2-\varepsilon$, and denote $\theta:=\frac{2n(2-\varepsilon)}{2n-2+\varepsilon}$ to be the index, we obtain
$\|v\|_{L^{\theta}(\omega^{n})}\leq C_{s}(\|\nabla v\|_{L^{2-\varepsilon}(\omega^{n})}+\|v\|_{L^{2-\varepsilon}(\omega^{n})}).$
It follows that
$\left(\int_{X}u^{(p+\frac{1}{2})\theta}\omega^{n}\right)^{\frac{1}{\theta}}\leq$ $C_{s}\left(\Big{(}\int_{X}|\nabla(u^{p+\frac{1}{2}})|^{2-\varepsilon}\omega^{n}\Big{)}^{\frac{1}{2-\varepsilon}}+\Big{(}\int_{X}u^{(p+\frac{1}{2})(2-\varepsilon)}\omega^{n}\Big{)}^{\frac{1}{2-\varepsilon}}\right)$
$\leq$ $C_{s}\left(C_{\mathrm{v},\xi}p\sqrt{K_{\varepsilon}}\Big{(}\int_{X}u^{(p+\frac{1}{2})(2+\varepsilon)}\omega^{n}\Big{)}^{\frac{1}{2+\varepsilon}}+\Big{(}\int_{X}u^{(p+\frac{1}{2})(2-\varepsilon)}\omega^{n}\Big{)}^{\frac{1}{2-\varepsilon}}\right)$
$\leq$ $C_{\varepsilon}p\left(\int_{X}u^{(p+\frac{1}{2})(2+\varepsilon)}\omega^{n}\right)^{\frac{1}{2+\varepsilon}}.$
One can choose $\varepsilon>0$ sufficiently small such that $\theta>(2+\varepsilon)$ (e.g. $\varepsilon=\frac{1}{2n}$). We deduce
$\|u\|_{L^{(p+\frac{1}{2})\theta}}\leq(C_{\varepsilon}p)^{\frac{1}{p+\frac{1}{2}}}\|u\|_{L^{(p+\frac{1}{2})(2+\varepsilon)}}.$
Denote $\chi:=\frac{\theta}{2+\varepsilon}>1$ and choose $p+\frac{1}{2}=\chi^{i}$, then we have
$\|u\|_{L^{(2+\varepsilon)\chi^{i+1}}}\leq C^{\frac{1}{\chi^{i}}}(\chi^{i})^{\frac{1}{\chi^{i}}}\|u\|_{L^{(2+\varepsilon)\chi^{i}}}.$
It follows that
$\|u\|_{L^{\infty}}\leq C^{\sum_{i\geq 0}\frac{1}{\chi^{i}}}\chi^{\sum_{i\geq 0}\frac{i}{\chi^{i}}}\|u\|_{L^{(2+\varepsilon)\chi^{i}}}$
Then we
$\|u\|_{L^{\infty}}\leq C\|u\|_{L^{2+\varepsilon}}\leq C\|u\|_{L^{1}}^{\frac{1}{2+\varepsilon}}\|u\|_{L^{\infty}}^{\frac{1+\varepsilon}{2+\varepsilon}}.$ (3.30)
Thus, by (3.30), Theorem 3.5 and Lemma 3.9, we derive the $\|u\|_{L^{\infty}}$, which implies $\sup_{X}|\nabla^{\phi}(F+f_{*}-\log(\mathrm{v}(m_{\phi})))|_{\phi}\leq C$.
Furthermore, since $|\nabla^{\phi}\log\mathrm{v}|_{\phi}^{2}\leq C\cdot\mathrm{tr}\omega_{\phi}$, by Theorem 3.5, we have
$\int_{X}|\nabla^{\phi}(F+f_{*})|_{\phi}^{P}\omega^{n}<C_{p}.$ (3.31)
Lemma 3.9.
As above notation, the following estimate holds
$\|u\|_{L^{1}}\leq C.$ (3.32)
Proof.
Recall $u=e^{\frac{W}{2}-\log(\mathrm{v}(m_{\phi}))}|\nabla^{\phi}W|_{\phi}^{2}+1$. It suffices to show that
$\int_{X}|\nabla^{\phi}W|_{\phi}^{2}\omega_{\phi}^{n}\leq C.$
By (2.16) and (3.4), one observes that
$\Delta_{\phi}(F+f_{*}+\log\mathrm{v}(m_{\phi}))=$ $\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha}(F+f_{*})+\Delta_{\phi}\log(\mathrm{v}(m_{\phi}))+\widetilde{f}_{t}+\mathrm{tr}_{\phi}\widetilde{\eta}_{t}$
$=$ $\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha}(f_{*})+(\frac{\mathrm{v}_{,\alpha\beta}}{\mathrm{v}}-2\frac{\mathrm{v}_{,\alpha}\mathrm{v}_{\beta}}{\mathrm{v}^{2}})\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}+\widetilde{f}_{t}+\mathrm{tr}_{\phi}\widetilde{\eta}_{t}.$
Denote $\widetilde{W}=F+f_{*}+\log\mathrm{v}(m_{\phi})=W+2\log\mathrm{v}(m_{\phi})$. One has
$\Delta_{\phi}(\widetilde{W}^{2})=$ $2\widetilde{W}\Delta_{\phi}\widetilde{W}+2|\nabla^{\phi}\widetilde{W}|_{\phi}^{2}$
$=$ $2\widetilde{W}\left(\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha}(f_{*})+(\frac{\mathrm{v}_{,\alpha\beta}}{\mathrm{v}}-2\frac{\mathrm{v}_{,\alpha}\mathrm{v}_{\beta}}{\mathrm{v}^{2}})\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}+\widetilde{f}_{t}+\mathrm{tr}_{\phi}\widetilde{\eta}_{t}\right)+2|\nabla^{\phi}\widetilde{W}|_{\phi}^{2}.$
It follows that
$\int_{X}|\nabla^{\phi}\widetilde{W}|_{\phi}^{2}\omega_{\phi}^{n}\leq$ $\int_{X}\widetilde{W}(\widetilde{f}_{t}+\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha}(f_{*}))\omega_{\phi}^{n}+\int_{X}\widetilde{W}\left((\frac{\mathrm{v}_{,\alpha\beta}}{\mathrm{v}}-2\frac{\mathrm{v}_{,\alpha}\mathrm{v}_{\beta}}{\mathrm{v}^{2}})\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}+\mathrm{tr}_{\phi}\widetilde{\eta}_{t}\right)\omega_{\phi}^{n}$
$\leq$ $C+\|\widetilde{W}\|_{C^{0}}B_{\mathrm{v},\xi}\int_{X}(\mathrm{tr}\omega_{\phi}+\mathrm{tr}_{\phi}\omega)\omega_{\phi}^{n}$
$\leq$ $C+C\int_{X}(\mathrm{tr}\omega_{\phi})^{n-1}\omega_{\phi}^{n}\leq C.$
Therefore, by Theorem 3.5 and (3.28), we obtain
$\int_{X}|\nabla^{\phi}W|_{\phi}^{2}\omega_{\phi}^{n}=$ $\int_{X}|\nabla^{\phi}(\widetilde{W}-2\log\mathrm{v}(m_{\phi}))|_{\phi}^{2}\omega_{\phi}^{n}$
$\leq$ $\int_{X}|\nabla^{\phi}\widetilde{W}|_{\phi}^{2}\omega_{\phi}^{n}+2\int_{X}|\nabla^{\phi}\log\mathrm{v}(m_{\phi})|_{\phi}^{2}\omega_{\phi}^{n}$
$\leq$ $C+4\int_{X}\frac{\mathrm{v}_{,\alpha}(m_{\phi})\mathrm{v}_{,\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})^{2}}\langle\xi^{\beta},\xi^{\alpha}\rangle_{\phi}\omega_{\phi}^{n}$
$\leq$ $C+C_{\mathrm{v},\xi}\int_{X}\mathrm{tr}\omega_{\phi}\omega_{\phi}^{n}\leq C.$
4. Proof of Continuity Method
In this section, we prove the Theorem 1.2 by finishing the continuity method, concluding the openness and closedness of continuity path.
The weighted Mabuchi functional associated to the continuity path (2.13a), (2.13b) is
${\mathbb{M}}_{\mathrm{v},\mathrm{w},t}={\mathbb{H}}_{\mathrm{v}}(\phi)-{\mathbb{E}}_{\mathrm{v}}^{\mathrm{Ric}}(\phi)+{\mathbb{E}}_{\mathrm{w}\cdot\ell_{\mathrm{ext}}}(\phi)+\frac{1-t}{t}({\mathbb{E}}_{\mathrm{v}}^{\theta}(\phi)-\underline{\theta}\cdot{\mathbb{E}}_{\mathrm{w}}(\phi))-C_{0},$ (4.1)
where
$C_{0}=\int_{X}\log(\mathrm{v}(m_{\omega}))\mathrm{v}(m_{\omega})\omega^{n},$
$\int_{X}\mathrm{tr}_{\mathrm{v},\omega}(\theta)\frac{\omega^{n}}{n!}=\int_{X}\mathrm{v}(m_{\omega})\theta\wedge\frac{\omega^{n-1}}{(n-1)!}+\int_{X}\langle d\mathrm{v}(m_{\omega}),m_{\theta}\rangle\frac{\omega^{n}}{n!}=\underline{\theta}\cdot\int_{X}\mathrm{w}(m_{\omega})\frac{\omega^{n}}{n!}.$
Denote ${\mathbb{J}}^{\mathrm{Ric}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}}}(\phi)={\mathbb{E}}_{\mathrm{v}}^{\mathrm{Ric}}(\phi)-{\mathbb{E}}_{\mathrm{w}\cdot\ell_{\mathrm{ext}}}(\phi)$, ${\mathbb{J}}^{\theta}_{\mathrm{v},\mathrm{w}}(\phi)={\mathbb{E}}_{\mathrm{v}}^{\theta}(\phi)-\underline{\theta}\cdot{\mathbb{E}}_{\mathrm{w}}(\phi)$.
Then
${\mathbb{M}}_{\mathrm{v},\mathrm{w},t}={\mathbb{H}}_{\mathrm{v}}(\phi)-{\mathbb{J}}^{\mathrm{Ric}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}}}(\phi)+\frac{1-t}{t}\cdot{\mathbb{J}}^{\theta}_{\mathrm{v},\mathrm{w}}(\phi)-C_{0}.$ (4.2)
4.A. Openness
In this subsection, we show the openness of the continuity path (1.3), following the standard implicit function theory ([Che18], [Has19]). Consider
${\mathcal{H}}^{4,\alpha}_{K}:=\{\phi\in C_{K}^{4,\alpha}(X)~{}|~{}\omega_{\phi}>0\}$
and define
${\mathcal{F}}:{\mathcal{H}}^{4,\alpha}_{K}\times[0,1]$ $\rightarrow C^{\alpha}(X)$
$(\phi,t)$ $\mapsto t\left(\frac{S_{\mathrm{v}}(\phi)}{\mathrm{w}(m_{\phi})}-\ell_{\mathrm{ext}}(m_{\phi})\right)-(1-t)\left(\frac{\mathrm{tr}_{\mathrm{v},\phi}(\theta)}{\mathrm{w}(m_{\phi})}-\underline{\theta}\right).$
Suppose there exists $(\phi_{0},t_{0})\in{\mathcal{H}}^{4,\alpha}_{K}\times[0,1]$ such that ${\mathcal{F}}(\phi_{0},t_{0})=0$. We will consider the following modified map with a normalization:
~$\widetilde{{\mathcal{F}}}:{\mathcal{H}}^{4,\alpha}_{K,0}\times[0,1]$ $\rightarrow C^{\alpha}(X)_{K,0}$
$(\phi,t)$ $\mapsto{\mathcal{F}}(\phi,t)-\int_{X}{\mathcal{F}}(\phi,t)\omega_{\phi_{0}}^{n},$
where
${\mathcal{H}}^{4,\alpha}_{K,0}:=\{\phi\in{\mathcal{H}}_{K}^{4,\alpha}~{}|~{}\int_{X}\phi\omega_{\phi_{0}}^{n}=0\}.$
Note that $\int_{X}{\mathcal{F}}(\phi,t)\omega_{\phi}^{n}=0$. Therefore if ~$\widetilde{{\mathcal{F}}}(\phi,t)=0$, ${\mathcal{F}}(\phi,t)$=0. Thus it suffices to apply implicit function theorem on the map $\widetilde{{\mathcal{F}}}$ near $(\phi_{0},t_{0})$. Denote
(ϕ):=assignitalic-$\mathcal{R}(\phi):=$ $\frac{S_{\mathrm{v}}(\phi)}{\mathrm{w}(m_{\phi})}-\ell_{\rm ext}(m_{\phi}),\text{ and }\mathcal{T}(\phi):=\frac{\mathrm{tr}_{\mathrm{v},\phi}(\theta)}{\mathrm{w}(m_{\phi})}.$
Thus
${\mathcal{F}}(\phi,t)=t{\mathcal{R}}(\phi)-(1-t)({\mathcal{T}}(\phi)-\underline{\theta}).$
Compute the linearization of $\widetilde{{\mathcal{F}}}$:
$D_{(\phi,t)}\widetilde{{\mathcal{F}}}:C^{4,\alpha}(X)_{K,0}\times{\mathbb{R}}$ $\rightarrow C^{\alpha}(X)_{K,0}$
$(u,s)$ $\mapsto D_{(\phi,t)}{\mathcal{F}}(u,s)-\int_{X}D_{(\phi,t)}{\mathcal{F}}(u,s)\omega_{\phi_{0}}^{n},$
$(D_{(\phi,t)}{\mathcal{F}})(u,s)=$ $\lim_{\delta\rightarrow 0}\frac{{\mathcal{F}}(\phi+\delta u,t+\delta s)-{\mathcal{F}}(\phi,t)}{\delta}$
$=$ $tD_{\phi}{\mathcal{R}}(u)-(1-t)D_{\phi}{\mathcal{T}}(u)+s({\mathcal{R}}(\phi)+{\mathcal{T}}(\phi)-\underline{\theta}).$
Using [Lah19, Lemma B.1], one computes
$\begin{split}t(D_{\phi}{\mathcal{R}})(u)=&-t\frac{(\mathcal{D}^{*}_{\phi}\mathrm{v}(m_{\phi})\mathcal{D}_{\phi})u}{\mathrm{w}(m_{\phi})}+t\langle\mathrm{d}({\mathcal{R}}(\phi)),\mathrm{d}u\rangle_{\phi},\\ =&-t\frac{(\mathcal{D}^{*}_{\phi}\mathrm{v}(m_{\phi})\mathcal{D}_{\phi})u}{\mathrm{w}(m_{\phi})}+\langle\mathrm{d}({\mathcal{F}}(\phi,t)),\mathrm{d}u\rangle_{\phi}+(1-t)\langle\mathrm{d}(\mathcal{T}(\phi)),\mathrm{d}u\rangle_{\phi},\end{split}$ (4.3)
where $\mathcal{D}_{\phi}u:=\sqrt{2}(\nabla^{\phi}\mathrm{d}u)^{-}$ is the $J$-anti-invariant part of the tensor $(\nabla^{\phi}\mathrm{d}u)$, with $\nabla^{\phi}$ the $g_{\phi}$-Levi-Civita connection, and $\mathcal{D}^{*}_{\phi}$ is the formal adjoint of $\mathcal{D}_{\phi}$.
We denote the operator $F_{\phi}:C^{4,\alpha}(X)_{K,0}\to C^{\alpha}(X)_{K,0}$ by
$F_{\phi}(u)=-D_{\phi}{\mathcal{T}}(u)+\langle\mathrm{d}(\mathcal{T}(\phi)),\mathrm{d}u\rangle_{\phi}.$
Then
$(D_{(\phi,t)}{\mathcal{F}})(u,s)$
$=$ $-t\frac{(\mathcal{D}^{*}_{\phi}\mathrm{v}(m_{\phi})\mathcal{D}_{\phi})u}{\mathrm{w}(m_{\phi})}+\langle\mathrm{d}({\mathcal{F}}(\phi,t)),\mathrm{d}u\rangle_{\phi}+(1-t)F_{\phi}(u)+s({\mathcal{R}}(\phi)+{\mathcal{T}}(\phi)-\underline{\theta}).$
In particular at $(\phi_{0},t_{0})$, we have
$(D_{(\phi_{0},t_{0})}{\mathcal{F}})(u,s)=-t_{0}\frac{(\mathcal{D}^{*}_{\phi_{0}}\mathrm{v}(m_{\phi_{0}})\mathcal{D}_{\phi_{0}})u}{\mathrm{w}(m_{\phi_{0}})}+(1-t_{0})F_{\phi}(u)+s({\mathcal{R}}(\phi_{0})+{\mathcal{T}}(\phi_{0})-\underline{\theta}).$
We denote
${\mathcal{L}}_{(\phi_{0},t_{0})}u:=-t_{0}\frac{(\mathcal{D}^{*}_{\phi_{0}}\mathrm{v}(m_{\phi_{0}})\mathcal{D}_{\phi_{0}})u}{\mathrm{w}(m_{\phi_{0}})}+(1-t_{0})F_{\phi_{0}}(u).$
Thus
$D_{(\phi_{0},t_{0})}\widetilde{{\mathcal{F}}}(u)={\mathcal{L}}_{(\phi_{0},t_{0})}u+s({\mathcal{R}}(\phi_{0})+{\mathcal{T}}(\phi_{0})-\underline{\theta}).$
By Lemma 5.4, we have that the operator ${\mathcal{L}}_{(\phi_{0},t_{0})}$ is a $(\cdot,\cdot)_{\mathrm{w},\phi}$-self-adjoint $K$-invariant linear operator satisfying
$({\mathcal{L}}_{(\phi_{0},t_{0})}u_{1},u_{2})_{\mathrm{w},\phi}=$ $\int_{X}{\mathcal{L}}_{(\phi_{0},t_{0})}u_{1}\cdot u_{2}\mathrm{w}(m_{\phi_{0}})\omega_{\phi_{0}}^{n}$
$=$ $-t_{0}\int_{X}\mathrm{v}(m_{\phi_{0}}){\mathcal{D}}_{\phi_{0}}u_{1}{\mathcal{D}}_{\phi_{0}}u_{2}\omega_{\phi_{0}}^{n}-(1-t_{0})\int_{X}\langle\theta,\mathrm{d}u_{1}\wedge\mathrm{d}^{c}u_{2}\rangle_{\phi_{0}}\mathrm{v}(m_{\phi_{0}})\omega_{\phi_{0}}^{n}$
$=$ $(u_{1},{\mathcal{L}}_{(\phi_{0},t_{0})}u_{2})_{\mathrm{w},\phi}.$
When $t_{0}>0$, following the argument in [Che18], we divide the remaining argument into three steps:
Step 1: show the inverse operator ${\mathcal{L}}_{(\phi_{0},t_{0})}^{-1}$ is bounded, that is, if $u\in C_{K}^{4,\alpha}(X)$ satisfies the equation ${\mathcal{L}}_{(\phi_{0},t_{0})}u=f$ for some $f\in C_{K}^{\alpha}(X)$ and $\int_{X}u\omega_{\phi_{0}}^{n}=0$, then there exists a constant $C$ depending on $\omega_{\phi_{0}}$ such that
$\|u\|_{C^{4,\alpha}(X)}\leq C\|f\|_{C^{\alpha}(X)}.$
The proof follows from [Che18, Lemma 4.3] by using Schauder estimates, Poincaré inequality.
Step 2: Show that $D_{(\phi_{0},t_{0})}\widetilde{{\mathcal{F}}}(\cdot,0)={\mathcal{L}}_{(\phi_{0},t_{0})}(\cdot):C^{4,\alpha}(X)_{K,0}\rightarrow C^{\alpha}(X)_{K,0}$ is an isomorphism.
Similar to [Che18, Lemma 4.4], the proof is given by a standard Fredholm theorem for elliptic operators.
Step 3: By the standard implicit function theory, there exists $\phi(t)\in{\mathcal{H}}^{4,\alpha}_{K}$ such that ${\mathcal{F}}(\phi(t),t)=0$.
When $t_{0}=0$, following [Has19], we modify the operator ${\mathcal{F}}$ as follows
${\mathcal{F}}_{0}:{\mathcal{H}}^{4,\alpha}_{K,0}$ $\rightarrow C^{\alpha}(X)_{K,0}$
$\phi$ $\mapsto\frac{S_{\mathrm{v}}(\phi)}{\mathrm{w}(m_{\phi})}-\ell_{\mathrm{ext}}(m_{\phi})-r\left(\frac{\mathrm{tr}_{\mathrm{v},\phi}(\theta)}{\mathrm{w}(m_{\phi})}-\underline{\theta}\right)$
$\qquad-\int_{X}\frac{S_{\mathrm{v}}(\phi)}{\mathrm{w}(m_{\phi})}-\ell_{\mathrm{ext}}(m_{\phi})-r\left(\frac{\mathrm{tr}_{\mathrm{v},\phi}(\theta)}{\mathrm{w}(m_{\phi})}-\underline{\theta}\right)\omega_{\phi_{0}}^{n}$
for some $r>>0$. Its linearization is
${\mathcal{L}}_{0,\phi_{0}}(u):=(D_{\phi_{0}}{\mathcal{F}}_{0})(u)=-\frac{(\mathcal{D}^{*}_{\phi_{0}}\mathrm{v}(m_{\phi_{0}})\mathcal{D}_{\phi_{0}})u}{\mathrm{w}(m_{\phi_{0}})}+rF_{\phi_{0}}(u),$
which is a $(\cdot,\cdot)_{\mathrm{w},\phi}$-self-adjoint $K$-invariant linear operator with a trivial kernal (see Lemma 5.4).
By the detailed computation of the linearized operators in Appendix 5.B, we find that $F_{\phi_{0}}$ has a negative first eigenvalue. Thus, for $r$ sufficiently large, the first eigenvalue $\lambda_{1,r}$ of ${\mathcal{L}}_{0,\phi_{0}}$ is always negative.
At the end, the implicit function theorem argument in [Has19] implies that there exists a solution $\phi\in{\mathcal{H}}_{K}$ of ${\mathcal{F}}_{0}(\phi)=0$ for $r$ large enough.
Combining the two cases of $t_{0}=0$ and $10<t_{0}<$ 1, we obtain the following openness property.
Theorem 4.1.
Suppose ${\mathcal{F}}(\phi_{0},t_{0})=0$ for some $t_{0}\in[0,1)$. Then for any $t\in[0,1)$ that is sufficiently close to $t_{0}$, there exists $\phi(t)\in{\mathcal{H}}^{4,\alpha}_{K}$ such that ${\mathcal{F}}(\phi(t),t)=0$.
4.B. Closedness
In this subsection, we will always assume the weighted Mabuchi functional ${\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}}}$ is $\mathbb{G}$-coercive. Recall that ${\mathbb{G}}=K_{\mathbb{C}}$ is a reductive Lie subgroup of ${\rm Aut}_{T}(X)$ and $K$ contains a maximal torus, and ${\mathbb{T}}$ is the center of ${\mathbb{G}}$.
Proposition 4.2.
  1. (1)
    The weighted Mabuchi functional ${\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}}}$ is bounded from below. Furthermore, ${\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}},t}$ is coercive for $10<t<$ 1.
  2. (2)
    The weighted Mabuchi functional ${\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}}}$ is invariant under the group action $\sigma\in{\mathbb{G}}$.
Proof.
  1. (1)
    The lower bound is a direct consequence of ${\mathbb{G}}$-coercivity.
    For $t\in[0,1)$, by the definition of ${\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}},t}$, it suffices to show that ${\mathbb{J}}_{\mathrm{v},\mathrm{w}}^{\theta}$ is coercive. We take a suitable choice of $\theta:=\omega+\sqrt{-1}\partial\overline{\partial}\psi_{0}$ such that
    $\frac{\mathrm{tr}_{\mathrm{v},\omega}\theta}{\mathrm{w}(m_{\omega})}=\underline{\theta}.$
    Then $\theta\geq\varepsilon\omega$ for some $\varepsilon>0$. Recall that for any $\phi\in{\mathcal{H}}_{K,0}$,
    ${\mathbb{J}}^{\omega}_{\mathrm{v},\mathrm{w}}(\phi)=({\mathbb{I}}_{\mathrm{v}}-{\mathbb{J}}_{\mathrm{v}})(\phi)-\underline{\theta}{\mathbb{E}}_{\mathrm{w}}(\phi)\geq C\cdot{\mathbb{J}}_{\mathrm{v}}(\phi)-C$
    Thus we have
    ${\mathbb{J}}_{\mathrm{v},\mathrm{w}}^{\theta}(\phi)\geq\varepsilon{\mathbb{J}}_{\mathrm{v},\mathrm{w}}^{\omega}(\phi)\geq C\epsilon\cdot{\mathbb{J}}_{\mathrm{v}}(\phi)-C,$
    which shows that ${\mathbb{J}}^{\theta}_{\mathrm{v},\mathrm{w}}$ is coercive.
  2. (2)
    Let $\sigma\in{\mathbb{T}}$, then there exists a holomorphic vector field $\xi$ which generates a one-parameter subgroup $\{\sigma(t)\}_{t\in{\mathbb{R}}}$ such that $\sigma(0)={\rm id}$ and $\sigma(1)=\sigma$. By the definition of weighted Mabuchi functional and weighted Futaki invariant (see [Lah19, Section 6]), we have
    $\frac{\mathrm{d}}{\mathrm{d}t}{\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}}}(\sigma(t)^{*}\phi)={\mathbb{F}}_{\mathrm{v},\mathrm{w}}(\xi)=a,$
    where $a$ is a constant and ${\mathbb{F}}_{\mathrm{v},\mathrm{w}}$ is the weighted Futaki invariant. Since ${\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}}}$ is bounded from below, then $a=0$. It implies ${\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}}}(\sigma^{*}\phi)={\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}}}(\phi)$.
Proposition 4.3.
Suppose ${\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}},t}$ is coercive. Then there exists a solution $\varphi_{t}$ to (1.3), which is also the minimizer of ${\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}},t}$ for $t\leq 1$.
Proof.
This corresponds to the case ${\rm Aut}_{0}(X)=0$ in [CC21b, Section 4.1]. We have the higher order estimate of equations (2.13a), (2.13b) (see Theorem 5.2 and Corollary 5.3). Exactly, using the standard conitinuity method in [CC21b, Section 4.1] implies the existence of the solution to (2.13a), (2.13b). ∎
Proposition 4.4.
There exists a sequence $\{t_{i}\}_{i\in\mathbb{N}}$, $t_{i}\to 1$, satisfying
${\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}},t_{i}}(\varphi_{i})\to\inf_{\phi\in{\mathcal{E}}^{1}_{{K}}}{\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}}}(\phi)$ (4.4)
and
${\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}}}(\varphi_{i})\rightarrow\inf_{\phi\in{\mathcal{E}}^{1}_{{K}}}{\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}}}(\phi)$ (4.5)
where $\varphi_{i}=\varphi_{t_{i}}$ is the solution of (2.13a), (2.13b) for $t=t_{i}<1$.
Proof.
By Proposition 4.3, one has ${\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}},t_{i}}(\varphi_{i})=\inf_{{\mathcal{E}}^{1}_{K}}{\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}},t_{i}}$. By Proposition 4.2 (1), $\inf_{\phi\in{\mathcal{E}}^{1}_{{K}}}{\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}}}(\phi)>-\infty$. Let $\varphi^{\varepsilon}\in{\mathcal{E}}^{1}_{K}$ such that ${\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}}}(\varphi^{\varepsilon})\leq\inf_{\phi\in{\mathcal{E}}^{1}_{{K}}}{\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}}}(\phi)+\varepsilon$. We obtain
$\limsup_{i\rightarrow\infty}{\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}},t_{i}}(\varphi_{i})\leq\limsup_{i\rightarrow\infty}{\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}},t_{i}}(\varphi^{\varepsilon})={\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}}}(\varphi^{\varepsilon})\leq\inf_{\phi\in{\mathcal{E}}^{1}_{{K}}}{\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}}}(\phi)+\varepsilon.$ (4.6)
On the other hand,
${\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}},t_{i}}(\varphi_{i})={\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}}}(\varphi_{i})+\frac{1-t_{i}}{t_{i}}\cdot{\mathbb{J}}^{\theta}_{\mathrm{v},\mathrm{w}}(\varphi_{i})\geq\inf_{\phi\in{\mathcal{E}}^{1}_{{K}}}{\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}}}(\phi)+\frac{1-t_{i}}{t_{i}}{\mathbb{J}}_{\mathrm{v},\mathrm{w}}^{\theta}(0),$
where we take a suitable choice of $\theta:=\omega+\sqrt{-1}\partial\overline{\partial}\psi_{0}$ such that
$\frac{\mathrm{tr}_{\mathrm{v},\omega}\theta}{\mathrm{w}(m_{\omega})}=\underline{\theta},$
of which the solution $\phi=0$ is the minimizer of ${\mathbb{J}}_{\mathrm{v},\mathrm{w}}^{\theta}(\cdot)$. Then we obtain
$\liminf_{i\rightarrow\infty}{\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}},t_{i}}(\varphi_{i})\geq\inf_{\phi\in{\mathcal{E}}^{1}_{{K}}}{\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}}}(\phi).$
Together with (4.6), we obtain (4.4).
The remaining part is the (4.5). By (4.6), we have
$\inf_{\phi\in{\mathcal{E}}^{1}_{{K}}}{\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}}}(\phi)+\varepsilon\geq$ ${\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}}}(\varphi_{i})+\frac{1-t_{i}}{t_{i}}\cdot{\mathbb{J}}^{\theta}_{\mathrm{v},\mathrm{w}}(\varphi_{i})$
$\geq$ ${\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}}}(\varphi_{i})+\frac{1-t_{i}}{t_{i}}{\mathbb{J}}_{\mathrm{v},\mathrm{w}}^{\theta}(0).$
Taking the limsup on the right hand side, we obtain (4.5). ∎
Corollary 4.5.
There exists a constant $C>0$ such that ${\mathbb{J}}_{\mathrm{v},\mathbb{T}}(\varphi_{i})<C$ for any $i\in\mathbb{N}$.
Proof.
It follows from the assumption of ${\mathbb{G}}$-coercivity of weighted $K$-energy and (4.5). ∎
Let $\phi_{i}=\sigma_{i}^{*}\varphi_{i}$, such that ${\mathbb{J}}_{\mathrm{v}}(\phi_{i})={\mathbb{J}}_{\mathrm{v},\mathbb{T}}(\varphi_{i})$.
Lemma 4.6.
There exists a constant $C>0$, such that ${\mathbb{H}}_{\mathrm{v}}(\phi_{i})<C$ for any $i\in\mathbb{N}$.
Proof.
Due to the ${\mathbb{G}}$-invariance of weighted $K$-energy as in Proposition 4.2 (2) and (4.5), we have
$\sup_{i}{\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}}}(\phi_{i})=\sup_{i}{\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}}}(\varphi_{i})<\infty.$
One has
$\sup_{i}|{\mathbb{J}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}}}^{\mathrm{Ric}}(\phi_{i})|\leq C|\mathrm{Ric}|_{\omega}\sup_{i}{\mathbb{J}}_{\mathrm{v}}(\phi_{i})<\infty.$
By the formula of weighted $K$-energy, we finish the proof. ∎
Lemma 4.7.
Let $u_{i}$ be the potential function which satisfies $\sigma_{i}^{*}\theta=\theta+{\rm dd^{c}}u_{i}$, $\sup_{X}u_{i}=0$. Then
$\mathrm{v}(m_{\phi_{i}})\omega_{\phi_{i}}^{n}=$ $e^{\widetilde{F}_{i}}\omega^{n},$ (4.7)
$\left(\Delta_{\phi_{i}}-\frac{\mathrm{v}_{,\alpha}(m_{\phi_{i}})}{\mathrm{v}(m_{\phi_{i}})}J\xi^{\alpha}\right)\widetilde{F}_{i}=$ $-\ell_{\mathrm{ext}}(m_{\phi_{i}})\frac{\mathrm{w}(m_{\phi_{i}})}{\mathrm{v}(m_{\phi_{i}})}+(1-\frac{1}{t})\frac{\mathrm{tr}_{\mathrm{v},\phi_{i}}(\theta_{u_{i}})}{\mathrm{v}(m_{\phi_{i}})}$ (4.8)
$-(1-\frac{1}{t})\underline{\theta}\frac{\mathrm{w}(m_{\phi_{i}})}{\mathrm{v}(m_{\phi_{i}})}+\frac{\mathrm{tr}_{\mathrm{v},\phi_{i}}\mathrm{Ric}(\omega)}{\mathrm{v}(m_{\phi_{i}})}.$
Proof.
Recall that $\phi_{i}=\sigma_{i}^{*}\varphi_{i}$. Then
$\sigma_{i}^{*}(\omega_{\varphi_{i}}^{n})=(\sigma_{i}^{*}\omega_{\varphi_{i}})^{n}=\omega_{\phi_{i}}^{n},\quad\sigma_{i}^{*}(\mathrm{v}(m_{\varphi_{i}})\omega_{\varphi_{i}}^{n})=\mathrm{v}(m_{\phi_{i}})\omega_{\phi_{i}}^{n}$
and
$\sigma_{i}^{*}(e^{F_{i}}\omega^{n})=e^{F_{i}\circ\sigma_{i}}(\sigma_{i}^{*}\omega)^{n}.$
Taking pullback in the both sides of the first equation (2.13a), we obtain
$\mathrm{v}(m_{\phi_{i}})\omega_{\phi_{i}}^{n}=\sigma_{i}^{*}(e^{F_{i}}\omega^{n})=e^{\widetilde{F}_{i}}\omega^{n}$
where $\widetilde{F}_{i}=F_{i}\circ\sigma_{i}+\log(\frac{\sigma_{i}^{*}\omega^{n}}{\omega^{n}})$.
Take $\sigma_{i}^{*}$-action on the second equation (2.13b), one has
[…]
and
[…]
On the other hand,
$\mathrm{tr}_{\mathrm{v},\phi_{i}}(\mathrm{Ric}(\sigma_{i}^{*}\omega)-\mathrm{Ric}(\omega))$
$=$ $\mathrm{v}(m_{\phi_{i}})\mathrm{tr}_{\phi_{i}}\left(\sqrt{-1}\partial\overline{\partial}\log\frac{\omega^{n}}{\sigma_{i}^{*}\omega^{n}}\right)+\langle\mathrm{d}\mathrm{v}(m_{\phi_{i}}),m_{\mathrm{Ric}(\sigma_{i}^{*}\omega)}-m_{\mathrm{Ric}(\omega)}\rangle$
$=$ $\mathrm{v}(m_{\phi_{i}})\Delta_{\phi_{i}}\log\frac{\omega^{n}}{\sigma_{i}^{*}\omega^{n}}+\mathrm{v}_{,\alpha}(m_{\phi_{i}})J\xi^{\alpha}\left(\log\frac{\omega^{n}}{\sigma_{i}^{*}\omega^{n}}\right),$ (4.11)
where we have used (2.11) for the last equality. Compare (4.9), (4.10) and (4.11), we derive (4.8). ∎
Proposition 4.8.
Let $f_{i}=\frac{1-t_{i}}{t_{i}}u_{i}$. Let $\alpha=\alpha(X,\omega)$ be the alpha-invariant of the Kähler class$[\omega]$. Then there exists $C>0$ such that $\int_{X}e^{-p\cdot f_{i}}\omega^{n}<C$ for any $p<\frac{t_{i}\alpha}{1-t_{i}}$. Moreover, we have $e^{-f_{i}}\rightarrow 1$ in $L^{p}(\omega^{n})$ as $t_{i}\rightarrow 1$ for any $p<\infty$.
Proof.
Given $p>1$, suppose $t_{i}$ is sufficiently close to $1$ such that $p\frac{1-t_{i}}{t_{i}}<\alpha$, due to Tian’s $\alpha$-invariant, then
$\int_{X}e^{-pf_{i}}\omega^{n}=\int_{X}e^{-p\frac{1-t_{i}}{t_{i}}u_{i}}<C.$
By (4.4), we have
$\inf_{\phi\in{\mathcal{E}}^{1}_{{K}}}{\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}}}(\phi)+\varepsilon_{i}\geq$ ${\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}},t_{i}}(\varphi_{i})={\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}}}(\varphi_{i})+\frac{1-t_{i}}{t_{i}}\cdot{\mathbb{J}}^{\theta}_{\mathrm{v},\mathrm{w}}(\varphi_{i})$
$\geq$ $\inf_{\phi\in{\mathcal{E}}^{1}_{{K}}}{\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}}}(\phi)+\frac{1-t_{i}}{t_{i}}{\mathbb{J}}_{\mathrm{v}}(\varphi_{i}).$
It follows that $(1-t_{i}){\mathbb{J}}_{\mathrm{v}}(\varphi_{i})\rightarrow 0$. In particular, $\int_{X}|f_{i}|\omega^{n}\rightarrow 0$. Then for any $\epsilon>0$, for $i$ sufficiently large, there exists $A_{\epsilon}\subset X$, such that $|f_{i}|\leq\epsilon$ on $A_{\epsilon}$, and $Vol_{\omega}(X\setminus A_{\epsilon})\leq\epsilon$. Then
$\int_{X}|e^{-f_{i}}-1|^{p}\omega^{n}$ $\leq C\cdot\big{(}\epsilon^{p}\cdot Vol_{\omega}(X)+\int_{X\setminus A_{\epsilon}}1+e^{-pf_{i}}\omega^{n}\big{)}$
$\leq C\big{(}\epsilon+(\int_{X\setminus A_{\epsilon}}e^{-pqf_{i}}\omega^{n})^{\frac{1}{q}}\cdot\epsilon^{\frac{1}{q^{*}}}\big{)}$
$\leq C\epsilon^{\frac{1}{q^{*}}}.$
Then $e^{-f_{i}}\rightarrow 1$ in $L^{p}(\omega^{n})$ as $t_{i}\rightarrow 1$ for any $p<\infty$. ∎
Applying Theorem 3.7 to equations (4.7), (4.8), we obtain
Proposition 4.9.
By apriori estimates, $\phi_{i}\in W^{2,p},\widetilde{F}_{i}+f_{i}-\log(\mathrm{v}(m_{\phi_{i}}))\in W^{1,p}$.
Hence we may take a subsequence of $\phi_{i}$ (without relabeled), and a function $\phi_{*}\in W^{2,p}$ for any $p<\infty$, and another function $F_{*}\in W^{1,p}$ for any $p<\infty$, such that
$\phi_{i}\rightarrow\phi_{*}\text{ in }C^{1,\alpha}\text{ for any }0<\alpha<1\text{ and }\sqrt{-1}\partial\overline{\partial}\phi_{i}\rightarrow\sqrt{-1}\partial\overline{\partial}\phi_{*}\text{ weakly in }L^{p},$ (4.12)
$\widetilde{F}_{i}-\log\mathrm{v}(m_{\phi_{i}})+f_{i}\rightarrow F_{*}\text{ in }C^{\alpha}\text{ for any }0<\alpha<1\text{ and }$ (4.13)
$\qquad\qquad\qquad\qquad\nabla(\widetilde{F}_{i}-\log\mathrm{v}(m_{\phi_{i}})+f_{i})\rightarrow\nabla F_{*}\text{ weakly in }L^{p}.$
It follows from [CC21b, Lemma 4.22] that
$\omega_{\phi_{i}}^{k}\rightarrow\omega_{\phi_{i}}^{k}\text{ weakly in }L^{p}\text{ for any }1\leq k\leq n\text{ and }p<\infty.$
Proposition 4.10.
Let $\phi_{*}$, $F_{*}$ be the limits as above. Then $\phi_{*}$ is a weak solution of the weighted cscK equation in the following sense:
  1. (1)
    $\omega_{\phi_{*}}^{n}=e^{F_{*}}\omega^{n}$,
  2. (2)
    For any $u\in C^{\infty}(X)$, we have
    $\int_{X}F_{*}\sqrt{-1}\partial\overline{\partial}u\wedge\omega_{\phi_{*}}^{n-1}=$ $\int_{X}u\frac{\mathrm{v}_{,\alpha}(m_{\phi_{*}})}{\mathrm{v}(m_{\phi_{*}})}J\xi^{\alpha}(F_{*})\omega_{\phi_{*}}^{n}-\int_{X}\log\mathrm{v}(m_{\phi_{*}})\sqrt{-1}\partial\overline{\partial}u\wedge\omega_{\phi_{*}}^{n-1}$
    $-\int_{X}u\left(\frac{\mathrm{v}_{,\alpha}(m_{\phi_{*}})\mathrm{v}_{,\beta}(m_{\phi_{*}})}{\mathrm{v}(m_{\phi_{*}})^{2}}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi_{*}}+\ell_{\mathrm{ext}}(m_{\phi_{*}})\frac{\mathrm{w}(m_{\phi_{*}})}{\mathrm{v}(m_{\phi_{*}})}\right)\omega_{\phi_{*}}^{n}$
    $+\int_{X}u\mathrm{Ric}(\omega)\wedge\omega_{\phi_{*}}^{n-1}-\int_{X}u\frac{\mathrm{v}_{,\alpha}(m_{\phi_{*}})}{\mathrm{v}(m_{\phi_{*}})}\Delta_{\omega}m_{\omega}^{\xi^{\alpha}}\omega_{\phi_{*}}^{n}.$ (4.14)
Proof.
For each $i$, we have $\omega_{\phi_{i}}^{n}=e^{\widetilde{F}_{i}-\log\mathrm{v}(m_{\phi_{i}})}\omega^{n}$. By (4.12), we have $\omega_{\phi_{i}}^{n}\rightarrow\omega_{\phi_{*}}^{n}$ converge weakly in $L^{p}$ for any $p<\infty$. Note that $e^{\widetilde{F}_{i}-\log\mathrm{v}(m_{\phi_{i}})}\omega^{n}=e^{\widetilde{F}_{i}-\log\mathrm{v}(m_{\phi_{i}})+f_{i}}e^{-f_{i}}\omega^{n}$. Since $\widetilde{F}_{i}-\log\mathrm{v}(m_{\phi_{i}})+f_{i}$ is uniformly bounded and converges to $F_{*}$, then $e^{F_{i}-\log\mathrm{v}(m_{\phi_{i}})+f_{i}}\rightarrow e^{F_{*}}$ in $L^{p}$ for any $p<\infty$. On the other hand, by Proposition 4.8, one has $e^{-f_{i}}\rightarrow 1$ in $L^{p}$ for any $p<\infty$. Thus, $e^{\widetilde{F}_{i}-\log\mathrm{v}(m_{\phi_{i}})}\rightarrow e^{F_{*}}$ in $L^{p}$ for any $p<\infty$. We obtain the first equation $\omega_{\phi_{*}}^{n}=e^{F_{*}}\omega^{n}$.
To obtain the second equation, by (2.10), we first re-write the equation (4.8) to
$\Delta_{\phi_{i}}(\widetilde{F}_{i}+f_{i}-\log\mathrm{v}(m_{\phi_{i}}))$
$=$ $\Delta_{\phi_{i}}(\widetilde{F}_{i}+f_{i})-\Delta_{\phi_{i}}\log\mathrm{v}(m_{\phi_{i}})$
$=$ $\frac{\mathrm{v}_{,\alpha}(m_{\phi_{i}})}{\mathrm{v}(m_{\phi_{i}})}J\xi^{\alpha}(\widetilde{F}_{i}+f_{i}-\log\mathrm{v}(m_{\phi_{i}}))+\frac{\mathrm{v}_{,\alpha}(m_{\phi_{i}})}{\mathrm{v}(m_{\phi_{i}})}J\xi^{\alpha}(\log\mathrm{v}(m_{\phi_{i}}))-\Delta_{\phi_{i}}\log\mathrm{v}(m_{\phi_{i}})$ (4.15)
$-\ell_{\mathrm{ext}}(m_{\phi_{i}})\frac{\mathrm{w}(m_{\phi_{i}})}{\mathrm{v}(m_{\phi_{i}})}+(1-\frac{1}{t_{i}})\frac{\mathrm{tr}_{\mathrm{v},\phi_{i}}(\theta)}{\mathrm{v}(m_{\phi_{i}})}-(1-\frac{1}{t_{i}})\underline{\theta}\frac{\mathrm{w}(m_{\phi_{i}})}{\mathrm{v}(m_{\phi_{i}})}+\frac{\mathrm{tr}_{\mathrm{v},\phi_{i}}\mathrm{Ric}(\omega)}{\mathrm{v}(m_{\phi_{i}})}.$
Denote $W_{i}=\widetilde{F}_{i}+f_{i}-\log\mathrm{v}(m_{\phi_{i}})$. For any $u\in C^{\infty}(X)$, one has
$\int_{X}W_{i}\sqrt{-1}\partial\overline{\partial}u\wedge\omega_{\phi_{i}}^{n-1}$
$=$ $\int_{X}u\frac{\mathrm{v}_{,\alpha}(m_{\phi_{i}})}{\mathrm{v}(m_{\phi_{i}})}J\xi^{\alpha}(W_{i})\omega_{\phi_{i}}^{n}-\int_{X}\log\mathrm{v}(m_{\phi_{i}})\sqrt{-1}\partial\overline{\partial}u\wedge\omega_{\phi_{i}}^{n-1}$
$-\int_{X}u\left(\frac{\mathrm{v}_{,\alpha}(m_{\phi_{i}})\mathrm{v}_{,\beta}(m_{\phi_{i}})}{\mathrm{v}(m_{\phi_{i}})^{2}}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi_{i}}+\ell_{\mathrm{ext}}(m_{\phi_{i}})\frac{\mathrm{w}(m_{\phi_{i}})}{\mathrm{v}(m_{\phi_{i}})}+(1-\frac{1}{t_{i}})\underline{\theta}\frac{\mathrm{w}(m_{\phi_{i}})}{\mathrm{v}(m_{\phi_{i}})}\right)\omega_{\phi_{i}}^{n}$
$+\int_{X}u\left((1-\frac{1}{t_{i}})\frac{\mathrm{tr}_{\mathrm{v},\phi_{i}}(\theta)}{\mathrm{v}(m_{\phi_{i}})}+\frac{\mathrm{tr}_{\mathrm{v},\phi_{i}}\mathrm{Ric}(\omega)}{\mathrm{v}(m_{\phi_{i}})}\right)\omega_{\phi_{i}}^{n}.$
For the left hand side, since $W_{i}\rightarrow F_{*}$ strongly in $L^{p}$, $\omega_{\phi_{i}}^{n-1}\rightarrow\omega_{\phi_{*}}^{n-1}$ weakly in $L^{p}$ for any $p<\infty$, we obtain
$\int_{X}W_{i}\sqrt{-1}\partial\overline{\partial}u\wedge\omega_{\phi_{i}}^{n-1}\rightarrow\int_{X}F_{*}\sqrt{-1}\partial\overline{\partial}u\wedge\omega_{\phi_{*}}^{n-1}.$
For the right hand side, by (4.13), (4.12), $\phi_{i}\rightarrow\phi_{*}$ in $C^{1,\alpha}$ for any $10<\alpha<$ 1. Then following the construction in the proof of [HL23, Proposition 10], we have
$\frac{\mathrm{v}_{,\alpha}(m_{\phi_{i}})\mathrm{v}_{,\beta}(m_{\phi_{i}})}{\mathrm{v}(m_{\phi_{i}})^{2}}\omega_{\phi_{i}}^{n}\rightarrow\frac{\mathrm{v}_{,\alpha}(m_{\phi_{*}})\mathrm{v}_{,\beta}(m_{\phi_{*}})}{\mathrm{v}(m_{\phi_{*}})^{2}}\omega_{\phi_{*}}^{n}$
converges weakly as signed measures. In addition, $J\xi^{\alpha}(W_{i})\in L^{p}$ and $\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi_{i}}\in L^{p}$ for any $p<\infty$. Thus, we have
$\int_{X}u\frac{\mathrm{v}_{,\alpha}(m_{\phi_{i}})}{\mathrm{v}(m_{\phi_{i}})}J\xi^{\alpha}(W_{i})\omega_{\phi_{i}}^{n}\rightarrow\int_{X}u\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}(F_{*})\omega_{\phi_{*}}^{n}$
and
$-\int_{X}u\left(\frac{\mathrm{v}_{,\alpha}(m_{\phi_{i}})\mathrm{v}_{,\beta}(m_{\phi_{i}})}{\mathrm{v}(m_{\phi_{i}})^{2}}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi_{i}}+\ell_{\mathrm{ext}}(m_{\phi_{i}})\frac{\mathrm{w}(m_{\phi_{i}})}{\mathrm{v}(m_{\phi_{i}})}+(1-\frac{1}{t_{i}})\underline{\theta}\frac{\mathrm{w}(m_{\phi_{i}})}{\mathrm{v}(m_{\phi_{i}})}\right)\omega_{\phi_{i}}^{n}$
$\longrightarrow-\int_{X}u\left(\frac{\mathrm{v}_{,\alpha}(m_{\phi_{*}})\mathrm{v}_{,\beta}(m_{\phi_{*}})}{\mathrm{v}(m_{\phi_{*}})^{2}}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi_{*}}+\ell_{\mathrm{ext}}(m_{\phi_{*}})\frac{\mathrm{w}(m_{\phi_{*}})}{\mathrm{v}(m_{\phi_{*}})}\right)\omega_{\phi_{*}}^{n}.$
We also have
$\int_{X}u\left((1-\frac{1}{t_{i}})\frac{\mathrm{tr}_{\mathrm{v},\phi_{i}}(\theta)}{\mathrm{v}(m_{\phi_{i}})}+\frac{\mathrm{tr}_{\mathrm{v},\phi_{i}}\mathrm{Ric}(\omega)}{\mathrm{v}(m_{\phi_{i}})}\right)\omega_{\phi_{i}}^{n}$
$=$ $\int_{X}u(1-\frac{1}{t_{i}})\theta\wedge\omega_{\phi_{i}}^{n-1}+\int_{X}u\mathrm{Ric}(\omega)\wedge\omega_{\phi_{i}}^{n-1}$
$+\int_{X}u\left((1-\frac{1}{t_{i}})\frac{\langle\mathrm{d}\mathrm{v}(m_{\phi_{i}}),m_{\theta}\rangle}{\mathrm{v}(m_{\phi})}-\frac{\mathrm{v}_{,\alpha}(m_{\phi_{i}})}{\mathrm{v}(m_{\phi_{i}})}\Delta_{\omega}m_{\omega}^{\xi^{\alpha}}\right)\omega_{\phi_{i}}^{n}$
$\longrightarrow\int_{X}u\mathrm{Ric}(\omega)\wedge\omega_{\phi_{*}}^{n-1}-\int_{X}u\frac{\mathrm{v}_{,\alpha}(m_{\phi_{*}})}{\mathrm{v}(m_{\phi_{*}})}\Delta_{\omega}m_{\omega}^{\xi^{\alpha}}\omega_{\phi_{*}}^{n}$
and
$\int_{X}\log\mathrm{v}(m_{\phi_{i}})\sqrt{-1}\partial\overline{\partial}u\wedge\omega_{\phi_{i}}^{n-1}\rightarrow\int_{X}\log\mathrm{v}(m_{\phi_{*}})\sqrt{-1}\partial\overline{\partial}u\wedge\omega_{\phi_{*}}^{n-1}.$
Thus, we obtain (2). ∎
Lemma 4.11.
There exists a constant $C>0$ such that $C^{-1}\omega\leq\omega_{\phi_{*}}\leq C\omega$.
Proof.
We have obtained that $F_{*}\in W^{1,p}$ for any $p<\infty$, then we may find a sequence $G_{k}\in C^{\infty}(X)$, uniformly bounded, and $G_{k}\rightarrow F_{*}$ in $W^{1,p}$. Let $\psi_{k}$ be the solution of $\omega_{\psi_{k}}^{n}=e^{G_{k}}\omega^{n}$ with $\sup_{X}\psi_{k}=0$. By [CH12, Theorem 1.1], for any $p<\infty$, one has
$\sup_{k}\|\psi_{k}\|_{W^{3,p}}<\infty.$
Up to a subsequence, we have that there exists $\psi_{*}\in W^{3,p}$ such that $\psi_{k}\rightarrow\psi_{*}$ in $W^{2,p}$ for any finite $p$. It follows that $\omega_{\psi_{*}}^{n}=e^{F_{*}}\omega^{n}$. By uniqueness of Monge-Ampère equations, we obtain that $\phi_{*}=\psi_{*}+c$ for some constant $c$. Hence $\omega_{\phi_{*}}=\omega_{\psi_{*}}\leq C\omega$ and $\omega_{\phi_{*}}=\omega_{\psi_{*}}\geq C^{-1}\omega$ follows from $F_{*}$ is bounded from below. ∎
Corollary 4.12.
$\phi_{*}$ is the smooth solution to the weighted cscK equation (1.1).
Proof.
In the proof of Lemma 4.11, we have obtain that $\phi_{*}\in W^{3,p}$ for any $p<\infty$. Hence $\omega_{\phi_{*}}\in C^{\alpha}$ for any $10<\alpha<$ 1. By (4.14) and Schauder estimate, we derive $F_{*}\in C^{2,\alpha}$ for any $10<\alpha<$ 1. Then the higher regularity follows from bootstrap (see Corollary 5.3). ∎
5. Appendix
5.A. Apriori estimates for weighted cscK equation
In this subsection, we provide the apriori estimates required to show that the weighted-cscK metric is smooth.
We consider the weighted cscK equations (1.2a), (1.2b). Applying Theorem 3.5 with $f_{*}=0$ induces
Proposition 5.1.
Assume $\mathrm{v}$ is log-concave. Let $\phi$ be a solution of (1.2a), (1.2b), for any $p\geq 1$, then
$\|\mathrm{tr}\omega_{\phi}\|_{L^{p}(\omega^{n})}\leq C=C(n,p,\mathrm{v},\mathrm{w},\ell_{\mathrm{ext}},A_{1},B,C_{0}),$
where $A_{1}$, $B$, $C_{0}$ are constant
$\mathrm{Bisec}(\omega)\geq-B,\quad\mathrm{Ric}(\omega)\leq A_{1}\omega,\quad\|F\|_{L^{\infty}}+\|\phi\|_{L^{\infty}}\leq C_{0}.$
Adapting the similar argument in the proof of Theorem 3.7, we obtain the gradient estimate of $F$ and $C^{2}$-estimate of $\phi$ for equations (1.2a), (1.2b).
Theorem 5.2.
Assume $\mathrm{v}$ is log-concave. Let $\phi$ be a solution of (1.2a), (1.2b), then
$\max_{X}(|\nabla^{\phi}F|_{\phi}+\mathrm{tr}\omega_{\phi})\leq C,$
where $C$ is a constant depending on $\omega$, $\mathrm{v}$, $\mathrm{w}$, $\|\phi\|_{C^{0}}$, $\|F\|_{C^{0}}$, $\|\mathrm{tr}\omega_{\phi}\|_{L^{p}}$ for some $p$.
Proof.
Comparing with the setting in the proof of Theorem 3.7, $f_{*}=0$. Applying $f_{*}=0$ in (3.16), we have
$\Delta_{\phi}(e^{-\log\mathrm{v}(m_{\phi})}\cdot e^{\frac{F^{\prime}}{2}}|\nabla^{\phi}F^{\prime}|_{\phi}^{2})$
$\geq$ $e^{\frac{F^{\prime}}{2}-\log\mathrm{v}(m_{\phi})}|\nabla^{\phi}F^{\prime}|_{\phi}^{2}|\nabla^{\phi}\log\mathrm{v}(m_{\phi})|_{\phi}^{2}-2e^{-\log\mathrm{v}(m_{\phi})}\nabla^{\phi}\log\mathrm{v}(m_{\phi})\cdot_{\phi}\nabla^{\phi}(e^{\frac{F^{\prime}}{2}}|\nabla^{\phi}F^{\prime}|_{\phi}^{2})$
$+e^{\frac{F^{\prime}}{2}-\log\mathrm{v}(m_{\phi})}2\nabla^{\phi}F^{\prime}\cdot_{\phi}\nabla^{\phi}\Delta_{\phi}F^{\prime}-C_{\mathrm{v},\xi}(1+(\mathrm{tr}\omega_{\phi})^{n-1})e^{\frac{F^{\prime}}{2}-\log\mathrm{v}(m_{\phi})}|\nabla^{\phi}F^{\prime}|_{\phi}^{2}$
$+\frac{|F^{\prime}_{i\overline{j}}|^{2}}{(1+\phi_{i\overline{i}})(1+\phi_{j\overline{j}})}e^{\frac{F^{\prime}}{2}-\log\mathrm{v}(m_{\phi})}+2\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}(F^{\prime})e^{\frac{F^{\prime}}{2}-\log\mathrm{v}(m_{\phi})}|\nabla^{\phi}F^{\prime}|_{\phi}^{2}.$ (5.1)
By [CC21a, formula (3.6)],
$\Delta_{\phi}\mathrm{tr}\omega_{\phi}\geq-C(\mathrm{tr}\omega_{\phi})^{n}+\frac{\sum_{i}|\phi_{p\overline{q}i}|^{2}}{(1+\phi_{p\overline{p}})(1+\phi_{q\overline{q}})}+\Delta F^{\prime}-R(g).$
By Young inequality,
$|\Delta F^{\prime}|\leq$ $\frac{\delta}{2}\frac{|F^{\prime}_{i\overline{i}}|^{2}}{(1+\phi_{i\overline{i}})^{2}}+(2\delta)^{-1}(1+\phi_{i\overline{i}})^{2}$
$\leq$ $\frac{\delta}{2}\frac{|F^{\prime}_{i\overline{i}}|^{2}}{(1+\phi_{i\overline{i}})^{2}}+n(2\delta)^{-1}(\mathrm{tr}\omega_{\phi})^{2}.$
Note that, by the arithmetic-geometric inequality, we have the lower bound
$\mathrm{tr}\omega_{\phi}\geq ne^{\frac{F^{\prime}}{n}}\geq ne^{-\frac{\|F^{\prime}\|_{C_{0}}}{n}}>0.$
Then there exists a constant $C$ such that $1\leq C(\mathrm{tr}\omega_{\phi})^{n-1}$. Now, by taking a suitable constant $\delta$ and (5.1), one has
$\Delta_{\phi}(e^{-\log\mathrm{v}(m_{\phi})}\cdot e^{\frac{F^{\prime}}{2}}|\nabla^{\phi}F^{\prime}|_{\phi}^{2}+\mathrm{tr}\omega_{\phi})$
$\geq$ $e^{\frac{F^{\prime}}{2}-\log\mathrm{v}(m_{\phi})}|\nabla^{\phi}F^{\prime}|_{\phi}^{2}|\nabla^{\phi}\log\mathrm{v}(m_{\phi})|_{\phi}^{2}-2e^{-\log\mathrm{v}(m_{\phi})}\nabla^{\phi}\log\mathrm{v}(m_{\phi})\cdot_{\phi}\nabla^{\phi}(e^{\frac{F^{\prime}}{2}}|\nabla^{\phi}F^{\prime}|_{\phi}^{2})$
$+e^{\frac{F^{\prime}}{2}-\log\mathrm{v}(m_{\phi})}2\nabla^{\phi}F^{\prime}\cdot_{\phi}\nabla^{\phi}\Delta_{\phi}F^{\prime}-C_{\mathrm{v},\xi}(\mathrm{tr}\omega_{\phi})^{n-1}(e^{\frac{F^{\prime}}{2}-\log\mathrm{v}(m_{\phi})}|\nabla^{\phi}F^{\prime}|_{\phi}^{2}+\mathrm{tr}\omega_{\phi})$
$+2\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}(F^{\prime})(e^{\frac{F^{\prime}}{2}-\log\mathrm{v}(m_{\phi})}|\nabla^{\phi}F^{\prime}|_{\phi}^{2}+\mathrm{tr}\omega_{\phi})-2\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}(F^{\prime})\mathrm{tr}\omega_{\phi}.$
Set
$u=e^{\frac{F^{\prime}}{2}-\log\mathrm{v}(m_{\phi})}|\nabla^{\phi}F^{\prime}|_{\phi}^{2}+\mathrm{tr}\omega_{\phi}+1,$
then we have
$\Delta_{\phi}u\geq$ $-C_{\mathrm{v},\xi}(\mathrm{tr}\omega_{\phi})^{n-1}u+2e^{\frac{F^{\prime}}{2}-\log\mathrm{v}(m_{\phi})}\nabla^{\phi}F^{\prime}\cdot_{\phi}\nabla^{\phi}\Delta_{\phi}F^{\prime}+2\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}(F^{\prime})u$
$+e^{\frac{F^{\prime}}{2}-\log\mathrm{v}(m_{\phi})}|\nabla^{\phi}F^{\prime}|_{\phi}^{2}|\nabla^{\phi}\log\mathrm{v}(m_{\phi})|_{\phi}^{2}-2e^{-\log\mathrm{v}(m_{\phi})}\nabla^{\phi}\log\mathrm{v}(m_{\phi})\cdot_{\phi}\nabla^{\phi}(e^{\frac{F^{\prime}}{2}}|\nabla^{\phi}F^{\prime}|_{\phi}^{2}).$ (5.2)
For any $p>0$,
$\frac{1}{2p+1}\Delta_{\phi}(u^{2p+1})=u^{2p}\Delta_{\phi}u+2pu^{2p-1}|\nabla^{\phi}u|_{\phi}^{2}.$
Taking the integration, one has
$\int_{X}2pu^{2p-1}|\nabla^{\phi}u|_{\phi}^{2}e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}$
$=$ $\int_{X}u^{2p}(-\Delta_{\phi}u)e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}+\frac{1}{2p+1}\int_{X}\Delta_{\phi}(u^{2p+1})e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}.$ (5.3)
On the one hand, by (3.C),
$\frac{1}{2p+1}\int_{X}\Delta_{\phi}(u^{2p+1})e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}=$ $-2\int_{X}u^{2p}\nabla^{\phi}(e^{\frac{F^{\prime}}{2}}|\nabla^{\phi}F^{\prime}|_{\phi}^{2})\cdot_{\phi}\nabla^{\phi}\log\mathrm{v}(m_{\phi})e^{\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}$
$+2\int_{X}u^{2p}e^{\frac{F^{\prime}}{2}}|\nabla^{\phi}F^{\prime}|_{\phi}^{2}|\nabla^{\phi}\log\mathrm{v}(m_{\phi})|_{\phi}^{2}e^{\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}$
$-2\int_{X}u^{2p}\nabla^{\phi}\mathrm{tr}\omega_{\phi}\cdot_{\phi}\nabla^{\phi}\log\mathrm{v}(m_{\phi})e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}.$ (5.4)
We estimate the third term on the right hand side of (5.A)
$-2\int_{X}u^{2p}\nabla^{\phi}\mathrm{tr}\omega_{\phi}\cdot_{\phi}\nabla^{\phi}\log\mathrm{v}(m_{\phi})e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}$
$=$ $2\int_{X}2pu^{2p-1}\mathrm{tr}\omega_{\phi}\nabla^{\phi}u\cdot_{\phi}\nabla^{\phi}\log\mathrm{v}(m_{\phi})e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}$
$+2\int_{X}u^{2p}\mathrm{tr}\omega_{\phi}\Delta_{\phi}\log\mathrm{v}(m_{\phi})e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}+4\int_{X}u^{2p}\mathrm{tr}\omega_{\phi}|\nabla^{\phi}\log\mathrm{v}(m_{\phi})|_{\phi}^{2}e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}$
$\leq$ $\frac{p}{2}\int_{X}u^{2p-1}|\nabla^{\phi}u|_{\phi}^{2}e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}+\int_{X}8pu^{2p-1}|\nabla^{\phi}\log\mathrm{v}(m_{\phi})|_{\phi}^{2}(\mathrm{tr}\omega_{\phi})^{2}e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}$
$+2\int_{X}u^{2p}\mathrm{tr}\omega_{\phi}\left(-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}(F^{\prime})+\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}\Delta_{\omega}(m_{\omega}^{\xi^{\alpha}})\right.$
$\left.+\frac{\mathrm{v}_{,\alpha\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}-\frac{\mathrm{v}_{,\alpha}(m_{\phi})\mathrm{v}_{,\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})^{2}}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}\right)e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}$
$+4\int_{X}u^{2p}\mathrm{tr}\omega_{\phi}|\nabla^{\phi}\log\mathrm{v}(m_{\phi})|_{\phi}^{2}e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}.$
$\leq$ $\frac{p}{2}\int_{X}u^{2p-1}|\nabla^{\phi}u|_{\phi}^{2}e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}+C_{\mathrm{v},\xi}(p+1)\int_{X}u^{2p+1}(\mathrm{tr}\omega_{\phi})^{2n}e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n},$ (5.5)
where we have used (2.16) for the second inequality, (2.18),
$|J\xi^{\alpha}(F^{\prime})|\leq C|\nabla F^{\prime}|\leq C(\mathrm{tr}\omega_{\phi})^{\frac{1}{2}}|\nabla^{\phi}F^{\prime}|\leq C(\mathrm{tr}\omega_{\phi})^{\frac{1}{2}}C\sqrt{u}\leq C(\mathrm{tr}\omega_{\phi})^{\frac{1}{2}}u\leq C(\mathrm{tr}\omega_{\phi})^{n-1}u,$
and (2.9) for the third inequality.
On the other hand, by (5.2),
$\int_{X}u^{2p}(-\Delta_{\phi}u)e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}$
$=$ $C\int_{X}u^{2p+1}(\mathrm{tr}\omega_{\phi})^{n-1}e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}-2\int_{X}u^{2p+1}\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}(F^{\prime})e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}$
$-\int_{X}u^{2p}e^{\frac{F^{\prime}}{2}}|\nabla^{\phi}F^{\prime}|_{\phi}^{2}|\nabla^{\phi}\log\mathrm{v}(m_{\phi})|_{\phi}^{2}e^{\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}$
$+2\int_{X}u^{2p}\nabla^{\phi}\log\mathrm{v}(m_{\phi})\cdot_{\phi}\nabla^{\phi}(e^{\frac{F^{\prime}}{2}}|\nabla^{\phi}F^{\prime}|_{\phi}^{2})e^{\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}$
$-2\int_{X}u^{2p}e^{\frac{F^{\prime}}{2}}\nabla^{\phi}F^{\prime}\cdot_{\phi}\nabla^{\phi}\Delta_{\phi}F^{\prime}e^{\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}.$ (5.6)
Applying the same computation as in (3.23), (3.24), (3.25), (3.C) and (3.27) implies
$-2\int_{X}u^{2p}e^{\frac{F^{\prime}}{2}}\nabla^{\phi}F^{\prime}\cdot_{\phi}\nabla^{\phi}\Delta_{\phi}F^{\prime}e^{\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}$
$\leq$ $\int_{X}pu^{2p-1}|\nabla^{\phi}u|_{\phi}^{2}e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}+C(p+1)\int_{X}(\mathrm{tr}\omega_{\phi})^{3(n-1)}u^{2p+1}e^{\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}$
$+2\int_{X}u^{2p+1}\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}(F^{\prime})e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}+\int_{X}u^{2p+1}e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}$ (5.7)
Therefore, by (5.A), (5.A), (5.A), (5.6) together with the above inequality, we deduce
$\int_{X}pu^{2p-1}|\nabla^{\phi}u|_{\phi}^{2}e^{2\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}\leq C(p+1)\int_{X}(\mathrm{tr}\omega_{\phi})^{3(n-1)}u^{2p+1}e^{\log\mathrm{v}(m_{\phi})}\omega_{\phi}^{n}.$
It follows that
$\int_{X}|\nabla^{\phi}(u^{p+\frac{1}{2}})|_{\phi}^{2}\omega^{n}\leq\frac{C_{\mathrm{v},\xi}(p+\frac{1}{2})^{2}(p+1)}{p}\int_{X}C_{\xi}(\mathrm{tr}\omega_{\phi})^{3(n-1)}u^{2p+1}\omega_{\phi}^{n}.$
Then applying the standard Moser iteration steps as in the proof of Theorem 3.7 and $f_{*}=0$ in Lemma 3.9, we have $\|u\|_{C^{0}}\leq C$, which implies that $|\nabla^{\phi}F^{\prime}|_{\phi}^{2}+\mathrm{tr}\omega_{\phi}\leq C$.
Note that
$|\nabla^{\phi}F|_{\phi}^{2}\leq|\nabla^{\phi}F^{\prime}|_{\phi}^{2}+|\nabla^{\phi}\log\mathrm{v}(m_{\phi})|_{\phi}^{2}\leq C.$
We finish the proof. ∎
Once we have the $C^{2}$-estimate, then, by standard bootstraping argument, we obtain the higher order estimates.
Corollary 5.3.
Suppose $C^{-1}\omega\leq\omega_{\phi}\leq C\omega$ for some constant $C>0$, then all higher derivatives can be estimated in terms of $C$.
Proof.
By the assumption, then (2.13b) is uniformly elliptic with bounded right hand side. It follows that $\|F\|_{C^{\alpha}}\leq C_{1}$ where $\alpha$ and $C_{1}$ depend on $C$. Then going back to (2.13a), we conclude that $\|\phi\|_{C^{2,\alpha^{\prime}}}\leq C_{2}$ for $\alpha^{\prime}<\alpha$, thanks to Evans-Krylov estimate. Hence the coefficients of the elliptic operator of (2.13b) are in $C^{\alpha^{\prime}}$. By [GT77, Theorem 8.32], we may obtain $\|F\|_{C^{1,\alpha^{\prime}}}\leq C_{3}$.
By differentiating both sides of (2.13a), we obtain the linear elliptic equation with $C^{\alpha^{\prime}}$ coefficient
$\Delta_{\phi}(\partial_{k}\phi)-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}\omega_{\phi}(\xi^{\alpha},\partial_{k})=\partial_{k}F-g_{\phi}^{i\overline{j}}(\partial_{k}g_{i\overline{j}})+g^{i\overline{j}}(\partial_{k}g_{i\overline{j}}).$ (5.8)
By Schauder estimate, we deduce $\|\phi\|_{C^{3,\alpha^{\prime}}}\leq C_{4}$. Going back to (2.13b) again, the coefficient are in $C^{\gamma}$ for any $10<\gamma<$ 1, hence we have $\|F\|_{C^{1,\gamma}}\leq C_{5}$. Now, the coefficients of (5.8) are in $C^{\gamma}$ for any $10<\gamma<$ 1. Thus we derive $\|\phi\|_{C^{3,\gamma}}\leq C_{6}$ for any $10<\gamma<$ 1.
Now the second equation (2.13b) has $C^{1,\gamma}$ coefficients and bounded right hand side, the classical $L^{p}$ estimate ([GT77, Theorem 9.11]) implies $F\in W^{2,p}$ for any finite $p$. Differentiating the equation (5.8), we obtain a linear elliptic equation in terms of the second derivatives of $\phi$, which has $C^{2,\gamma}$ coefficients and $L^{p}$ right hand side. Thus $\phi\in W^{4,p}$. The bootstrapping argument concludes the higher order estimates of $\phi$ and $F$. ∎
5.B. Linearization of opeartors
In this subsection, we present the details of the computation for the linearization of operators required in Subsection 4.A. The computation basically follows [Has19, Lah19].
Let ${\mathcal{R}}=\frac{\rm Scal_{\mathrm{v}}}{\mathrm{w}}-\ell_{\mathrm{ext}}$, ${\mathcal{T}}=\frac{\mathrm{tr}_{\mathrm{v},\phi}\theta}{\mathrm{w}}=\frac{\mathrm{v}\cdot\mathrm{tr}_{\phi}\theta+\langle d\mathrm{v},m_{\theta}\rangle}{\mathrm{w}}$. Locally we can write $\theta={\rm dd^{c}}\psi$. Then $m_{\theta}^{\alpha}=-J\xi^{\alpha}(\psi)$, and $\langle\mathrm{d}\mathrm{v},m_{\theta}\rangle=-\sum_{\alpha}\mathrm{v}_{,\alpha}J\xi^{\alpha}(\psi)$. Let ${\mathcal{F}}_{0}={\mathcal{R}}-r{\mathcal{T}}$ for some $r>0$ large enough.
Then
$\frac{d}{dt}|_{t=0}{\mathcal{F}}_{0}(\dot{\phi})$ $=\frac{-{\mathcal{D}}^{*}(\mathrm{v}{\mathcal{D}}\dot{\phi})}{\mathrm{w}}+\langle\partial{\mathcal{R}},\overline{\partial}\dot{\phi}\rangle_{\phi}-r\frac{d}{dt}{\mathcal{T}}$
$=\frac{-{\mathcal{D}}^{*}(\mathrm{v}{\mathcal{D}}\dot{\phi})}{\mathrm{w}}+\langle\partial{\mathcal{F}},\overline{\partial}\dot{\phi}\rangle_{\phi}-r\Big{(}\frac{d}{dt}{\mathcal{T}}-\langle\partial{\mathcal{T}},\overline{\partial}\dot{\phi}\rangle_{\phi}\Big{)}.$
By direct computation,
$-\frac{d}{dt}{\mathcal{T}}=\frac{{\mathcal{T}}}{\mathrm{w}}\langle\partial\mathrm{w},\overline{\partial}\dot{\phi}\rangle_{\phi}-\frac{1}{\mathrm{w}}\frac{d}{dt}\Big{(}\mathrm{v}\cdot\mathrm{tr}_{\phi}\theta-\sum_{\alpha}\mathrm{v}_{,\alpha}J\xi^{\alpha}(\psi)\Big{)},$
$-\frac{1}{\mathrm{w}}\frac{d}{dt}\Big{(}\mathrm{v}\cdot\mathrm{tr}_{\phi}\theta-\sum_{\alpha}\mathrm{v}_{,\alpha}J\xi^{\alpha}(\psi)\Big{)}$
$=$ $\frac{1}{\mathrm{w}}\Big{(}-\langle\partial\mathrm{v},\overline{\partial}\dot{\phi}\rangle_{\phi}\cdot\mathrm{tr}_{\phi}\theta+\mathrm{v}\langle\theta,\sqrt{-1}\partial\overline{\partial}\dot{\phi}\rangle_{\phi}+\sum_{\alpha}\langle\partial\mathrm{v}_{,\alpha},\overline{\partial}\dot{\phi}\rangle_{\phi}J\xi^{\alpha}(\psi)\Big{)}.$
Then
$-\frac{d}{dt}{\mathcal{T}}=\frac{{\mathcal{T}}}{\mathrm{w}}\langle\partial\mathrm{w},\overline{\partial}\dot{\phi}\rangle_{\phi}+\frac{1}{\mathrm{w}}\Big{(}-\langle\partial\mathrm{v},\overline{\partial}\dot{\phi}\rangle_{\phi}\cdot\mathrm{tr}_{\phi}\theta+\mathrm{v}\langle\theta,\sqrt{-1}\partial\overline{\partial}\dot{\phi}\rangle_{\phi}+\sum_{\alpha}\langle\partial\mathrm{v}_{,\alpha},\overline{\partial}\dot{\phi}\rangle_{\phi}J\xi^{\alpha}(\psi)\Big{)}.$
And
$\langle\partial{\mathcal{T}},\overline{\partial}\dot{\phi}\rangle_{\phi}=\frac{-{\mathcal{T}}}{\mathrm{w}}\langle\partial\mathrm{w},\overline{\partial}\dot{\phi}\rangle_{\phi}+\frac{\mathrm{tr}_{\phi}\theta}{\mathrm{w}}\langle\partial\mathrm{v},\overline{\partial}\dot{\phi}\rangle_{\phi}+\frac{\mathrm{v}}{\mathrm{w}}\langle\partial(\mathrm{tr}_{\phi}\theta),\overline{\partial}\dot{\phi}\rangle_{\phi}-\frac{1}{\mathrm{w}}\langle\partial\big{(}\sum_{\alpha}\mathrm{v}_{,\alpha}J\xi^{\alpha}(\psi)\big{)},\overline{\partial}\dot{\phi}\rangle_{\phi}.$
Thus
$\langle\partial{\mathcal{T}},\overline{\partial}\dot{\phi}\rangle_{\phi}-\frac{d}{dt}{\mathcal{T}}=\frac{\mathrm{v}}{\mathrm{w}}\langle\theta,\sqrt{-1}\partial\overline{\partial}\dot{\phi}\rangle_{\phi}+\frac{\mathrm{v}}{\mathrm{w}}\langle\partial(\mathrm{tr}_{\phi}\theta),\overline{\partial}\dot{\phi}\rangle_{\phi}-\frac{1}{\mathrm{w}}\sum_{\alpha}\mathrm{v}_{,\alpha}\langle\partial\big{(}J\xi^{\alpha}(\psi)\big{)},\overline{\partial}\dot{\phi}\rangle_{\phi}.$
Let $f$ be a test function. Then
$\int_{X}\big{(}\langle\partial{\mathcal{T}},\overline{\partial}\dot{\phi}\rangle_{\phi}-\frac{d}{dt}{\mathcal{T}}\big{)}\cdot f\cdot\mathrm{w}\frac{\omega_{\phi}^{n}}{n!}$
$=$ $\int_{X}\mathrm{v}\langle\theta,\sqrt{-1}\partial\overline{\partial}\dot{\phi}\rangle_{\phi}\cdot f\cdot\frac{\omega_{\phi}^{n}}{n!}+\int_{X}\mathrm{v}\cdot\langle\partial(\mathrm{tr}_{\phi}\theta),\overline{\partial}\dot{\phi}\rangle_{\phi}\cdot f\cdot\frac{\omega_{\phi}^{n}}{n!}$
$-\int_{X}\sum_{\alpha}\mathrm{v}_{,\alpha}\langle\partial\big{(}J\xi^{\alpha}(\psi)\big{)},\overline{\partial}\dot{\phi}\rangle_{\phi}\cdot f\cdot\frac{\omega_{\phi}^{n}}{n!},$
where $-\sum_{\alpha}\mathrm{v}_{,\alpha}\langle\partial\big{(}J\xi^{\alpha}(\psi)\big{)},\overline{\partial}\dot{\phi}\rangle_{\phi}=\langle\theta,\partial\mathrm{v}\wedge\overline{\partial}\dot{\phi}\rangle$.
For two $(1,1)$-forms $\alpha,\beta$, we have the formula
$n(n-1)\alpha\wedge\beta\wedge\omega_{\phi}^{n-2}=\big{(}\mathrm{tr}_{\phi}\alpha\cdot\mathrm{tr}_{\phi}\beta-\langle\alpha,\beta\rangle_{\phi}\big{)}\omega_{\phi}^{n}.$
Apply this formula, we have
$\int_{X}\mathrm{v}\langle\theta,\sqrt{-1}\partial\overline{\partial}\dot{\phi}\rangle_{\phi}\cdot f\cdot\frac{\omega_{\phi}^{n}}{n!}$
$=\int\mathrm{v}\cdot\mathrm{tr}_{\phi}\theta\cdot\Delta_{\phi}(\dot{\phi})\cdot f\frac{\omega_{\phi}^{n}}{n!}-\int_{X}\mathrm{v}\cdot f\cdot\theta\wedge\sqrt{-1}\partial\overline{\partial}\dot{\phi}\wedge\frac{\omega_{\phi}^{n-2}}{(n-2)!},$
Among which,
$\int\mathrm{v}\cdot\mathrm{tr}_{\phi}\theta\cdot\Delta_{\phi}(\dot{\phi})\cdot f\frac{\omega_{\phi}^{n}}{n!}=\int_{X}\mathrm{v}\cdot\mathrm{tr}_{\phi}\theta\cdot f\cdot\sqrt{-1}\partial\overline{\partial}\dot{\phi}\wedge\frac{\omega_{\phi}^{n-1}}{(n-1)!}$
$=$ $-\int_{X}\mathrm{v}\cdot\mathrm{tr}_{\phi}\theta\cdot\partial f\wedge\overline{\partial}\dot{\phi}\wedge\frac{\omega_{\phi}^{n-1}}{(n-1)!}-\int_{X}\mathrm{tr}_{\phi}\theta\cdot f\cdot\partial\mathrm{v}\wedge\overline{\partial}\dot{\phi}\wedge\frac{\omega_{\phi}^{n-1}}{(n-1)!}$
$-\int_{X}\mathrm{v}\cdot f\cdot\partial(\mathrm{tr}_{\phi}\theta)\wedge\overline{\partial}\dot{\phi}\wedge\frac{\omega_{\phi}^{n-1}}{(n-1)!},$
and we also have
$-\int_{X}\mathrm{v}\cdot f\cdot(\theta\wedge\sqrt{-1}\partial\overline{\partial}\dot{\phi})\wedge\frac{\omega_{\phi}^{n-2}}{(n-2)!}$
$=$ $\int_{X}\mathrm{v}\cdot\partial f\wedge\overline{\partial}\dot{\phi}\wedge\theta\wedge\frac{\omega_{\phi}^{n-2}}{(n-2)!}+\int_{X}f\cdot(\partial\mathrm{v}\wedge\overline{\partial}\dot{\phi})\wedge\theta\wedge\frac{\omega_{\phi}^{n-2}}{(n-2)!}$
$=$ $\int_{X}\mathrm{v}\cdot\partial f\wedge\overline{\partial}\dot{\phi}\wedge\theta\wedge\frac{\omega_{\phi}^{n-2}}{(n-2)!}+\int_{X}f\cdot\mathrm{tr}_{\phi}(\partial\mathrm{v}\wedge\overline{\partial}\dot{\phi})\cdot\mathrm{tr}_{\phi}\theta\frac{\omega_{\phi}^{n}}{n!}-\int_{X}f\langle\partial\mathrm{v}\wedge\overline{\partial}\dot{\phi},\theta\rangle_{\phi}\frac{\omega_{\phi}^{n}}{n!}.$
Therefore
$\int_{X}\Big{(}\langle\partial{\mathcal{T}},\overline{\partial}\dot{\phi}\rangle_{\phi}-\frac{d}{dt}{\mathcal{T}}\Big{)}\cdot f\cdot\mathrm{w}\cdot\frac{\omega_{\phi}^{n}}{n!}$
$=$ $\int_{X}-\mathrm{v}\cdot\mathrm{tr}_{\phi}\theta\cdot\mathrm{tr}_{\phi}(\partial f\wedge\overline{\partial}\dot{\phi})\frac{\omega_{\phi}^{n}}{n!}+\int_{X}\mathrm{v}\cdot\partial f\wedge\overline{\partial}\dot{\phi}\wedge\theta\wedge\frac{\omega_{\phi}^{n-2}}{(n-2)!}$
$=$ $-\int_{X}\mathrm{v}\cdot\langle\theta,\partial f\wedge\overline{\partial}\dot{\phi}\rangle\frac{\omega_{\phi}^{n}}{n!}.$
By the computation above, we can see that
Lemma 5.4.
The linearized operator $\frac{d}{dt}{\mathcal{T}}-\langle\partial{\mathcal{T}},\overline{\partial}\cdot\rangle_{\phi}$ is a $(\cdot,\cdot)_{\mathrm{w},\phi}$- self-adjoint second order elliptic linear operator.

### v3

@@
 In this paper, we prove that on a smooth Kähler manifold, the $\mathbb{G}$}-coercivity of the weighted Mabuchi functional implies the existence of the $(\mathrm{v},\mathrm{w})$-weighted-cscK (extremal) metric with $\mathrm{v}$ log-concave (firstly studied in [Lah19]), e.g, extremal metrics, Kähler–Ricci solitons, $\mu$-cscK metrics.
-[…]
-[…]
 1. Introduction
@@
 The (generalized) Yau–Tian–Donaldson conjecture predicts an equivalence between the existence of cscK metric(s) and the $\mathbb{G}$}-uniform K-stability condition.
-By the work of many authors (see an incomplete list: [Tia97, DR17, BDL20, CC21a, CC21b, Li22, BJ23]), much progress has been made, and it is known that the existence of cscK metric(s) is equivalent to the ${\mathbb{G}}$-coercivity of the Mabuchi functional, and can be implied by the ${\mathbb{G}}$-uniformly K-stability for models.
+By the work of many authors (see an incomplete list: [Tian97, DR17, BDL20, CC21a, CC21b, Li22a, BJ23]), much progress has been made, and it is known that the existence of cscK metric(s) is equivalent to the $\mathbb{G}$}-coercivity of the Mabuchi functional, and can be implied by the $\mathbb{G}$}-uniformly K-stability for models.
 The category of weighted-cscK metrics includes a large class of important metrics, such as Kähler–Einstein metrics, cscK metrics, Kähler–Ricci solitons, Mabuchi solitons, $\mu$-cscK metrics, etc.
@@
 where for a $(1,1)$-form $\eta$, $\mathrm{tr}_{\mathrm{v},\phi}(\eta)$ is defined by
-$\mathrm{tr}_{\mathrm{v},\phi}(\eta):=\mathrm{v}(m_{\phi})\mathrm{tr}_{\phi}\eta+\langle(\mathrm{d}\mathrm{v})(m_{\phi}),m_{\eta}\rangle,$
-where $\langle(\mathrm{d}\mathrm{v})(m_{\phi}),m_{\eta}\rangle$ is uniformly bounded independent of $\phi$. For the detail definition of $m_{\eta}$ see Subsection 2.A.
+$\mathrm{tr}_{\mathrm{v},\phi}(\eta):=\mathrm{v}(m_{\phi})\mathrm{tr}_{\phi}\eta+\langle\mathrm{d}\mathrm{v}(m_{\phi}),m_{\eta}\rangle,$
+where the definition of $\langle\mathrm{d}\mathrm{v}(m_{\phi}),m_{\eta}\rangle$ can be found in Definition 2.2.
 From the variational point of view, the equation (1.1) can also be considered as the Euler-Lagrangian of the weighted Mabuchi functional ${\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}}}$ (see Definition 2.5).
@@
 for some constants $\delta>0$, $C>0$, and any $\phi\in{\mathcal{E}}^{1}_{K}(X,\omega)$.
-To show (2) implies (1), we consider the following weighted version of the continuity path of Chen [Che18] for weighted extremal metrics,
+To show (2) implies (1), we consider the following weighted version of the continuity path of Chen [Ch18] for weighted extremal metrics,
 $t\left(\frac{S_{\mathrm{v}}(\phi)}{\mathrm{w}(m_{\phi})}-\ell_{\rm ext}(m_{\phi})\right)-(1-t)\left(\frac{\mathrm{tr}_{\mathrm{v},\phi}(\theta)}{\mathrm{w}(m_{\phi})}-\underline{\theta}\right)=0,$ (1.3)
-where $\theta$ is a $K$-invariant smooth positive $(1,1)$-form in the same class of $\omega$.
+where $\theta$ is a $K$-invariant smooth positive $(1,1)$-form in the same class of $\omega$ and $\underline{\theta}$ is defined by (4.2), which is a constant only depending on $\omega]$ by [Lah19, Lemma 2].
 Remark 1.3.
 The idea of the continuity path (1.3) can be seen from the variational picture.
-Since $-\mathrm{Ric}(\omega)$ may not be positive, replacing a small part of the term $-{\mathbb{E}}_{\mathrm{v}}^{\mathrm{Ric}}$ in the weighted Mabuchi functional by ${\mathbb{E}}_{\mathrm{v}}^{\theta}$ for some positive (1,1)-form $\theta$ will make the weighted Mabuchi functional more coercive.
+Since $\mathrm{Ric}(\omega)$ may not be positive, replacing a small part of the term ${\mathbb{E}}_{\mathrm{v}}^{\mathrm{Ric}}$ in the weighted Mabuchi functional by ${\mathbb{E}}_{\mathrm{v}}^{\theta}$ for some positive (1,1)-form $\theta$ will make the weighted Mabuchi functional more coercive.
 Intuitively, this makes (1.3) always solvable for $t\in[0,1)$.
@@
 $\left(\Delta_{\phi}-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}\right)F=-\ell_{\mathrm{ext}}(m_{\phi})\frac{\mathrm{w}(m_{\phi})}{\mathrm{v}(m_{\phi})}+(1-\frac{1}{t})\frac{\mathrm{tr}_{\mathrm{v},\phi}(\theta)}{\mathrm{v}(m_{\phi})}$
-$\qquad\qquad\qquad\qquad\qquad\qquad-(1-\frac{1}{t})\underline{\theta}\frac{\mathrm{w}(m_{\phi})}{\mathrm{v}(m_{\phi})}+\frac{\mathrm{tr}_{\mathrm{v},\phi}\mathrm{Ric}(\omega)}{\mathrm{v}(m_{\phi})}.$ (1.4b)
+$\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt-(1-\frac{1}{t})\underline{\theta}\frac{\mathrm{w}(m_{\phi})}{\mathrm{v}(m_{\phi})}+\frac{\mathrm{tr}_{\mathrm{v},\phi}\mathrm{Ric}(\omega)}{\mathrm{v}(m_{\phi})}.$ (1.4b)
 Compared with the equations in [CC21a], the first equation (1.4a) is a weighted Monge-Ampère equation, and the second linear equation (1.4b) also contains first-order derivative term of $F$ (which is a third-order derivative of $\phi$).
 This makes the $C^{0}$-estimate and $C^{2}$-estimate of $\phi$, the gradient estimate of $F$ more challenging.
-We prove a uniform $C^{0}$-estimate by modifying Guo-Phong’s argument in [GP22].
+We prove a uniform $C^{0}$-estimate by modifying Guo-Phong’s argument in [guo2022].
 In addition, in the calculation of the $C^{2}$-estimate of $\phi$ and the gradient estimate of $F$, we need to modify Chen-Cheng’s method in [CC21b] (in particular, we need to choose new test functions and integral weights in order to deal with the third-order derivative terms (of $\phi$)).
@@
 In Section 4, we prove Theorem 1.2 by using the continuity method.
-We give a sketched proof of the openness part in Subsection 4.A, where we follow the argument in [Che18], [Has19].
+We give a sketched proof of the openness part in Subsection 4.A, where we follow the argument in [Ch18], [Has19].
 We prove the closedness part in Subsection 4.B, where we follow the strategy of [CC21b].
@@
 Gang Tian, Akito Futaki, Chi Li, Eiji Inoue, Kewei Zhang for their interest and helpful comments.
+The authors would like to thank the anonymous referees for many helpful comments.
 J.
-Han is supported by National Key R&D Program of China 2023YFA1009900, NSFC-20701003A012401.
+Han is supported by National Key R&D Program of China 2023YFA1009900, NSFC No.12301059, XHD23A0101.
 2. Preliminaries
@@
 2.A. Weighted Monge-Ampère measures, space of potentials, and functionals
-The variational approach has successful applications in the resolution of Kähler-Einstein problem, see [Tia97, BBGZ13, BBE+19, BBJ21].
+The variational approach has successful applications in the resolution of Kähler-Einstein problem, see [Tian97, BBGZ13, BBEGZ19, BBJ21].
 In this subsection, we will define the notions needed in the variational approach.
-We will follow the lines in [BN14, HL23, AJL23].
+We will follow the lines in [BWN14, HL20, AJL23].
 We will consider $K$-invariant $\omega$-psh functions:
-[…]
-In [BN14], the authors define the weighted Monge-Ampère measure by using Kiselman’s partial Legendre transform and approximation by series of step functions.
-Then in [HL23], the authors present an alternative definition by using a more direct fibration construction, which is more adapted in defining the non-Archimedean functionals.
-Proposition 2.1 ([BN14, HL23]).
+$\mathrm{Psh}_{K}(X,\omega)=\{\phi\in\mathrm{Psh}(X,\omega):\phi\text{ is invariant under the }K\text{-action}\}.$ (2.1)
+In [BWN14], the authors define the weighted Monge-Ampère measure by using Kiselman’s partial Legendre transform and approximation by series of step functions.
+Then in [HL20], the authors present an alternative definition by using a more direct fibration construction, which is more adapted in defining the non-Archimedean functionals.
+Proposition 2.1 ([BWN14, HL20]).
 Let $\mathrm{v}$ be a smooth positive function on the moment polytope $P$.
 For any $\omega$-psh function $\phi\in\mathrm{Psh}_{K}(X,\omega)$, the weighted Monge-Ampère measure ${\rm MA}_{\mathrm{v}}(\phi):=\mathrm{v}(m_{\phi})\omega_{\phi}^{n}$ is well-defined as a Radon measure.
-The following weighted functionals and potential spaces are defined in [BN14, HL23, Lah19, AJL23].
+It should be noted that such weighted Monge–Ampère measure is understood in the sense of non-pluripolar product.
+The following weighted functionals and potential spaces are defined in [BWN14, HL20, Lah19, AJL23].
 Definition 2.2.
@@
 ${\mathbb{E}}_{\mathrm{v}}(\phi,\phi_{0})=\frac{1}{\mathbb{V}_{\mathrm{v}}(X)}\int_{0}^{1}\int_{X}(\phi-\phi_{0})\mathrm{v}(m_{\phi_{t}})\frac{\omega_{\phi_{t}}^{n}}{n!}\mathrm{d}t,$ (2.2)
-${\mathbb{E}}_{\mathrm{v}}^{\eta}(\phi,\phi_{0})=\frac{1}{\mathbb{V}_{\mathrm{v}}(X)}\int_{0}^{1}\int_{X}(\phi-\phi_{0})\Big{(}\mathrm{v}(m_{\phi_{t}})\eta\wedge\frac{\omega_{\phi_{t}}^{n-1}}{(n-1)!}+\langle\mathrm{d}\mathrm{v}(m_{\phi_{t}}),m_{\eta}\rangle_{\phi_{t}}\frac{\omega_{\phi_{t}}^{n}}{n!}\Big{)}\mathrm{d}t,$ (2.3)
+${\mathbb{E}}_{\mathrm{v}}^{\eta}(\phi,\phi_{0})=\frac{1}{\mathbb{V}_{\mathrm{v}}(X)}\int_{0}^{1}\int_{X}(\phi-\phi_{0})\Big(\mathrm{v}(m_{\phi_{t}})\eta\wedge\frac{\omega_{\phi_{t}}^{n-1}}{(n-1)!}+\langle\mathrm{d}\mathrm{v}(m_{\phi_{t}}),m_{\eta}\rangle\frac{\omega_{\phi_{t}}^{n}}{n!}\Big)\mathrm{d}t,$ (2.3)
 where $\eta$ is a smooth $(1,1)$-form, which can be locally expressed as $\eta={\rm dd^{c}}\psi$.
 And $m_{\eta}(\xi):=-J\xi(\psi)$, which is globally well-defined.
-And $\langle d\mathrm{v}(m_{\phi},m_{\eta}\rangle_{\phi}=-\sum_{\alpha}\mathrm{v}_{,\alpha}J\xi^{\alpha}(\psi)$.
+And $\langle\mathrm{d}\mathrm{v}(m_{\phi}),\cdot\rangle=\mathrm{d}\mathrm{v}(m_{\phi})^{\sharp}=\sum_{\alpha}\mathrm{v}_{,\alpha}\xi^{\alpha}$ as a $\mathfrak{t}$-valued function on $X$, where $(\cdot)^{\sharp}$ is with respect to the Kähler form.
+Therefore, $\langle\mathrm{d}\mathrm{v}(m_{\phi}),m_{\eta}\rangle=\sum_{\alpha}\mathrm{v}_{,\alpha}\cdot\xi^{\alpha}(m_{\eta})=-\sum_{\alpha}\mathrm{v}_{,\alpha}J\xi^{\alpha}(\psi)$.
 We denote ${\mathbb{E}}_{\mathrm{v}}(\phi)={\mathbb{E}}_{\mathrm{v}}(\phi,0)$. (The same convention applies for other energy functionals.) We define the weighted I-functional:
@@
 and
-$\frac{1}{C}\big{(}{\mathbb{I}}(\phi)-{\mathbb{J}}(\phi)\big{)}\leq{\mathbb{I}}_{\mathrm{v}}(\phi)-{\mathbb{J}}_{\mathrm{v}}(\phi)\leq C\big{(}{\mathbb{I}}(\phi)-{\mathbb{J}}(\phi)\big{)}.$
+$\frac{1}{C}\big({\mathbb{I}}(\phi)-{\mathbb{J}}(\phi)\big)\leq{\mathbb{I}}_{\mathrm{v}}(\phi)-{\mathbb{J}}_{\mathrm{v}}(\phi)\leq C\big({\mathbb{I}}(\phi)-{\mathbb{J}}(\phi)\big),$
+where $\mathbb{I}$}, $\mathbb{J}$} are the classical $I$-functional and $J$-functional:
+${\mathbb{I}}(\phi):=\frac{1}{\mathbb{V}(X)}\int_{X}\phi\omega^{n}-\frac{1}{\mathbb{V}(X)}\int_{X}\phi\omega_{\phi}^{n}$
+and
+${\mathbb{J}}(\phi):=\frac{1}{\mathbb{V}(X)}\int_{X}\phi\omega^{n}-E(\phi)=\frac{1}{\mathbb{V}(X)}\int_{X}\phi\omega^{n}-\frac{1}{(n+1)\mathbb{V}(X)}\sum_{j=0}^{n}\int_{X}\phi\omega_{\phi}^{j}\wedge\omega^{n-j}.$
 Definition 2.4.
@@
 ${\mathbb{H}}_{\mathrm{v}}(\phi)=\int_{X}\log(\frac{\mathrm{v}(m_{\phi})\omega_{\phi}^{n}}{\omega^{n}})\mathrm{v}(m_{\phi})\frac{\omega_{\phi}^{n}}{n!},\;C_{0}=\int_{X}\log(\mathrm{v}(m_{\omega}))\mathrm{v}(m_{\omega})\frac{\omega^{n}}{n!}.$
-In [Che00, Dar15, BN14, BB17, BDL17, Lah19, HL23, Lah23], the authors show that the space ${\mathcal{E}}^{1}_{K}(X,\omega)$ is granted with a Finsler structure.
+In [Chen00, Dar15, BWN14, BB17, BDL17, Lah19, HL20, Lah23], the authors show that the space ${\mathcal{E}}^{1}_{K}(X,\omega)$ is granted with a Finsler structure.
 The functionals have controlled behaviours with respect to such Finsler structure.
@@
 Remark 2.9.
-All the definitions and propositions of this subsection can be generalized to the setting of projective klt varieties. See for example [HL23, Section 2].
+All the definitions and propositions of this subsection can be generalized to the setting of projective klt varieties. See for example [HL20, Section 2].
 2.B. Weighted-cscK equations
-In this subsection, following the idea in [CC21a], we re-write weighted-cscK equation (1.1) to a coupled PDE system, in which the first one is a $g$-Monge-Ampère equation (as studied in [BN14], [HL23]) and the second one is a linear elliptic equation.
+In this subsection, following the idea in [CC21a], we re-write weighted-cscK equation (1.1) to a coupled PDE system, in which the first one is a $g$-Monge-Ampère equation (as studied in [BWN14], [HL20]) and the second one is a linear elliptic equation.
 Compared with the cscK equation studied in [CC21a], one main difference is that, there are first order derivative terms appearing in the second equation.
@@
 Let
-$F=\log\left(\frac{\mathrm{v}(m_{\phi})\omega_{\phi}^{n}}{\omega^{n}}\right)\quad i.e.,\mathrm{v}(m_{\phi})\omega_{\phi}^{n}=e^{F}\omega^{n}.$
+$F=\log\left(\frac{\mathrm{v}(m_{\phi})\omega_{\phi}^{n}}{\omega^{n}}\right)\qquad i.e.,\mathrm{v}(m_{\phi})\omega_{\phi}^{n}=e^{F}\omega^{n}.$
 Then the linearization of L.H.S. is given by
@@
 where $\phi_{t}=\phi+t\psi$ and $\Delta_{\phi}:=\mathrm{tr}_{\phi}\sqrt{-1}\partial\overline{\partial}$, we have used [Lah19, Lemma 1]
-$m_{\phi}^{\xi}=m_{\omega}^{\xi}+t\mathrm{d}^{\rm c}\phi(\xi).$
+$m_{\phi}^{\xi}=m_{\omega}^{\xi}+t\mathrm{d}^{\rm c}\phi(\xi).$ (2.9)
 Note that
@@
 One has
-$\Delta_{\phi}\log\mathrm{v}(m_{\phi})=\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}\Delta_{\phi}(m^{\xi^{\alpha}}_{\phi})+\frac{\mathrm{v}_{,\alpha\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}-\frac{\mathrm{v}_{,\alpha}(m_{\phi})\mathrm{v}_{,\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})^{2}}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}$ (2.10)
+$\Delta_{\phi}\log\mathrm{v}(m_{\phi})=$ $g_{\phi}^{i\overline{j}}\partial_{i}\partial_{\overline{j}}(\log\mathrm{v}(m_{\phi}))=g_{\phi}^{i\overline{j}}\partial_{i}\frac{\partial_{\overline{j}}\mathrm{v}(m_{\phi})}{\mathrm{v}(\phi)}=g_{\phi}^{i\overline{j}}\partial_{i}\frac{\mathrm{v}_{,\alpha}(m_{\phi})\partial_{\overline{j}}m_{\phi}^{\xi^{\alpha}}}{\mathrm{v}(\phi)}$
+$=$ $g_{\phi}^{i\overline{j}}\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}\partial_{i}\partial_{\overline{j}}m_{\phi}^{\xi^{\alpha}}+g_{\phi}^{i\overline{j}}\frac{\mathrm{v}_{,\alpha\beta}(m_{\phi})\partial_{i}m_{\phi}^{\xi^{\beta}}\partial_{\overline{j}}m_{\phi}^{\xi^{\alpha}}}{\mathrm{v}(\phi)}-g_{\phi}^{i\overline{j}}\frac{\mathrm{v}_{,\beta}(m_{\phi})\partial_{i}m_{\phi}^{\xi^{\alpha}}\mathrm{v}_{,\alpha}(m_{\phi})\partial_{\overline{j}}m_{\phi}^{\xi^{\alpha}}}{\mathrm{v}(\phi)^{2}}$
+$=$ $\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}\Delta_{\phi}(m^{\xi^{\alpha}}_{\phi})+\frac{\mathrm{v}_{,\alpha\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}-\frac{\mathrm{v}_{,\alpha}(m_{\phi})\mathrm{v}_{,\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})^{2}}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}$ (2.11)
 and
@@
 $=$ $-\frac{S_{\mathrm{v}}(\phi)}{\mathrm{v}(m_{\phi})}-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}\Delta_{\omega}(m^{\xi^{\alpha}}_{\omega})+\mathrm{tr}_{\phi}\mathrm{Ric}(\omega)$
-$=$ $-\frac{S_{\mathrm{v}}(\phi)}{\mathrm{v}(m_{\phi})}+\mathrm{v}(m_{\phi})^{-1}\Big{(}\langle(\mathrm{d}\mathrm{v})(m_{\phi}),m_{\mathrm{Ric}(\omega)}\rangle+\mathrm{v}(m_{\phi})\mathrm{tr}_{\phi}\mathrm{Ric}(\omega)\Big{)}$
-$=$ $-\frac{S_{\mathrm{v}}(\phi)}{\mathrm{v}(m_{\phi})}+\frac{\mathrm{tr}_{\mathrm{v},\phi}\mathrm{Ric}(\omega)}{\mathrm{v}(m_{\phi})}$
+$=$ $-\frac{S_{\mathrm{v}}(\phi)}{\mathrm{v}(m_{\phi})}+\mathrm{v}(m_{\phi})^{-1}\Big(\langle\mathrm{d}\mathrm{v}(m_{\phi}),m_{\mathrm{Ric}(\omega)}\rangle+\mathrm{v}(m_{\phi})\mathrm{tr}_{\phi}\mathrm{Ric}(\omega)\Big)$
+$=$ $-\frac{S_{\mathrm{v}}(\phi)}{\mathrm{v}(m_{\phi})}+\frac{\mathrm{tr}_{\mathrm{v},\phi}\mathrm{Ric}(\omega)}{\mathrm{v}(m_{\phi})},$
 where
-$S_{\mathrm{v}}(\phi)=\mathrm{v}(m_{\phi})S(\phi)-2\mathrm{v}_{,\alpha}\Delta_{\phi}(m_{\phi}^{\xi^{\alpha}})-\mathrm{v}_{,\alpha\beta}(m_{\phi})\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}.$
-and $m_{\mathrm{Ric}(\omega)}:=\Delta_{\omega}(m_{\omega})$.
+$S_{\mathrm{v}}(\phi)=\mathrm{v}(m_{\phi})S(\phi)-2\mathrm{v}_{,\alpha}\Delta_{\phi}(m_{\phi}^{\xi^{\alpha}})-\mathrm{v}_{,\alpha\beta}(m_{\phi})\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi},$
+$m_{\mathrm{Ric}(\omega)}=\Delta_{\omega}(m_{\omega})$ and
+$-\mathrm{v}_{,\alpha}(m_{\phi})\Delta_{\omega}(m^{\xi^{\alpha}}_{\omega})=\langle\mathrm{d}\mathrm{v}(m_{\phi}),m_{\mathrm{Ric}(\omega)}\rangle$
+by [Lah19, Lemma 5].
+Indeed, locally we can write $\omega={\rm dd^{c}}\psi$, $\mathrm{Ric}(\omega)=-{\rm dd^{c}}\log\det(\psi_{i\overline{j}})$, and Ricci potential is $\log\det(\psi_{i\overline{j}})$.
+Then $\Delta_{\omega}(m^{\xi^{\alpha}}_{\omega})=g^{i\overline{j}}\partial_{i\overline{j}}(-J\xi^{\alpha}\psi)=J\xi^{\alpha}(-\log\det(\psi_{i\overline{j}}))=m_{\mathrm{Ric}(\omega)}$.
 The continuity path (1.3) can be reformulated into the following coupled PDE system
@@
 $\Delta_{\phi}F=$ $-\ell_{\mathrm{ext}}(m_{\phi})\frac{\mathrm{w}(m_{\phi})}{\mathrm{v}(m_{\phi})}+\frac{\mathrm{tr}_{\mathrm{v},\phi}(\eta_{t})}{\mathrm{v}(m_{\phi})}-(1-\frac{1}{t})\underline{\theta}\frac{\mathrm{w}(m_{\phi})}{\mathrm{v}(m_{\phi})}$
-$-\frac{\mathrm{v}_{,\alpha}(m_{\phi})\mathrm{v}_{,\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})^{2}}\langle\xi^{\beta},\xi^{\alpha}\rangle_{\phi}-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}\Delta_{\phi}(m_{\phi}^{\xi^{\alpha}}-m_{\omega}^{\xi^{\alpha}})$
+$-\frac{\mathrm{v}_{,\alpha}(m_{\phi})\mathrm{v}_{,\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})^{2}}\langle\xi^{\beta},\xi^{\alpha}\rangle_{\phi}-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}(\Delta_{\phi}m_{\phi}^{\xi^{\alpha}}-\Delta_{\omega}m_{\omega}^{\xi^{\alpha}})$
 and
@@
 $-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}\Delta_{\phi}(m^{\xi^{\alpha}}_{\phi})-\frac{\mathrm{v}_{,\alpha\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}+\frac{\mathrm{v}_{,\alpha}(m_{\phi})\mathrm{v}_{,\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})^{2}}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}$
-$=$ $-\left((\ell_{\mathrm{ext}}+(1-\frac{1}{t})\underline{\theta})\frac{\mathrm{w}}{\mathrm{v}}-(1-\frac{1}{t})\frac{\langle(\mathrm{d}\mathrm{v})(m_{\phi}),m_{\theta}\rangle}{\mathrm{v}}\right)+\mathrm{tr}_{\phi}\eta_{t}$
+$=$ $-\left((\ell_{\mathrm{ext}}+(1-\frac{1}{t})\underline{\theta})\frac{\mathrm{w}}{\mathrm{v}}-(1-\frac{1}{t})\frac{\langle\mathrm{d}\mathrm{v}(m_{\phi}),m_{\theta}\rangle}{\mathrm{v}}\right)+\mathrm{tr}_{\phi}\eta_{t}$
 $-2\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}\Delta_{\phi}m_{\phi}^{\xi^{\alpha}}-\frac{\mathrm{v}_{,\alpha\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}.$
@@
 where
-$f_{t}^{\prime}:=(\ell_{\mathrm{ext}}+(1-\frac{1}{t})\underline{\theta})\frac{\mathrm{w}}{\mathrm{v}}-(1-\frac{1}{t})\frac{\langle(\mathrm{d}\mathrm{v})(m_{\phi}),m_{\theta}\rangle}{\mathrm{v}}+2\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}\Delta_{\omega}(m_{\omega}^{\xi^{\alpha}}).$
-In addition, by (2.10) and (2.14)
+$f_{t}^{\prime}:=(\ell_{\mathrm{ext}}+(1-\frac{1}{t})\underline{\theta})\frac{\mathrm{w}}{\mathrm{v}}-(1-\frac{1}{t})\frac{\langle\mathrm{d}\mathrm{v}(m_{\phi}),m_{\theta}\rangle}{\mathrm{v}}+2\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}\Delta_{\omega}(m_{\omega}^{\xi^{\alpha}}).$
+In addition, by (2.B) and (2.15)
 $\Delta_{\phi}\log\mathrm{v}(m_{\phi})=$ $-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}(F-\log\mathrm{v}(m_{\phi}))+\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}\Delta_{\omega}(m_{\omega}^{\xi^{\alpha}}).$
@@
 where
-$f_{t}=(\ell_{\mathrm{ext}}+(1-\frac{1}{t})\underline{\theta})\frac{\mathrm{w}}{\mathrm{v}}-(1-\frac{1}{t})\frac{\langle(\mathrm{d}\mathrm{v})(m_{\phi}),m_{\theta}\rangle}{\mathrm{v}}.$
+$f_{t}=(\ell_{\mathrm{ext}}+(1-\frac{1}{t})\underline{\theta})\frac{\mathrm{w}}{\mathrm{v}}-(1-\frac{1}{t})\frac{\langle\mathrm{d}\mathrm{v}(m_{\phi}),m_{\theta}\rangle}{\mathrm{v}}.$
 We have
@@
 $\left(\Delta_{\phi}-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}\right)F=-\ell_{\mathrm{ext}}(m_{\phi})\frac{\mathrm{w}(m_{\phi})}{\mathrm{v}(m_{\phi})}+(1-\frac{1}{t})\frac{\mathrm{tr}_{\mathrm{v},\phi}(\theta)}{\mathrm{v}(m_{\phi})}$
-$\qquad\qquad\qquad\qquad\qquad\quad-\frac{\mathrm{tr}_{\mathrm{v},\phi}({\rm dd^{c}}f_{*})}{\mathrm{v}(m_{\phi})}-(1-\frac{1}{t})\underline{\theta}\frac{\mathrm{w}(m_{\phi})}{\mathrm{v}(m_{\phi})}+\frac{\mathrm{tr}_{\mathrm{v},\phi}\mathrm{Ric}(\omega)}{\mathrm{v}(m_{\phi})},$ (2.20b)
+$\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\qquad-\frac{\mathrm{tr}_{\mathrm{v},\phi}({\rm dd^{c}}f_{*})}{\mathrm{v}(m_{\phi})}-(1-\frac{1}{t})\underline{\theta}\frac{\mathrm{w}(m_{\phi})}{\mathrm{v}(m_{\phi})}+\frac{\mathrm{tr}_{\mathrm{v},\phi}\mathrm{Ric}(\omega)}{\mathrm{v}(m_{\phi})},$ (2.21b)
 where $f_{*}\in\mathrm{Psh}(X,(\frac{1-t}{t})\theta)$, $\sup_{X}f_{*}=0$.
@@
 $\log(\mathrm{Vol}^{\lambda})(\xi)=\frac{1}{\int_{X}e^{m_{\phi}^{\xi}}\omega_{\phi}^{n}}\int_{X}\Big(S(\omega_{\phi})-\Delta_{\phi}m_{\phi}^{\xi}-\lambda m_{\phi}^{\xi}\Big)e^{m_{\phi}^{\xi}}\omega_{\phi}^{n}+\lambda\cdot\log(\int_{X}e^{m_{\phi}^{\xi}}\omega_{\phi}^{n}).$
-After a normalization, $\log(\mathrm{Vol}^{\lambda})$ has the same expression as Perelman’s $W$-functional [Per02].
+After a normalization, $\log(\mathrm{Vol}^{\lambda})$ has the same expression as Perelman’s $W$-functional [Pe02].
 When restricted on Fano manifolds, $\log(\mathrm{Vol}^{\lambda})$ is equal to Tian–Zhu’s volume functional [TZ02] up to the difference of a constant.
-Some other related works can be found in [He19, Ino21, Liu22, SD22, Jub23, Zha23, Fut24, GJSS24, His24, NN24].
+Some other related works can be found in [He19, Ino21, Liu22, Dyr22, Jub23, Zha23, Fut24, GJSS24, His24, NN24].
 3. Apriori estimates
 In this section, we provide uniform $C^{0}$-estimate, $W^{2,p}$-estimate of $\phi$ and gradient estimate of $F+f_{*}-\log(\mathrm{v}(m_{\phi}))$ for equation (2.21a), (2.21b).
-The proof of the uniform $C^{0}$-estimate is based on a modification of Guo–Phong’s method [GP22].
+The proof of the uniform $C^{0}$-estimate is based on a modification of Guo–Phong’s method [guo2022].
 The proof of $W^{2,p}$-estimate and gradient estimate follows from Chen–Cheng’s method [CC21b] with a new choice of test functions and integral weights.
@@
 Let $\omega_{X}$ be a fixed smooth Kähler form, and let $\omega=\chi+s\omega_{X}$ be a smooth Kähler form, where $\chi$ is a smooth $(1,1)$-form in a nef class, $s>0$.
-We will first consider the uniform $C^{0}$-estimate of the following PDE system
-$\begin{cases}\mathrm{v}(m_{\phi})\;\omega_{\phi}^{n}=c_{\omega,\mathrm{v}}\;e^{\widetilde{F}}\omega_{X}^{n},\\ (\Delta_{\phi}-\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha})\big{(}\widetilde{F}+f_{*}\big{)}=-f+\mathrm{tr}_{\phi}\eta,\end{cases}$ (3.1)
+We simplify the equations (2.21a), (2.21b).
+The equation (2.21a) can be rewritten as
+$\mathrm{v}(m_{\phi})\;\omega_{\phi}^{n}=e^{F+\log\frac{\omega^{n}}{\omega_{X}^{n}}}\omega_{X}^{n}=e^{\widetilde{F}+\log c_{\omega,\mathrm{v}}}\omega_{X}^{n},$
+where we denote $\widetilde{F}=F+\log\frac{\omega^{n}}{\omega_{X}^{n}}-\log c_{\omega,\mathrm{v}}$. On the other hand, the equation (2.21b) can be re-written as
+$(\Delta_{\phi}-\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha})(F+f_{*})=-f_{t}-\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}\Delta_{\omega}(m_{\omega}^{\xi^{\alpha}})+\mathrm{tr}_{\phi}((1-\frac{1}{t})\theta+\mathrm{Ric}(\omega)).$
+Thus, by (2.12), one has
+$(\Delta_{\phi}-\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha})(\widetilde{F}+f_{*})=$ $-f_{t}-\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}\Delta_{\omega}(m_{\omega}^{\xi^{\alpha}})+\mathrm{tr}_{\phi}((1-\frac{1}{t})\theta+\mathrm{Ric}(\omega))$
+$+\Delta_{\phi}\log\frac{\omega^{n}}{\omega_{X}^{n}}-\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha}\log\frac{\omega^{n}}{\omega_{X}^{n}}$
+$=$ $-f_{t}-\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}\Delta_{\omega_{X}}(m_{\omega_{X}}^{\xi^{\alpha}})+\mathrm{tr}_{\phi}\Big((1-\frac{1}{t})\theta+\mathrm{Ric}(\omega_{X})\Big).$
+We will consider the uniform $C^{0}$-estimate of the following PDE system
+$\begin{cases}\mathrm{v}(m_{\phi})\;\omega_{\phi}^{n}=c_{\omega,\mathrm{v}}\;e^{\widetilde{F}}\omega_{X}^{n},\\ (\Delta_{\phi}-\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha})\big(\widetilde{F}+f_{*}\big)=-f+\mathrm{tr}_{\phi}\eta,\end{cases}$ (3.1)
 where $c_{\omega,\mathrm{v}}=\int_{X}\mathrm{v}\;\omega^{n}/{\int_{X}\omega_{X}^{n}}$.
@@
 In the modified continuity path, for the $(\frac{1-t}{t})\theta$-psh function $f_{*}$, for arbitrarily large $p$, we have $\int_{X}e^{-pf_{*}}\omega^{n}$ uniformly bounded from above when $t$ is sufficiently close to $1$ (see Proposition 4.8).
-Furthermore, the function $\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha}(f_{*})$ is determined by the class $[\theta]$, which is bounded.
-Therefore the modified continuity path (2.20a), (2.20b) can be reformulated into (3.1) when $t$ is sufficiently close to $1$.
+Furthermore, the function $\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha}(f_{*})$ is bounded by a constant which only depends on the class $\theta]$.
+Indeed, denote $\theta_{t}:=\frac{1-t}{t}\theta$ and by (2.9),
+$-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}(f_{*})=\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}\mathrm{d}^{c}f_{*}(\xi^{\alpha})=\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}(m_{\theta_{t}+{\rm dd^{c}}f_{*}}^{\xi^{\alpha}}-m_{\theta_{t}}^{\xi^{\alpha}}),$
+where the right hand side only depends on the moment polytope determined by $\theta]$.
+Therefore the modified continuity path (2.21a), (2.21b) can be reformulated into (3.1) when $t$ is sufficiently close to 11.
 Theorem 3.2.
@@
 Proof.
-By applying [GP22, Theorem 2] to equation $\omega_{\phi}^{n}=c_{\omega}e^{\widetilde{F}}\omega_{X}^{n}$, for $l>n$, we have $\sup_{X}|\phi|\leq C$, where $C$ depends on $\omega_{X},n,l,\kappa,{\rm Ent}_{l}(e^{\widetilde{F}})$.
-Therefore, we just need to modify [GP22, Theorem 3] to use $\int_{X}\widetilde{F}e^{\widetilde{F}}\omega_{X}^{n}$ to bound ${\rm Ent}_{l}(\widetilde{F})$.
+[…]
+[…]
 We need to use the auxiliary function $\psi_{k}$ which satisfies the following equation
@@
 $\leq\frac{c_{\omega,\mathrm{v}}}{V_{\omega}}\int_{\Omega}\Big(-\phi+\delta\cdot\widetilde{F}\Big)e^{\widetilde{F}}\omega_{X}^{n},$
-where $\Omega=\{x\in X:\widetilde{F}+f_{*}>0\}$ and $0\leq A_{\infty}\leq C(K_{1})$, which is due to the entropy bound and the weak compactness implied by it.
+where $\Omega=\{x\in X:-\phi+\delta(\widetilde{F}+f_{*})>0\}$ and $0\leq A_{\infty}\leq C(K_{1})$, which is due to the entropy bound assumption, [guo2022, Theorem 1 (ii)] for $p=1$ and Hölder inequality.
 Consider
 $\Psi=-\epsilon(-\psi_{k}+\Lambda)^{\frac{n}{n+1}}-\phi+\delta(\widetilde{F}+f_{*}),$ (3.3)
-where $\epsilon=(\frac{(n+1)(n+C_{f,\delta})}{n^{2}})^{\frac{n}{n+1}}A_{k}^{\frac{1}{n+1}}$, $\Lambda^{\frac{1}{n+1}}=\frac{2n}{n+1}\epsilon$, where the constant $C_{f,\delta}$ is chosen such that $\left|-\delta f+\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha}(\phi)\right|\leq C_{f,\delta}$.
+where $\epsilon=2^{\frac{n-1}{n+1}}(\frac{(n+1)(n+C_{f,\delta})}{n^{2}})^{\frac{n}{n+1}}A_{k}^{\frac{1}{n+1}}$, $\Lambda^{\frac{1}{n+1}}=\frac{2n}{n+1}\epsilon$, where the constant $C_{f,\delta}$ is chosen such that $\left|-\delta f+\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha}(\phi)\right|\leq C_{f,\delta}$.
 The next is to show that $\Psi\leq 0$. It suffices to assume $\sup_{X}\Psi$ is achieved at a point $x_{0}\in\Omega$. We have
 $\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha}(\Psi)(x_{0})=0,$
-which implies
-$\delta\cdot\Delta_{\phi}(\widetilde{F}+f_{*})=-\delta\cdot f+\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha}(\phi)+\delta\cdot\mathrm{tr}_{\phi}\eta-\frac{n\epsilon}{n+1}\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha}(\psi_{k})\cdot(-\psi_{k}+\Lambda)^{-\frac{1}{n+1}}.$
+which together with the second equation of (3.1) implies
+$\delta\cdot\Delta_{\phi}(\widetilde{F}+f_{*})=-\delta\cdot f+\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha}(\phi)+\delta\cdot\mathrm{tr}_{\phi}\eta-\frac{n\epsilon}{n+1}\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha}(\psi_{k})\cdot(-\psi_{k}+\Lambda)^{-\frac{1}{n+1}}.$ (3.4)
 Then at $x_{0}$, we have
-00 0 $\geq\Delta_{\phi}\Psi$
+$0$ $\geq\Delta_{\phi}\Psi$
 $\geq\frac{\epsilon n}{n+1}(-\psi_{k}+\Lambda)^{\frac{-1}{n+1}}({\rm tr}_{\phi}\omega_{\psi_{k}}-{\rm tr}_{\phi}\omega)-(n-{\rm tr}_{\phi}\omega)+\delta\Delta_{\phi}(\widetilde{F}+f_{*})$
 $\geq\frac{\epsilon n^{2}}{n+1}(-\psi_{k}+\Lambda)^{\frac{-1}{n+1}}\Big(\frac{-\phi+\delta(\widetilde{F}+f_{*})}{A_{k}}\Big)^{\frac{1}{n}}-n-C_{f,\delta}$
-$\quad+\left(1-\frac{\epsilon n}{n+1}(-\psi_{k}+\Lambda)^{\frac{-1}{n+1}}-\frac{1}{10}\right){\rm tr}_{\phi}\omega-\frac{n\epsilon}{n+1}\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha}(\psi_{k})(-\psi_{k}+\Lambda)^{\frac{-1}{n+1}}$
+$\qquad+\left(1-\frac{\epsilon n}{n+1}(-\psi_{k}+\Lambda)^{\frac{-1}{n+1}}-\frac{1}{10}\right){\rm tr}_{\phi}\omega-\frac{n\epsilon}{n+1}\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha}(\psi_{k})(-\psi_{k}+\Lambda)^{\frac{-1}{n+1}}$
 $\geq\frac{\epsilon n^{2}}{n+1}(-\psi_{k}+\Lambda)^{\frac{-1}{n+1}}\left(\Big(\frac{-\phi+\delta(\widetilde{F}+f_{*})}{A_{k}}\Big)^{\frac{1}{n}}-\frac{\mathrm{v}_{\alpha}}{n\mathrm{v}}J\xi^{\alpha}(\psi_{k})\right)-(n+C_{f,\delta}).$
 Note that $|\frac{\mathrm{v}_{\alpha}}{n\mathrm{v}}J\xi^{\alpha}(\psi_{k})|$ is bounded by some constant $C^{\prime}$ uniformly. This implies
-$-\phi+\delta(\widetilde{F}+f_{*})\leq\left(\frac{(n+1)(n+C_{f,\delta})}{\epsilon n^{2}}\right)^{n}A_{k}(-\psi_{k}+\Lambda)^{\frac{n}{n+1}}+C^{\prime}=\epsilon(-\psi_{k}+\Lambda)^{\frac{n}{n+1}}+C^{\prime}.$
+$-\phi+\delta(\widetilde{F}+f_{*})\leq$ $\left[C^{\prime}+\frac{(n+1)(n+C_{f,\delta})}{\epsilon n^{2}}(-\psi_{k}+\Lambda)^{\frac{1}{n+1}}\right]^{n}A_{k}$
+$\leq$ $2^{n-1}\left(\frac{(n+1)(n+C_{f,\delta})}{\epsilon n^{2}}\right)^{n}A_{k}(-\psi_{k}+\Lambda)^{\frac{n}{n+1}}+2^{n-1}C^{\prime n}A_{k}$
+$=$ $\epsilon(-\psi_{k}+\Lambda)^{\frac{n}{n+1}}+C^{\prime}.$
 This concludes that $\sup_{X}\Psi=\Psi(x_{0})\leq C^{\prime}$. Therefore
@@
 $\leq C\Big(\int_{X}e^{-qr\widetilde{\epsilon}\psi_{k}}\omega_{X}^{n}\Big)^{\frac{1}{q}}\leq C^{\prime}.$
-In particular, this implies $\mathrm{Ent}_{l}(\widetilde{F})$ is uniformly bounded.
-Then by [GPT23, Theorem 1] we have $\sup_{X}|\phi|\leq C$, where the constant $C$ depends on parameters listed in the statement of the theorem.
-(Alternatively, we can also use the $C^{0}$-estimate in [EGZ09]).
-In order to prove the upper and lower bounds of $\widetilde{F}+f_{*}$, we need to use the mean-value type inequality proved in [GP22].
+[…]
+In order to prove the upper and lower bounds of $\widetilde{F}+f_{*}$, we need to use the mean-value type inequality proved in [guo2022].
 Let $u:=\widetilde{F}+f_{*}-K_{2}\phi$, which satisfies $(\Delta_{\phi}-\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha})u\geq-a$ for some constant $a\geq 0$.
-Apply [GP22, Lemma 3] to $u$, we can obtain the upper bound of $\widetilde{F}+f_{*}$.
+Apply [guo2022, Lemma 3] to $u$, we can obtain the upper bound of $\widetilde{F}+f_{*}$.
 Similarly, let $u:=-(\widetilde{F}+f_{*})-K_{3}\phi$, which satisfies $(\Delta_{\phi}-\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha})u\geq-a$ for some $a\geq 0$, we can obtain a lower bound of $\widetilde{F}+f_{*}$.
@@
 For the convenience of the computation, since
-$\frac{\mathrm{tr}_{\mathrm{v},\phi}({\rm dd^{c}}f_{*})}{\mathrm{v}(m_{\phi})}=\Delta_{\phi}f_{*}+\frac{\langle(\mathrm{d}\mathrm{v})(m_{\phi}),\mathrm{d}^{c}f_{*}\rangle}{\mathrm{v}}=\Delta_{\phi}f_{*}-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}f_{*},$
+$\frac{\mathrm{tr}_{\mathrm{v},\phi}({\rm dd^{c}}f_{*})}{\mathrm{v}(m_{\phi})}=\Delta_{\phi}f_{*}+\frac{\langle\mathrm{d}\mathrm{v}(m_{\phi}),\mathrm{d}^{c}f_{*}\rangle}{\mathrm{v}}=\Delta_{\phi}f_{*}-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}f_{*},$
 we rewrite the coupled system (2.21a), (2.21b) into
@@
 where $f_{*}\in\mathrm{Psh}(X,(\frac{1}{t}-1)\theta)$, $\sup_{X}f_{*}=0$,
-$\widetilde{f}_{t}=-(\ell_{\mathrm{ext}}+(1-\frac{1}{t})\underline{\theta})\frac{\mathrm{w}}{\mathrm{v}}+(1-\frac{1}{t})\frac{\langle(\mathrm{d}\mathrm{v})(m_{\phi}),m_{\theta}\rangle}{\mathrm{v}}-\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}\Delta_{\omega}m_{\omega}^{\xi^{\alpha}}$
+$\widetilde{f}_{t}=-(\ell_{\mathrm{ext}}+(1-\frac{1}{t})\underline{\theta})\frac{\mathrm{w}}{\mathrm{v}}+(1-\frac{1}{t})\frac{\langle\mathrm{d}\mathrm{v}(m_{\phi}),m_{\theta}\rangle}{\mathrm{v}}-\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}\Delta_{\omega}m_{\omega}^{\xi^{\alpha}}$
 which is a bounded function, $\eta_{t}=(1-\frac{1}{t})\theta+\mathrm{Ric}(\omega)$, and $\int_{X}e^{-pf_{*}}\omega^{n}<C$ for a sufficiently large positive constant $p$.
 (which can be achieved when $t\to 1$ in the modified continuity path.)
+Note that by (2.B) and (2.13), the second equation of (3.5) can be rewritten as
+$\Delta_{\phi}(F+\log\mathrm{v}(m_{\phi}))=$ $\widetilde{f_{t}}+\mathrm{tr}_{\phi}(\eta_{t})-\Delta_{\phi}f_{*}+\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha}f_{*}+\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha}F+\Delta_{\phi}\log\mathrm{v}(m_{\phi})$
+$=$ $\widetilde{f_{t}}+\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}\Delta_{\omega}(m_{\omega}^{\xi^{\alpha}})+\mathrm{tr}_{\phi}(\eta_{t})-\Delta_{\phi}f_{*}+\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha}f_{*}$
+$-2|\nabla^{\phi}\log\mathrm{v}(m_{\phi})|_{\phi}^{2}+\frac{\mathrm{v}_{,\alpha\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}$
+$=$ $-f_{t}+\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha}f_{*}-\Delta_{\phi}f_{*}+\mathrm{tr}_{\phi}\eta_{t}-2|\nabla^{\phi}\log\mathrm{v}(m_{\phi})|_{\phi}^{2}+\frac{\mathrm{v}_{,\alpha\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}.$ (3.6)
 Lemma 3.4.
@@
 Proof.
-By [Szfrm[o]–4, Lemma 3.7], then
+By [Sze14, Lemma 3.7], then
 $\Delta_{\phi}\log\mathrm{tr}\omega_{\phi}\geq-B\mathrm{tr}_{\phi}\omega-\frac{g^{j\overline{k}}\mathrm{Ric}(\phi)_{j\overline{k}}}{\mathrm{tr}\omega_{\phi}},$
@@
 $=$ $-au\Delta_{\phi}(F+\log\mathrm{v}(m_{\phi})+\delta f_{*}+b\phi)+u\Delta_{\phi}\log\mathrm{tr}\omega_{\phi}+u^{-1}|\nabla^{\phi}u|_{\phi}^{2}.$ (3.9)
-By (2.17), (3.4) and Remark 3.1, we have
+By (3.6) and Remark 3.1, we have
 $\Delta_{\phi}(F+\log\mathrm{v}(m_{\phi})+\delta f_{*}+b\phi)$
@@
 $\Delta_{\phi}u$ $\geq e^{-a(F+\log\mathrm{v}+\delta f_{*}+b\phi)}\Big(a(f_{t}-bn-C+(1-\delta)\Delta_{\phi}f_{*})\mathrm{tr}\omega_{\phi}+(ab-aA_{1}-B)\mathrm{tr}_{\phi}\omega\mathrm{tr}\omega_{\phi}$
-$\qquad\quad+\Delta F^{\prime}-S(\omega)\Big{)}+2au|\nabla^{\phi}\log\mathrm{v}(m_{\phi})|_{\phi}^{2}-au\frac{\mathrm{v}_{,\alpha\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}+u^{-1}|\nabla^{\phi}u|_{\phi}^{2}\ ,$ (3.9)
+$\hskip 18.49988pt\qquad+\Delta F^{\prime}-S(\omega)\Big)+2au|\nabla^{\phi}\log\mathrm{v}(m_{\phi})|_{\phi}^{2}-au\frac{\mathrm{v}_{,\alpha\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}+u^{-1}|\nabla^{\phi}u|_{\phi}^{2}\ ,$ (3.11)
 where we denote $F^{\prime}:=F-\log\mathrm{v}(m_{\phi})$.
@@
 $\frac{1}{2p+1}\Delta_{\phi}u^{2p+1}=$ $u^{2p}\Delta_{\phi}u+2pu^{2p-1}|\nabla^{\phi}u|_{\phi}^{2}.$
-Take the integral on the above equality with respect to the weighted measure $e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$, together with (3.9), we obtain
-$\int_{X}2pu^{2p-1}\cdot|\nabla_{\phi}u|_{\phi}^{2}\cdot e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
-$=$ $\int_{X}u^{2p}\Big{(}-\Delta_{\phi}u-\langle\nabla_{\phi}u,2a\nabla_{\phi}\log\mathrm{v}\rangle_{\phi}\Big{)}\cdot e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
-$\leq$ $-2a\int_{X}u^{2p}\langle\nabla_{\phi}u,\nabla_{\phi}\log\mathrm{v}\rangle_{\phi}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
+Integrate the equality above with respect to the weighted measure $e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$, together with (3.11), we obtain
+$\int_{X}2pu^{2p-1}\cdot|\nabla^{\phi}u|_{\phi}^{2}\cdot e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
+$=$ $\int_{X}u^{2p}\Big(-\Delta_{\phi}u-\langle\nabla^{\phi}u,2a\nabla^{\phi}\log\mathrm{v}\rangle_{\phi}\Big)\cdot e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
+$\leq$ $-2a\int_{X}u^{2p}\langle\nabla^{\phi}u,\nabla^{\phi}\log\mathrm{v}\rangle_{\phi}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
 $-\int_{X}u^{2p}e^{-a(F+\log\mathrm{v}+\delta f_{*}+b\phi)}\Bigl\{a\Big(f_{t}-bn-C+(1-\delta)\Delta_{\phi}f_{*}\Big)\mathrm{tr}\omega_{\phi}+(ab-aA_{1}-B)\mathrm{tr}_{\phi}\omega\mathrm{tr}\omega_{\phi}$
 $+\Delta(F-\log\mathrm{v})-S(\omega)\Bigr\}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
-$-2a\int_{X}u^{2p+1}|\nabla_{\phi}\log\mathrm{v}|_{\phi}^{2}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}-\int_{X}u^{2p-1}|\nabla_{\phi}u|_{\phi}^{2}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
+$-2a\int_{X}u^{2p+1}|\nabla^{\phi}\log\mathrm{v}|_{\phi}^{2}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}-\int_{X}u^{2p-1}|\nabla^{\phi}u|_{\phi}^{2}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
 $+a\int_{X}u^{2p+1}\frac{\mathrm{v}_{,\alpha\beta}}{\mathrm{v}}\langle\xi_{\alpha},\xi_{\beta}\rangle_{\phi}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}.$
 Using the fact that $|S(\omega)|<C$, and by choosing $b>1$ sufficiently large so that $ab-aA_{1}-B\geq\frac{ab}{2}$ and $|a(f_{t}-bn-C)|\leq(n+1)ab$, the inequality above can be simplified into
-$(2p+1-a)\int_{X}u^{2p-1}\cdot|\nabla_{\phi}u|_{\phi}^{2}\cdot e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
+$(2p+1-a)\int_{X}u^{2p-1}\cdot|\nabla^{\phi}u|_{\phi}^{2}\cdot e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
 $\leq$ $(n+1)ab\int_{X}u^{2p+1}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}-a(1-\delta)\int_{X}u^{2p+1}\Delta_{\phi}f_{*}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
@@
 where we use the fact (2.20) which is ensured by the log-concavity of $\mathrm{v}$, and
-$-2a\int_{X}u^{2p}\langle\nabla_{\phi}u,\nabla_{\phi}\log\mathrm{v}\rangle_{\phi}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
-$\leq$ ${a}\int_{X}u^{2p+1}|\nabla_{\phi}\log\mathrm{v}|_{\phi}^{2}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}+a\int_{X}u^{2p-1}|\nabla_{\phi}u|_{\phi}^{2}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}.$
+$-2a\int_{X}u^{2p}\langle\nabla^{\phi}u,\nabla^{\phi}\log\mathrm{v}\rangle_{\phi}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
+$\leq$ ${a}\int_{X}u^{2p+1}|\nabla^{\phi}\log\mathrm{v}|_{\phi}^{2}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}+a\int_{X}u^{2p-1}|\nabla^{\phi}u|_{\phi}^{2}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}.$
 Next, we handle the term involving $\Delta(F-\log\mathrm{v})$ in the above inequality by repeating the integration by parts. Recall that $\omega_{\phi}^{n}=e^{F-\log\mathrm{v}}\omega^{n}$.
@@
 $-\int_{X}u^{2p}e^{(1-a)(F-\log\mathrm{v})-a\delta f_{*}-ab\phi}\cdot\frac{a\delta\Delta f_{*}+ab\Delta\phi}{1-a}\omega^{n}$
-=:absent:=:= : $I+II.$
+$=:$ $I+II.$
 One has
@@
 $=$ $\frac{p^{2}}{a-1}\int_{X}u^{2p-2}e^{-a(F+\log\mathrm{v}+\delta f_{*}+b\phi)}\cdot|\nabla u|^{2}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
-$\leq$ $\frac{p^{2}}{a-1}\int_{X}u^{2p-2}e^{-a(F+\log\mathrm{v}+\delta f_{*}+b\phi)}\cdot|\nabla_{\phi}u|_{\phi}^{2}(n+\Delta\phi)e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
-$=$ $\frac{p^{2}}{a-1}\int_{X}u^{2p-1}|\nabla_{\phi}u|_{\phi}^{2}e^{2a\log\mathrm{v}}\omega_{\phi}^{n},$
+$\leq$ $\frac{p^{2}}{a-1}\int_{X}u^{2p-2}e^{-a(F+\log\mathrm{v}+\delta f_{*}+b\phi)}\cdot|\nabla^{\phi}u|_{\phi}^{2}(n+\Delta\phi)e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
+$=$ $\frac{p^{2}}{a-1}\int_{X}u^{2p-1}|\nabla^{\phi}u|_{\phi}^{2}e^{2a\log\mathrm{v}}\omega_{\phi}^{n},$
 where we have used
@@
 Then the inequality (3.12) can be simplified into
-$\int_{X}(2p+1-a)u^{2p-1}\cdot|\nabla_{\phi}u|_{\phi}^{2}\cdot e^{2a\log\mathrm{v}}\omega_{\phi}^{n}+\frac{ab}{2}\int_{X}u^{2p+1}\mathrm{tr}_{\phi}\omega e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
+$\int_{X}(2p+1-a)u^{2p-1}\cdot|\nabla^{\phi}u|_{\phi}^{2}\cdot e^{2a\log\mathrm{v}}\omega_{\phi}^{n}+\frac{ab}{2}\int_{X}u^{2p+1}\mathrm{tr}_{\phi}\omega e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
 $\leq$ $(n+1)ab\int_{X}u^{2p+1}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}+C\int_{X}u^{2p}e^{-a(F+\log\mathrm{v}+\delta f_{*}+b\phi)}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
@@
 $\sup_{X}(|\nabla^{\phi}(F-\log\mathrm{v}(m_{\phi})+f_{*})|_{\phi})\leq C,$
-where $C$ is a constant depending on $\omega$, $[\theta]$, $\mathrm{v}$, $\mathrm{w}$, $\|\phi\|_{C^{0}}$, $\|F+f_{*}\|_{C^{0}}$, $\|\mathrm{tr}\omega_{\phi}\|_{L^{p}}$ for some $p$.
+where $C$ is a constant depending on $\omega$, $\theta]$, $\mathrm{v}$, $\mathrm{w}$, $\|\phi\|_{C^{0}}$, $\|F+f_{*}\|_{C^{0}}$, $\|\mathrm{tr}\omega_{\phi}\|_{L^{p}}$ for some $p$.
 Furthermore, for any $p>0$, with a constant $C_{p}$ that depends on the same parameters above,
@@
 $\geq-Ce^{-F^{\prime}}(\mathrm{tr}\omega_{\phi})^{n-1}|\nabla^{\phi}W|_{\phi}^{2}-\frac{W_{i\overline{j}}W_{\overline{i}}W_{j}}{(1+\phi_{i\overline{i}})(1+\phi_{j\overline{j}})}.$
-Since $f_{*}$ is a $(\frac{1-t}{t})\theta$-psh function , $\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha}(f_{*})$ is determined by class $[\theta]$, which is bounded.
+Since $f_{*}$ is a $(\frac{1-t}{t})\theta$-psh function , $\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha}(f_{*})$ is determined by class $\theta]$, which is bounded.
 And by Theorem 3.2, $|W|_{C^{0}}$ is bounded.
@@
 $-C(1+e^{-F^{\prime}}(\mathrm{tr}\omega_{\phi})^{n-1})\cdot|\nabla^{\phi}W|_{\phi}^{2}+\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha}(W)|\nabla^{\phi}W|_{\phi}^{2}.$ (3.16)
-Note that
+Note that (3.5) and (2.17),
 $\Delta_{\phi}W=$ $\Delta_{\phi}(F+f_{*})-\Delta_{\phi}\log\mathrm{v}$
 $=$ $\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha}(F+f_{*}-\log\mathrm{v})+\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha}\log\mathrm{v}-\Delta_{\phi}\log\mathrm{v}+\widetilde{f}_{t}+\mathrm{tr}_{\phi}\eta_{t}$
-$=$ $\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha}(W)+\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha}(F^{\prime}+f_{*})-\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha}(f_{*})+\left(\frac{\mathrm{v}_{,\alpha}\mathrm{v}_{,\beta}}{\mathrm{v}^{2}}-\frac{\mathrm{v}_{,\alpha\beta}}{\mathrm{v}}\right)\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}$ (3.15)
+$=$ $\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha}(W)+\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha}(F^{\prime}+f_{*})-\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha}(f_{*})+\left(\frac{\mathrm{v}_{,\alpha}\mathrm{v}_{,\beta}}{\mathrm{v}^{2}}-\frac{\mathrm{v}_{,\alpha\beta}}{\mathrm{v}}\right)\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}$
 $-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}\Delta_{\omega}(m_{\omega}^{\xi^{\alpha}})-\frac{\mathrm{v}_{,\alpha}\mathrm{v}_{,\beta}}{\mathrm{v}^{2}}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}+\widetilde{f}_{t}+\mathrm{tr}_{\phi}\eta_{t},$
-$2\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha}(W)-C-C\mathrm{tr}_{\phi}\omega-C\mathrm{tr}\omega_{\phi}\leq\Delta_{\phi}W\leq 2\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha}(W)+C+C\mathrm{tr}_{\phi}\omega+C\mathrm{tr}\omega_{\phi},$
-where we use (3.4) and (2.16) for the second equality, (2.18) for the last inequality.
+which together with (2.19) implies
+$2\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha}(W)-C-C\mathrm{tr}_{\phi}\omega-C\mathrm{tr}\omega_{\phi}\leq\Delta_{\phi}W\leq 2\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha}(W)+C+C\mathrm{tr}_{\phi}\omega+C\mathrm{tr}\omega_{\phi}$ (3.17)
 To control the derivative term $J\xi^{\alpha}(W)|\nabla^{\phi}W|_{\phi}^{2}$, we need to add another barrier function. One computes
@@
 $=$ $e^{\frac{W}{2}-\log\mathrm{v}(m_{\phi})}|\nabla^{\phi}W|_{\phi}^{2}|\nabla^{\phi}\log\mathrm{v}(m_{\phi})|_{\phi}^{2}-2e^{-\log\mathrm{v}(m_{\phi})}\nabla^{\phi}\log\mathrm{v}(m_{\phi})\cdot_{\phi}\nabla^{\phi}(e^{\frac{W}{2}}|\nabla^{\phi}W|_{\phi}^{2})$
-$+e^{-\log\mathrm{v}(m_{\phi})}\left(\left(\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}(W)-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(\phi)}J\xi^{\alpha}(f_{*})-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}\Delta_{\omega}(m_{\omega}^{\xi^{\alpha}})\right.\right.$
+[…]
 $\left.\left.-\frac{\mathrm{v}_{,\alpha\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}+\frac{\mathrm{v}_{,\alpha}(m_{\phi})\mathrm{v}_{,\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})^{2}}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}\right)e^{\frac{W}{2}}|\nabla^{\phi}W|_{\phi}^{2}+\Delta_{\phi}(e^{\frac{W}{2}}|\nabla^{\phi}W|_{\phi}^{2})\right)$
@@
 In the inequality above, we use $\omega^{n}=e^{f_{*}}\cdot e^{-(F^{\prime}+f_{*})}\omega_{\phi}^{n}\leq C\omega_{\phi}^{n}$, since $F^{\prime}+f_{*}$ is uniformly bounded, and $f_{*}\leq 0$.
-For any $20<\varepsilon<$ 2, by Hölder inequality, we have
+For any $0<\varepsilon<2$, by Hölder inequality, we have
 $\int_{X}(\mathrm{tr}\omega_{\phi})^{3(n-1)}u^{2p+1}\omega_{\phi}^{n}$ $=\int_{X}(\mathrm{tr}\omega_{\phi})^{3(n-1)}u^{2p+1}e^{-f_{*}}e^{F^{\prime}+f_{*}}\omega^{n}$
@@
 $C_{0}=\int_{X}\log(\mathrm{v}(m_{\omega}))\mathrm{v}(m_{\omega})\omega^{n},$
-$\int_{X}\mathrm{tr}_{\mathrm{v},\omega}(\theta)\frac{\omega^{n}}{n!}=\int_{X}\mathrm{v}(m_{\omega})\theta\wedge\frac{\omega^{n-1}}{(n-1)!}+\int_{X}\langle d\mathrm{v}(m_{\omega}),m_{\theta}\rangle\frac{\omega^{n}}{n!}=\underline{\theta}\cdot\int_{X}\mathrm{w}(m_{\omega})\frac{\omega^{n}}{n!}.$
+and $\underline{\theta}$ is a contant determined by the equality:
+$\int_{X}\mathrm{tr}_{\mathrm{v},\omega}(\theta)\frac{\omega^{n}}{n!}=\int_{X}\mathrm{v}(m_{\omega})\theta\wedge\frac{\omega^{n-1}}{(n-1)!}+\int_{X}\langle\mathrm{d}\mathrm{v}(m_{\omega}),m_{\theta}\rangle\frac{\omega^{n}}{n!}=\underline{\theta}\cdot\int_{X}\mathrm{w}(m_{\omega})\frac{\omega^{n}}{n!}.$ (4.2)
+In particular, $\underline{\theta}$ only depends on $\omega]$ ([Lah19, Lemma 3]).
 Denote ${\mathbb{J}}^{\mathrm{Ric}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}}}(\phi)={\mathbb{E}}_{\mathrm{v}}^{\mathrm{Ric}}(\phi)-{\mathbb{E}}_{\mathrm{w}\cdot\ell_{\mathrm{ext}}}(\phi)$, ${\mathbb{J}}^{\theta}_{\mathrm{v},\mathrm{w}}(\phi)={\mathbb{E}}_{\mathrm{v}}^{\theta}(\phi)-\underline{\theta}\cdot{\mathbb{E}}_{\mathrm{w}}(\phi)$.
@@
 4.A. Openness
-In this subsection, we show the openness of the continuity path (1.3), following the standard implicit function theory ([Che18], [Has19]). Consider
+In this subsection, we show the openness of the continuity path (1.3), following the standard implicit function theory ([Ch18], [Has19]).
+Firstly, note that as explained in the proof of Proposition 4.2 (1), we can choose $K$-invariant $\theta\in[\omega]$ such that $\phi=0$ is the solution of continuity path at $t=0$.
+Consider
 ${\mathcal{H}}^{4,\alpha}_{K}:=\{\phi\in C_{K}^{4,\alpha}(X)~|~\omega_{\phi}>0\}$
@@
 Suppose there exists $(\phi_{0},t_{0})\in{\mathcal{H}}^{4,\alpha}_{K}\times[0,1]$ such that ${\mathcal{F}}(\phi_{0},t_{0})=0$. We will consider the following modified map with a normalization:
-~$\widetilde{{\mathcal{F}}}:{\mathcal{H}}^{4,\alpha}_{K,0}\times[0,1]$ $\rightarrow C^{\alpha}(X)_{K,0}$
+$\widetilde{{\mathcal{F}}}:{\mathcal{H}}^{4,\alpha}_{K,0}\times[0,1]$ $\rightarrow C^{\alpha}(X)_{K,0}$
 $(\phi,t)$ $\mapsto{\mathcal{F}}(\phi,t)-\int_{X}{\mathcal{F}}(\phi,t)\omega_{\phi_{0}}^{n},$
@@
 ${\mathcal{H}}^{4,\alpha}_{K,0}:=\{\phi\in{\mathcal{H}}_{K}^{4,\alpha}~|~\int_{X}\phi\omega_{\phi_{0}}^{n}=0\}.$
-Note that $\int_{X}{\mathcal{F}}(\phi,t)\omega_{\phi}^{n}=0$.
-Therefore if ~$\widetilde{{\mathcal{F}}}(\phi,t)=0$, ${\mathcal{F}}(\phi,t)$=0.
+Note that $\int_{X}{\mathcal{F}}(\phi,t)\mathrm{w}(m_{\phi})\omega_{\phi}^{n}=0$ by the definition of the constant $\underline{\theta}$ and extremal function $\ell_{\mathrm{ext}}$.
+Therefore if $\widetilde{{\mathcal{F}}}(\phi,t)=0$, ${\mathcal{F}}(\phi,t)$=0.
+This is because $\widetilde{{\mathcal{F}}}(\phi,t)=0$ implies ${\mathcal{F}}(\phi,t)\equiv C$, then by integrating with respect to the weighted measure $\mathrm{w}(m_{\phi})\omega_{\phi}^{n}$, we obtain that $C$ must be zero.
 Thus it suffices to apply implicit function theorem on the map $\widetilde{{\mathcal{F}}}$ near $(\phi_{0},t_{0})$.
 Denote
-(ϕ):=assignitalic-$\mathcal{R}(\phi):=$ $\frac{S_{\mathrm{v}}(\phi)}{\mathrm{w}(m_{\phi})}-\ell_{\rm ext}(m_{\phi}),\text{ and }\mathcal{T}(\phi):=\frac{\mathrm{tr}_{\mathrm{v},\phi}(\theta)}{\mathrm{w}(m_{\phi})}.$
+$\mathcal{R}(\phi):=$ $\frac{S_{\mathrm{v}}(\phi)}{\mathrm{w}(m_{\phi})}-\ell_{\rm ext}(m_{\phi}),\text{ and }\mathcal{T}(\phi):=\frac{\mathrm{tr}_{\mathrm{v},\phi}(\theta)}{\mathrm{w}(m_{\phi})}.$
 Thus
@@
 $\begin{split}t(D_{\phi}{\mathcal{R}})(u)=&-t\frac{(\mathcal{D}^{*}_{\phi}\mathrm{v}(m_{\phi})\mathcal{D}_{\phi})u}{\mathrm{w}(m_{\phi})}+t\langle\mathrm{d}({\mathcal{R}}(\phi)),\mathrm{d}u\rangle_{\phi},\\ =&-t\frac{(\mathcal{D}^{*}_{\phi}\mathrm{v}(m_{\phi})\mathcal{D}_{\phi})u}{\mathrm{w}(m_{\phi})}+\langle\mathrm{d}({\mathcal{F}}(\phi,t)),\mathrm{d}u\rangle_{\phi}+(1-t)\langle\mathrm{d}(\mathcal{T}(\phi)),\mathrm{d}u\rangle_{\phi},\end{split}$ (4.4)
-where $\mathcal{D}_{\phi}u:=\sqrt{2}(\nabla^{\phi}\mathrm{d}u)^{-}$ is the $J$-anti-invariant part of the tensor $(\nabla^{\phi}\mathrm{d}u)$, with $\nabla^{\phi}$ the $g_{\phi}$-Levi-Civita connection, and $\mathcal{D}^{*}_{\phi}$ is the formal adjoint of $\mathcal{D}_{\phi}$.
+where $\mathcal{D}_{\phi}u:=\sqrt{2}(\nabla^{\phi}\mathrm{d}u)^{-}$ is the $J$-anti-invariant part of the tensor $(\nabla^{\phi}\mathrm{d}u)$, with $\nabla^{\phi}$ the $g_{\phi}$-Levi-Civita connection, and $\mathcal{D}^{*}_{\phi}$ is the formal adjoint of $\mathcal{D}_{\phi}$ with respect to $(\cdot,\cdot)_{\phi}$, where
+$(f,h)_{\phi}=\int_{X}f\;h\;\omega_{\phi}^{n},$
+for any $f,h\in C^{4,\alpha}(X)_{K,0}$.
 We denote the operator $F_{\phi}:C^{4,\alpha}(X)_{K,0}\to C^{\alpha}(X)_{K,0}$ by
@@
 $D_{(\phi_{0},t_{0})}\widetilde{{\mathcal{F}}}(u)={\mathcal{L}}_{(\phi_{0},t_{0})}u+s({\mathcal{R}}(\phi_{0})+{\mathcal{T}}(\phi_{0})-\underline{\theta}).$
-By Lemma 5.4, we have that the operator ${\mathcal{L}}_{(\phi_{0},t_{0})}$ is a $(\cdot,\cdot)_{\mathrm{w},\phi}$-self-adjoint $K$-invariant linear operator satisfying
-$({\mathcal{L}}_{(\phi_{0},t_{0})}u_{1},u_{2})_{\mathrm{w},\phi}=$ $\int_{X}{\mathcal{L}}_{(\phi_{0},t_{0})}u_{1}\cdot u_{2}\mathrm{w}(m_{\phi_{0}})\omega_{\phi_{0}}^{n}$
+By Lemma 5.4, we have that the operator ${\mathcal{L}}_{(\phi_{0},t_{0})}$ is a $(\cdot,\cdot)_{\mathrm{w},\phi_{0}}$-self-adjoint $K$-invariant linear operator satisfying
+$({\mathcal{L}}_{(\phi_{0},t_{0})}u_{1},u_{2})_{\mathrm{w},\phi_{0}}=$ $\int_{X}{\mathcal{L}}_{(\phi_{0},t_{0})}u_{1}\cdot u_{2}\mathrm{w}(m_{\phi_{0}})\omega_{\phi_{0}}^{n}$
 $=$ $-t_{0}\int_{X}\mathrm{v}(m_{\phi_{0}}){\mathcal{D}}_{\phi_{0}}u_{1}{\mathcal{D}}_{\phi_{0}}u_{2}\omega_{\phi_{0}}^{n}-(1-t_{0})\int_{X}\langle\theta,\mathrm{d}u_{1}\wedge\mathrm{d}^{c}u_{2}\rangle_{\phi_{0}}\mathrm{v}(m_{\phi_{0}})\omega_{\phi_{0}}^{n}$
-$=$ $(u_{1},{\mathcal{L}}_{(\phi_{0},t_{0})}u_{2})_{\mathrm{w},\phi}.$
-When $t_{0}>0$, following the argument in [Che18], we divide the remaining argument into three steps:
+$=$ $(u_{1},{\mathcal{L}}_{(\phi_{0},t_{0})}u_{2})_{\mathrm{w},\phi_{0}}.$
+When $t_{0}>0$, following the argument in [Ch18], we divide the remaining argument into three steps:
 Step 1: show the inverse operator ${\mathcal{L}}_{(\phi_{0},t_{0})}^{-1}$ is bounded, that is, if $u\in C_{K}^{4,\alpha}(X)$ satisfies the equation ${\mathcal{L}}_{(\phi_{0},t_{0})}u=f$ for some $f\in C_{K}^{\alpha}(X)$ and $\int_{X}u\omega_{\phi_{0}}^{n}=0$, then there exists a constant $C$ depending on $\omega_{\phi_{0}}$ such that
 $\|u\|_{C^{4,\alpha}(X)}\leq C\|f\|_{C^{\alpha}(X)}.$
-The proof follows from [Che18, Lemma 4.3] by using Schauder estimates, Poincaré inequality.
+The proof follows from [Ch18, Lemma 4.3] by using Schauder estimates, Poincaré inequality.
 Step 2: Show that $D_{(\phi_{0},t_{0})}\widetilde{{\mathcal{F}}}(\cdot,0)={\mathcal{L}}_{(\phi_{0},t_{0})}(\cdot):C^{4,\alpha}(X)_{K,0}\rightarrow C^{\alpha}(X)_{K,0}$ is an isomorphism.
-Similar to [Che18, Lemma 4.4], the proof is given by a standard Fredholm theorem for elliptic operators.
+Similar to [Ch18, Lemma 4.4], the proof is given by a standard Fredholm theorem for elliptic operators.
 Step 3: By the standard implicit function theory, there exists $\phi(t)\in{\mathcal{H}}^{4,\alpha}_{K}$ such that ${\mathcal{F}}(\phi(t),t)=0$.
@@
 $\phi$ $\mapsto\frac{S_{\mathrm{v}}(\phi)}{\mathrm{w}(m_{\phi})}-\ell_{\mathrm{ext}}(m_{\phi})-R\left(\frac{\mathrm{tr}_{\mathrm{v},\phi}(\theta)}{\mathrm{w}(m_{\phi})}-\underline{\theta}\right)$
-$\qquad-\int_{X}\frac{S_{\mathrm{v}}(\phi)}{\mathrm{w}(m_{\phi})}-\ell_{\mathrm{ext}}(m_{\phi})-r\left(\frac{\mathrm{tr}_{\mathrm{v},\phi}(\theta)}{\mathrm{w}(m_{\phi})}-\underline{\theta}\right)\omega_{\phi_{0}}^{n}$
+$\hskip 18.49988pt-\int_{X}\frac{S_{\mathrm{v}}(\phi)}{\mathrm{w}(m_{\phi})}-\ell_{\mathrm{ext}}(m_{\phi})-R\left(\frac{\mathrm{tr}_{\mathrm{v},\phi}(\theta)}{\mathrm{w}(m_{\phi})}-\underline{\theta}\right)\omega_{\phi_{0}}^{n}$
 for some $R>>0$. Its linearization is
-${\mathcal{L}}_{0,\phi_{0}}(u):=(D_{\phi_{0}}{\mathcal{F}}_{0})(u)=-\frac{(\mathcal{D}^{*}_{\phi_{0}}\mathrm{v}(m_{\phi_{0}})\mathcal{D}_{\phi_{0}})u}{\mathrm{w}(m_{\phi_{0}})}+rF_{\phi_{0}}(u),$
-which is a $(\cdot,\cdot)_{\mathrm{w},\phi}$-self-adjoint $K$-invariant linear operator with a trivial kernal (see Lemma 5.4).
+${\mathcal{L}}_{0,\phi_{0}}(u):=(D_{\phi_{0}}{\mathcal{F}}_{0})(u)=-\frac{(\mathcal{D}^{*}_{\phi_{0}}\mathrm{v}(m_{\phi_{0}})\mathcal{D}_{\phi_{0}})u}{\mathrm{w}(m_{\phi_{0}})}+RF_{\phi_{0}}(u),$ (4.5)
+which is a $(\cdot,\cdot)_{\mathrm{w},\phi_{0}}$-self-adjoint $K$-invariant linear operator with a trivial kernal (see Lemma 5.4).
 By the detailed computation of the linearized operators in Appendix 5.B, we find that $F_{\phi_{0}}$ has a negative first eigenvalue.
-Thus, for $r$ sufficiently large, the first eigenvalue $\lambda_{1,r}$ of ${\mathcal{L}}_{0,\phi_{0}}$ is always negative.
+Thus, the first eigenvalue $\lambda_{1,R}$ of ${\mathcal{L}}_{0,\phi_{0}}$ is always negative for $R>0$.
+Indeed, let $\phi_{1,R}$ be the first eigenfunction of ${\mathcal{L}}_{0,\phi_{0}}$, one has
+$\lambda_{1,R}=$ $\frac{({\mathcal{L}}_{0,\phi_{0}}(\phi_{1,R}),\phi_{1,R})_{\mathrm{w},\phi_{0}}}{(\phi_{1,R},\phi_{1,R})_{\mathrm{w},\phi_{0}}}$
+$=$ $\frac{1}{(\phi_{1,R},\phi_{1,R})_{\mathrm{w},\phi_{0}}}\left(-\int_{X}\phi_{1,R}({\mathcal{D}}_{\phi_{0}}^{*}\mathrm{v}(m_{\phi_{0}}){\mathcal{D}}_{\phi_{0}})\phi_{1,R}\omega_{\phi_{0}}^{n}+R(F_{\phi_{0}}(\phi_{1,R}),\phi_{1,R})_{\mathrm{w},\phi_{0}}\right)$
+$=$ $\frac{-1}{(\phi_{1,R},\phi_{1,R})_{\mathrm{w},\phi_{0}}}\int_{X}\mathrm{v}(m_{\phi_{0}})|{\mathcal{D}}_{\phi_{0}}\phi_{1,R}|^{2}_{\phi_{0}}\omega_{\phi_{0}}^{n}+R\frac{(F_{\phi_{0}}(\phi_{1,R}),\phi_{1,R})_{\mathrm{w},\phi_{0}}}{(\phi_{1,R},\phi_{1,R})_{\mathrm{w},\phi_{0}}}$
+$\leq$ $R\lambda_{1}(F_{\phi_{0}})\leq-C\cdot R<0.$ (4.6)
 At the end, the implicit function theorem argument in [Has19] implies that there exists a solution $\phi\in{\mathcal{H}}_{K}$ of ${\mathcal{F}}_{0}(\phi)=0$ for $R$ large enough.
-Combining the two cases of $t_{0}=0$ and $10<t_{0}<$ 1, we obtain the following openness property.
+We sketch the main steps of Hashimoto’s argument as follows.
+  1. (i)
+    As stated in (4.8), we can find a solution $\omega$ for the continuity path at $t=0$. Perturb $\omega$ to get
+    $\omega_{m}:=\omega+{\rm dd^{c}}(R^{-1}\phi_{1}+\cdots+R^{-m}\phi_{m})$
+    for some $\phi_{1},\ldots,\phi_{m}\in C^{\infty}(X)_{K}$, satisfying ${\mathcal{R}}(\omega_{m})-R{\mathcal{T}}(\omega_{m})=const+R^{-m}f_{m,R}$ for a function $f_{m,R}\in C^{\infty}(X)_{K}$ with $(f_{m,R},1)_{\mathrm{w},\phi_{0}}=0$ which is bounded in $C^{\infty}(X)_{K}$ for all sufficiently large $R$.
+The proof follows from [Has19, Lemma 3].
+    Then by solving a weighted Green’s function $\frac{\mathrm{d}^{*}(\mathrm{v}(m_{\omega_{m}})\mathrm{d}G_{m,R})}{\mathrm{w}(m_{\omega_{m}})}=f_{m,R}$, we can find that $\omega_{m}$ satisfies
+    ${\mathcal{R}}(\omega_{m})-R\frac{\mathrm{tr}_{\mathrm{v},\omega_{m}}\theta_{m}}{\mathrm{w}(m_{\omega_{m}})}=const$ (4.7)
+    where $\theta_{m}:=\theta+R^{-m-1}{\rm dd^{c}}G_{m,R}$.
+  2. (ii)
+    Consider the Banach spaces $B_{1}=\Omega^{1,1}(X)_{K}\times L^{2}_{p+4}(X)_{K,0}$ and $B_{2}=\Omega^{1,1}(X)_{K}\times L^{2}_{p}(X)_{K,0}$, and an open set $U:=\{(\chi,\phi)\in B_{1}~|~\omega_{m,\phi}:=\omega_{m}+{\rm dd^{c}}\phi>0\}$, where $\Omega^{1,1}(X)_{K}$ is the $L^{2}_{p}(X)$ Sobolev completion of the space of $K$-invariant real $(1,1)$-forms and $L^{2}_{p}(X)_{K,0}$ is the Sobolev completion of $C^{\infty}(X)_{K}$ with $(\phi,1)_{\mathrm{w},\phi_{0}}=0$.
+We can define a map $T:U\rightarrow B_{2}$ by
+    $T(\chi,\phi):=\left(\theta_{m}+\chi,{\mathcal{R}}(\omega_{m,\phi})-R\frac{\mathrm{tr}_{\mathrm{v},\omega_{m,\phi}}(\theta_{m}+\chi)}{\mathrm{w}(m_{\omega_{m,\phi}})}\right)$
+    Using (4.7), the linearization of $T$ at 0 is given by
+    $(D_{0}T)(\chi,u)=\left(\begin{array}[]{cc}1&0\\ -R\frac{\mathrm{tr}_{\mathrm{v},\omega_{m}}}{\mathrm{w}(m_{\omega_{m}})}&{\mathcal{L}}_{0,\omega_{m}}\end{array}\right)\left(\begin{array}[]{c}\chi\\ u\end{array}\right),$
+    where ${\mathcal{L}}_{0,\omega_{m}}$ is defined as in (4.5), where we substitute $\theta$ by $\theta_{m}$ and substitute $\omega_{\phi_{0}}$ by $\omega_{m}$.
+    By Appendix 5.B, we have ${\mathcal{L}}_{0,\omega_{m}}$ is a $(\cdot,\cdot)_{\mathrm{w},\omega_{m}}$-self-adjoint linear operator with trivial kernel.
+Then by Fredholm theory, we have that $D_{0}T:B_{1}\rightarrow B_{2}$ is an isomorphism with the inverse
+    $P=\left(\begin{array}[]{cc}1&0\\ \left({\mathcal{L}}_{0,\omega_{m}}\right)^{-1}\left(\frac{-R\mathrm{tr}_{\mathrm{v},\omega_{m}}}{\mathrm{w}(m_{\omega_{m}})}\right)&\left({\mathcal{L}}_{0,\omega_{m}}\right)^{-1}\end{array}\right).$
+    By the computation of (4.6) and Appendix 5.B (see also [Has19, lemma 4]), there is a constant $C=C(\omega,\theta)>0$ such that $\lambda_{1,m}<-CR$ for all sufficiently large $R$, where $\lambda_{1,m}<0$ is the largest non-zero eigenvalue of ${\mathcal{L}}_{0,\omega_{m}}$.
+By the above definition of $P$, one has $\|P\|_{\rm op}\leq CR^{2}$ (see [Has19, Below of Lemma 4]).
+  3. (iii)
+    By the definition of $T$, for $l\in{\mathbb{N}}$ and $l\geq 5$, for any $R$ sufficiently large, the higher order part $T-D_{0}T$ is Lipschitz with Lipschitz constant $1/(2\|P\|_{\rm op})$ in a ball centered at 0 of radius $\delta^{\prime}=R^{-l}$.
+  4. (iv)
+    We choose $\delta=\frac{\delta^{\prime}}{2\|P\|_{\rm op}}>\frac{1}{2C}R^{-l-2}$ for latter use.
+  5. (v)
+    Note that $T(0,0)=(\theta_{m},-R\underline{\theta})$. Since $\theta_{m}-\theta=O(R^{-m-1})$, we obtain
+    $\|T(0,0)-(\theta,-R\underline{\theta})\|_{L^{2}_{p}}\leq CR^{-m-1}\leq\frac{1}{2C}R^{-l-2}<\delta,$
+    for all large enough $R>0$, by taking $m$ to be sufficiently large.
+Choose $p>0$ to be sufficiently large such that $C^{4}(X)\hookrightarrow L^{2}_{p}(X)$.
+By the argument above, the assumptions of the quantitative inverse function theorem (see [Has19, Theorem 2]) are satisfied.
+By applying this theorem, there exists $(\chi,\phi)\in U$ such that $\theta_{m}+\chi=\theta$ and
+    ${\mathcal{R}}(\omega_{m,\phi})-R\frac{\mathrm{tr}_{\mathrm{v},\omega_{m,\phi}}(\theta)}{\mathrm{w}(m_{\omega_{m,\phi}})}=R\underline{\theta}$
+    for all large enough $R>0$. The standard elliptic PDE theory shows that $\phi$ is smooth.
+Combining the two cases of $t_{0}=0$ and $0<t_{0}<1$, we obtain the following openness property.
 Theorem 4.1.
@@
     The weighted Mabuchi functional ${\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}}}$ is bounded from below.
-Furthermore, ${\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}},t}$ is coercive for $10<t<$ 1.
+Furthermore, ${\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}},t}$ is coercive for $0<t<1$.
   2. $(2)$
@@
 We take a suitable choice of $\theta:=\omega+\sqrt{-1}\partial\overline{\partial}\psi_{0}$ such that
-    $\frac{\mathrm{tr}_{\mathrm{v},\omega}\theta}{\mathrm{w}(m_{\omega})}=\underline{\theta}.$
+    $\frac{\mathrm{tr}_{\mathrm{v},\omega}\theta}{\mathrm{w}(m_{\omega})}=\underline{\theta}.$ (4.8)
+    Indeed, the equation (4.8) can be rewritten as
+    $\frac{\mathrm{d}^{*}(\mathrm{v}(m_{\omega})\mathrm{d}\psi_{0})}{\mathrm{w}(m_{\omega})}=\frac{\mathrm{tr}_{\mathrm{v},\omega}(\omega)}{\mathrm{w}(m_{\omega})}-\underline{\theta}.$
+    Note that $\frac{\mathrm{d}^{*}(\mathrm{v}(m_{\omega})\mathrm{d}\psi_{0})}{\mathrm{w}(m_{\omega})}$ a second order $(\cdot,\cdot)_{\mathrm{w},\omega}$-self-adjoint elliptic operator.
+By standard elliptic theory it follows that there exist a solution $\psi_{0}\in C^{\infty}(X)_{K}$ of (4.8).
     Then $\theta\geq\varepsilon\omega$ for some $\varepsilon>0$. Recall that for any $\phi\in{\mathcal{H}}_{K,0}$,
@@
 Recall that $\phi_{i}=\sigma_{i}^{*}\varphi_{i}$. Then
-$\sigma_{i}^{*}(\omega_{\varphi_{i}}^{n})=(\sigma_{i}^{*}\omega_{\varphi_{i}})^{n}=\omega_{\phi_{i}}^{n},\quad\sigma_{i}^{*}(\mathrm{v}(m_{\varphi_{i}})\omega_{\varphi_{i}}^{n})=\mathrm{v}(m_{\phi_{i}})\omega_{\phi_{i}}^{n}$
+$\sigma_{i}^{*}(\omega_{\varphi_{i}}^{n})=(\sigma_{i}^{*}\omega_{\varphi_{i}})^{n}=\omega_{\phi_{i}}^{n},\qquad\sigma_{i}^{*}(\mathrm{v}(m_{\varphi_{i}})\omega_{\varphi_{i}}^{n})=\mathrm{v}(m_{\phi_{i}})\omega_{\phi_{i}}^{n}$
 and
@@
 Take $\sigma_{i}^{*}$-action on the second equation (2.14b), one has
-[…]
+$\text{L.H.S.}=\sigma_{i}^{*}\left(\left(\Delta_{\varphi_{i}}-\frac{\mathrm{v}_{,\alpha}(m_{\varphi_{i}})}{\mathrm{v}(m_{\varphi_{i}})}J\xi^{\alpha}\right)F_{i}\right)=\left(\Delta_{\phi_{i}}-\frac{\mathrm{v}_{,\alpha}(m_{\phi_{i}})}{\mathrm{v}(m_{\phi_{i}})}J\xi^{\alpha}\right)F_{i}\circ\sigma_{i},$ (4.14)
 and
-[…]
+$\text{R.H.S.}=-\ell_{\mathrm{ext}}(m_{\phi_{i}})\frac{\mathrm{w}(m_{\phi_{i}})}{\mathrm{v}(m_{\phi_{i}})}+(1-\frac{1}{t})\frac{\mathrm{tr}_{\mathrm{v},\phi_{i}}(\sigma_{i}^{*}\theta)}{\mathrm{v}(m_{\phi_{i}})}-(1-\frac{1}{t})\underline{\theta}\frac{\mathrm{w}(m_{\phi_{i}})}{\mathrm{v}(m_{\phi_{i}})}+\frac{\mathrm{tr}_{\mathrm{v},\phi_{i}}\mathrm{Ric}(\sigma_{i}^{*}\omega)}{\mathrm{v}(m_{\phi_{i}})}.$ (4.15)
 On the other hand,
@@
 Proposition 4.8.
-Let $f_{i}=\frac{1-t_{i}}{t_{i}}u_{i}$.
-Let $\alpha=\alpha(X,\omega)$ be the alpha-invariant of the Kähler class$[\omega]$.
-Then there exists $C>0$ such that $\int_{X}e^{-p\cdot f_{i}}\omega^{n}<C$ for any $p<\frac{t_{i}\alpha}{1-t_{i}}$.
-Moreover, we have $e^{-f_{i}}\rightarrow 1$ in $L^{p}(\omega^{n})$ as $t_{i}\rightarrow 1$ for any $p<\infty$.
+[…]
 Proof.
@@
 Hence we may take a subsequence of $\phi_{i}$ (without relabeled), and a function $\phi_{*}\in W^{2,p}$ for any $p<\infty$, and another function $F_{*}\in W^{1,p}$ for any $p<\infty$, such that
-$\phi_{i}\rightarrow\phi_{*}\text{ in }C^{1,\alpha}\text{ for any }0<\alpha<1\text{ and }\sqrt{-1}\partial\overline{\partial}\phi_{i}\rightarrow\sqrt{-1}\partial\overline{\partial}\phi_{*}\text{ weakly in }L^{p},$ (4.12)
+[…]
 $\widetilde{F}_{i}-\log\mathrm{v}(m_{\phi_{i}})+f_{i}\rightarrow F_{*}\text{ in }C^{\alpha}\text{ for any }0<\alpha<1\text{ and }$ (4.18)
-$\qquad\qquad\qquad\qquad\nabla(\widetilde{F}_{i}-\log\mathrm{v}(m_{\phi_{i}})+f_{i})\rightarrow\nabla F_{*}\text{ weakly in }L^{p}.$
+$\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\nabla(\widetilde{F}_{i}-\log\mathrm{v}(m_{\phi_{i}})+f_{i})\rightarrow\nabla F_{*}\text{ weakly in }L^{p}.$
 It follows from [CC21b, Lemma 4.22] that
-$\omega_{\phi_{i}}^{k}\rightarrow\omega_{\phi_{i}}^{k}\text{ weakly in }L^{p}\text{ for any }1\leq k\leq n\text{ and }p<\infty.$
+[…]
 Proposition 4.10.
@@
 We obtain the first equation $\omega_{\phi_{*}}^{n}=e^{F_{*}}\omega^{n}$.
-To obtain the second equation, by (2.10), we first re-write the equation (4.8) to
+To obtain the second equation, by (2.B), we first re-write the equation (4.13) to
 $\Delta_{\phi_{i}}(\widetilde{F}_{i}+f_{i}-\log\mathrm{v}(m_{\phi_{i}}))$
@@
 $\int_{X}W_{i}\sqrt{-1}\partial\overline{\partial}u\wedge\omega_{\phi_{i}}^{n-1}\rightarrow\int_{X}F_{*}\sqrt{-1}\partial\overline{\partial}u\wedge\omega_{\phi_{*}}^{n-1}.$
-For the right hand side, by (4.13), (4.12), $\phi_{i}\rightarrow\phi_{*}$ in $C^{1,\alpha}$ for any $10<\alpha<$ 1. Then following the construction in the proof of [HL23, Proposition 10], we have
+For the right hand side, by (4.18), (4.17), $\phi_{i}\rightarrow\phi_{*}$ in $C^{1,\alpha}$ for any $0<\alpha<1$. Then following the construction in the proof of [HL20, Proposition 10], we have
 $\frac{\mathrm{v}_{,\alpha}(m_{\phi_{i}})\mathrm{v}_{,\beta}(m_{\phi_{i}})}{\mathrm{v}(m_{\phi_{i}})^{2}}\omega_{\phi_{i}}^{n}\rightarrow\frac{\mathrm{v}_{,\alpha}(m_{\phi_{*}})\mathrm{v}_{,\beta}(m_{\phi_{*}})}{\mathrm{v}(m_{\phi_{*}})^{2}}\omega_{\phi_{*}}^{n}$
@@
 In the proof of Lemma 4.11, we have obtain that $\phi_{*}\in W^{3,p}$ for any $p<\infty$.
-Hence $\omega_{\phi_{*}}\in C^{\alpha}$ for any $10<\alpha<$ 1.
-By (4.14) and Schauder estimate, we derive $F_{*}\in C^{2,\alpha}$ for any $10<\alpha<$ 1.
+Hence $\omega_{\phi_{*}}\in C^{\alpha}$ for any $0<\alpha<1$.
+By (4.19) and Schauder estimate, we derive $F_{*}\in C^{2,\alpha}$ for any $0<\alpha<1$.
 Then the higher regularity follows from bootstrap (see Corollary 5.3).
@@
 where $A_{1}$, $B$, $C_{0}$ are constant
-$\mathrm{Bisec}(\omega)\geq-B,\quad\mathrm{Ric}(\omega)\leq A_{1}\omega,\quad\|F\|_{L^{\infty}}+\|\phi\|_{L^{\infty}}\leq C_{0}.$
+$\mathrm{Bisec}(\omega)\geq-B,\qquad\mathrm{Ric}(\omega)\leq A_{1}\omega,\qquad\|F\|_{L^{\infty}}+\|\phi\|_{L^{\infty}}\leq C_{0}.$
 Adapting the similar argument in the proof of Theorem 3.7, we obtain the gradient estimate of $F$ and $C^{2}$-estimate of $\phi$ for equations (1.2a), (1.2b).
@@
 By Schauder estimate, we deduce $\|\phi\|_{C^{3,\alpha^{\prime}}}\leq C_{4}$.
-Going back to (2.13b) again, the coefficient are in $C^{\gamma}$ for any $10<\gamma<$ 1, hence we have $\|F\|_{C^{1,\gamma}}\leq C_{5}$.
-Now, the coefficients of (5.8) are in $C^{\gamma}$ for any $10<\gamma<$ 1.
-Thus we derive $\|\phi\|_{C^{3,\gamma}}\leq C_{6}$ for any $10<\gamma<$ 1.
+Going back to (2.14b) again, the coefficient are in $C^{\gamma}$ for any $0<\gamma<1$, hence we have $\|F\|_{C^{1,\gamma}}\leq C_{5}$.
+Now, the coefficients of (5.8) are in $C^{\gamma}$ for any $0<\gamma<1$.
+Thus we derive $\|\phi\|_{C^{3,\gamma}}\leq C_{6}$ for any $0<\gamma<1$.
 Now the second equation (2.14b) has $C^{1,\gamma}$ coefficients and bounded right hand side, the classical $L^{p}$ estimate ([GT77, Theorem 9.11]) implies $F\in W^{2,p}$ for any finite $p$.
@@
 ∎
-5.B. Linearization of opeartors
+5.B. Linearization of operators
 In this subsection, we present the details of the computation for the linearization of operators required in Subsection 4.A. The computation basically follows [Has19, Lah19].
@@
 $-\int_{X}\sum_{\alpha}\mathrm{v}_{,\alpha}\langle\partial\big(J\xi^{\alpha}(\psi)\big),\overline{\partial}\dot{\phi}\rangle_{\phi}\cdot f\cdot\frac{\omega_{\phi}^{n}}{n!},$
-where $-\sum_{\alpha}\mathrm{v}_{,\alpha}\langle\partial\big{(}J\xi^{\alpha}(\psi)\big{)},\overline{\partial}\dot{\phi}\rangle_{\phi}=\langle\theta,\partial\mathrm{v}\wedge\overline{\partial}\dot{\phi}\rangle$.
+where $\sum_{\alpha}\mathrm{v}_{,\alpha}\langle\partial\big(J\xi^{\alpha}(\psi)\big),\overline{\partial}\dot{\phi}\rangle_{\phi}=\langle\theta,\partial\mathrm{v}\wedge\overline{\partial}\dot{\phi}\rangle$.
 For two $(1,1)$-forms $\alpha,\beta$, we have the formula
@@
 The linearized operator $\frac{d}{dt}{\mathcal{T}}-\langle\partial{\mathcal{T}},\overline{\partial}\cdot\rangle_{\phi}$ is a $(\cdot,\cdot)_{\mathrm{w},\phi}$- self-adjoint second order elliptic linear operator.
+In particular, if $\theta$ is positive, then the operator has negative first eigenvalue for $R>0$.